A Bibliography of Publications by, and about, Niklaus Wirth

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 March 2024 Version 1.07

Abstract

This bibliography records publications of Niklaus Wirth.

Title word cross-reference

+ [Wir76d]. $100M [BOB^{+}00]. $9.80 [Atk79a]. = [Wir76d]. N [Wir65e]. N ≤ 8 [Gol87]. SU(3) [Kae95].

* [Bie85].

-2 [Dub85]. -3 [Dub85]. -dienste [War03].

. [Bis81].

0 [FH15, SR15, Wal89]. 0-13-021999-1 [Wal89]. 0-387-9642 [Mee87]. 0-471-87834-0 [Ano87c]. 0-632-01609-4 [Bow88]. 07 [Kel21, Wir07e, Wir18].

1 [Dub85, Wal89, WW66b]. 1/3 [Lüt21]. 10 [GLN76, Nag76]. 10514-1 [ISO96, PS96, SP96]. 10514-2 [ISO98a]. 10514-3 [ISO98b]. 11’ [Hay80, BD76, Fei74, Hit79, Mah80, Mat80]. 11/45 [Fei74]. 1100 [Bal76]. 146 [Wir64b]. 15.95/$27.95 [Wal89]. 16BST [Dun84]. 17 [Ano86c]. 1900 [WQ72]. 1973 [GLL74]. 1983 [Ame83, Eur85]. 1988 [ISO88]. 1990 [AC90, ISO90a, ISO90b]. 1996 [ISO96]. 1998 [ISO98a, ISO98b].

2 [And85, And86, And89, AK92, Ano86b, Ano86c, Ano86d, Ano86a, Ano87a, Ano87b, Ano88, Pfi04, AJ90, Bai88a, Bai88b, Bat91, BJ86, BG88, BHHM86, Bie85, Bis91, Bla89, Bon87, Bös92, Bös93, BB91, BH86, Bru86a, Bru86b, Bru87a, Bru87b, Bud85, BDW88, BOB^{+}00, Car85, Cas91, Chr86, CLL^{+}86, CMM87, Col84, Cor88, Cra85, CI85, Det93, DNR90, DO85, Dot90, DSD97, Dun84, EBP93, Fel86a, Fel86b, Fel86c, Fel88, Floxx, For88, Fos86, GAK86a, GAK86b, GH93, Gol86, Gol87, GN87, GC88, Gut84, Gut86, Gut87, Har90, HS84, Hem88, Hir83, HH95a, HH95b, Hop80, ISO96, ISO98a, ISO98b, JPPM96a, JPPM96b, JPPM96c, JM84, Jor89a, Jor89b, Kel21, Kin89, KC90, Kin91, Kir95, KT90, Kof88a, Kof88b, Kre89, KP97a]. 2 [LG87, Lad89, Lea87, Lee85, LS87, Li88, Lo85, Lou90, MPT86, MS87, MS84, MS99a, Mof84, Moh89, MFMN89, Mös95a, Mös91, MW91b, MW91a, Mös93a, Mös93b, Mös93c, MW93, Mös95b, Mös95c, Mös98, MS99b, MLKK97a, MLKK97b, Mul84, Mul24, Nan89, Nie88, Nig93, Nik97, Ode89, Ohr84, Pad90, Pas85, PT92, Pir19, Pom85, Por88a, Por88c, Pou90, Pow84a, Pow84b, Pow84c, Pro92, PS96, PSSW97, Cre90a, Reh87, Rib90, RL01, Rov86, RD85, RD86, Sal92, SIRB16, Sav87, Sch88, Sch86, SH89, SP96, Sch89b, Sch89c, ST86, Sew87, Sew84a, Sew84b, Sew88, Sha86a, Sha86b, Sha87, SW86, Sou90, Spe82, Spe83, SW87, SG82, SM95, Tag88, Tem95, Ter86, Tho92a, Tho92b, Tor87a, WA86, Wie84, WS84, WS85, Wie86, Wir78a, Wir80b, Wir82, Wir83c]. 2 [Wir84a, Wir85c, Wir85b, Wir86a, Wir88c, Wir90a, Wir94b, Wir04b, Wir07c, Wir14, WF84, Woo86, Woo93, Woo94, WJ92a, WJ92b, Ano86c, Mee87, Ano86b, Ano87c, Bow88]. 2.25 [Atk83]. 2/86 [Ano86e]. 2/MUPE [MPT86]. 2003 [BS03]. 22.50 [Rob82]. 24 [Ano88, Wir65b]. 249 [Wir65e]. 26 [Ano86a]. 265 [Wir65c]. 28/29 [BD01]. 2nd [Wir76b].

3 [Aba94, Ano94, Bös92, Bös93, CDJ^{+}89, CDG^{+}92, ER92, Fer95a, Fer95b, FH98, Fre95, GW77, Gut94, Har90, Har92, Har94, HN91, HKMN93, Jor90a, Jor90b, Kin93, Kom98, LN98, Lev88, Mar94, Mas98, Nel91, RS95, Sed93, Vix92, Wya94]. 3.4 [Wir74c]. 3.50 [Bis79a]. 30 [Web67]. 30.11.201 [Wir14]. 360 [Fou75, Mal71, Mal72, RS76]. 360/370 [Fou75]. 370 [Fou75]. 3D [MTT83].

4.95 [Atk79b]. 41 [Ran87a]. 45

[Ano87a, Fei74]. 46 [Ano86d]. 4A [Nou05].

5 [Bar77]. 5-50 [Bar77]. 5.25 [Ros74]. 50 [Bar77]. 52 [Ano86b]. 55 [Ano87b].

6 [Des73]. 6.95 [Bis84]. 60 [Wir64a]. 6000 [Lid76, Wir72e, Wir74c]. 6000-3.4 [Wir74c].

7.50 [Bis79b]. 7185 [AC90, Eur85]. 7185-1982 [ISO82]. 7185-1983 [Eur85]. 7185-1990 [AC90]. 770.X3.97 [AC90]. 770.X3.97-1983 [AC90]. 770X3.97 [Ame83]. 770X3.97-1983 [Ame83]. 7th [LS06].

80th [Mar14]. 86 [Ano86e]. 8651-2 [ISO88]. ’88 [Gia89].

9.70 [Edw77]. 90 [MHT96, Mös95a]. ’97 [Mos97]. 99/4A [Nou05].

= [Edw77].

A. [LD74, Wir65a]. Abstract

[Bon87, EO80, GN87, KT90, Lea87, Sav87, SW87, Tor87b, Wie86]. Abstraction [BDW88, Moh89, RLN03, Sal79a, WA86, GH93]. abstractions [DNR90]. Access [Tag88]. accumulator [CRT80]. Achievements [GALW12]. ACM [Ash87]. acquiring [Ano80, Jos79, Jos80]. Action [BBS97]. Action-based [BBS97]. Active [DR97, Gut97]. Activity [Sal92]. Ada [Ano86b, Ano86c, Ano86d, Ano86a, Ano87a, Ano87b, Ano88, Gre86, Bie85, Col84, BBG^{+}81l, BK87, Eva84, Fel86a, FG84a, For88, Kin93, Sch86, SH89, ST86, SMB83, Wic84, WS84]. Adaptive [Sch01]. Addenda [HW74]. Adding [Kna97, LS94]. Addison [Bis79b]. Addison-Wesley [Bis79b]. Additional [Wir66a]. Additions [SP96]. address [CRT80]. Addyman [Bis79a]. ADTs [LS87]. advanced [Bai88a, Bai88b]. Again [Bon87, KPUG93, Mal72, Sal79b]. against [Tor87b, Wie86]. aged [Daw24]. aid [Bud85]. AIDA [CC87]. Air [DP85]. Akademgorodok [BBZ01, BBP96]. ALDES [Kre89]. ALDES/SAC [Kre89]. ALDES/SAC-2 [Kre89]. Algebra [KP97a, Kre89]. Algebraic [Zam96, Gia89]. algebras [Ran86, Ran87a, RWZ87, Ran87b]. ALGEC [Wir67e]. ALGOL [WH66b, WW66a, BBG68, BW71, Wir63a, Wir64a, Wir66a, WH66a, WW66b, WW66c, Wir68a, BBG^{+}69, Sit72, Wal86, Wir63b, WW65]. ALGOL-like [BW71]. Algorithm [Wir64b, Wir65c, Wir65e]. Algorithmen [Wir75e, Wir83a, Wir86a, Wir99a]. Algorithms [Sed93, SDK06, Wir76d, Wir84c, Wir84d, Wir86b, Wir04a, Wir75e, Wir83a, Wir86a, Wir99a, Edw77, Wal89]. Alias [MHT96]. Allow [Hei03]. Alone [FF00]. AlphaAXP [DG96]. Alternative [AK92, DO85, DSD97, Wir96e, Ano86b, CW82]. always [Lüt21]. Ambiguities [Spe82, Tor87a, WSH77]. ambiguity [Par85]. American [AI83, IEE83, Led84]. among [Sou90]. Analysis [OW83, SFIK80, WC81, BWA82, CL82, Coo89, CG94, MHT96]. Analysis/Synthesis [WC81]. Andra [GW84]. Andrei [BBZ01, BBP96]. Anniversary [Mar14]. annotations [Led84]. Anomaly [CC87]. ANSA [ER92]. ANSI [AI83, AC90, Ame83, Ten85]. ANSI/IEEE [AI83, AC90, Ame83]. ANSI/ISO [AC90]. Anwendung [War03]. APL [JF87, JF87]. APL2 [JF87, JF87].

Application [DP85, KS80, Szy92b].

Applications

[Mil89, MS99a, MS99b, Szy92b]. Applied [Amm74]. Approach [EBP93, Gut86, LS87, Li88, Ode89, Sch00a, Ano87c]. Approaches [WS85]. Approximation [CBG01]. Architect [Ros24]. Architecture [Wir10c, Wir15b, Wir17a]. Architectures [Wir86c, Wir87c, Wir87d, Wir87e, Wir87f]. Area [Tag88]. Arithmetic [Dot90]. Arm [KS84, Wir07d, Wir07f, Wir08g]. Array [LS79]. Arrays [Wir76a]. Art [Por88a, Sev12, Bös00, Mös00]. Article [Ano86b, Ano86c, Ano86d, Ano86a, Ano87a, Ano87b, Ano88]. Aspects [Mil89, NPJ79, PS96]. Assembler [GW78]. Assemblers [DS86]. Assembly [BW71, GW77, Ano86a]. assess [JF87, Wir76c, Wir84e]. Assessing [FG84a]. Assessment [CGHP79, Wir75f, Wir75g, Wir10a, Wir75a, Wir84b]. Association [Sof84, USE85]. assorted [Pou90]. Atkinson [Bis82]. attribute [KV15]. Aug [EN74]. August [Ano86a, BS76, BS03]. Austria [BS03, Mos97]. Author [Wir78c]. authors [Wir67a]. Automata [BMW91]. Automatic [FW65a, FW65b, Mar75]. avoids [JM84]. Award [Ash87]. Axiomatic [HW72, HW73, HW74, Wir70b, Hoa74].

B [Ano88, Bis84, Ree84]. B1800 [Hur84]. Baby [Aba94]. Back [BBG^{+}81a]. Backward [Shr79b, Shr79a]. Barrett [Bis81]. Base [ISO96, Dis93]. Based [AP84, ER92, LG87, MTT81, Rad94, Wir79c, Wir86c, BBS97, BS84, JF87, Mar79, KE85]. BASIC [JF87, Wir69a, Wir67b, JF87, Sha87]. basis [Ran87b]. Batch [Bri01a]. battle [Gut00a]. BC [HLW63]. Be [Coh91, Wir91a, Wir91c, PS94]. Beauty [FvGGM90]. Behaviour [Hur80]. benchmark [Dun84]. Benefits [Cas91, Com83]. Bericht [Wir76e]. Berlin [Atk79a]. between [IH71, MHT96, Mös91, MW91b, Mul84, Wir08d]. Beurteilung [JF87]. Beyond [Ano91, RW92]. binaries [FK97]. Binary [Gri90]. bindings [ISO88]. Birthday [FvGGM90]. Bit [And85, DG98]. Blackwell [Bow88]. block [Bri94a]. block-structured [Bri94a]. Blocks [Shr78]. Board [WL93]. boards [Pou90]. Bonn [BD01]. Book [Abr74, Ano79, Ano87c, Atk79a, Atk79b, Atk82b, Atk83, Bar77, Bis79a, Bis79b, Bis81, Bis82, Bis84, Bow88, Cyt97, Edw77, Fin77, Mee87, Ree84, Rob82, Ros74, Wal89, KPUG93]. Bookshelf [Mas98]. Bootstrap [GLN76]. both [Szy92a]. Bound [Wal86]. Bowles [Atk79a]. Brief [Wir08b, Wir08c]. BSD [War03]. Build [Rov86]. Building [Kom98, Ree00]. Business [FvGGM90, Wag00].

C [Ano86b, Bar77, Fin77, Rob82, AK92, Anoxxa, DO85, DSD97, Eva84, FG82a, FG84a, FG84b, HH95a, HH95b, Lad89, LP83, Mat84, Mös95a, SM95, Tem95]. calculate [Ran86, Ran87a]. Calls [Bri02a]. Cambridge [Bis81, Bis84, Fin77]. Can [Coh91, Wir77k, Wir91a, Wir91c, PS94]. canonical [Ran87b]. Case [CBG01, DS86, Gut00b, Hop80, Kof88a, Kof88b, Ree00, Sal81, Ano80, Ano86a, Atk82a, Jos79, Jos80, LF82, HM82]. Case-Study [Ree00]. CC [MFMN89]. CC-Modula [MFMN89]. CCNPASCAL [NPJ79]. CDC [Lid76, Wir72e]. Ceres [HN91, Wir90b]. Ceres-3 [HN91]. Ceres-Net [Wir90b]. certain [Wir67b]. Certification [Wir64b]. Challenges [GALW12, Sha87]. champion [VN24]. Changes [SP96, Wir72a]. Characters [Wir73a]. Charts [Mar14]. checker [LN98]. Checks [Wel78]. Chichester [Bis82, Ree84, Rob82]. Children [Fou91]. CHILL [SMB83]. chip [BOB^{+}00]. Christian [Mee87]. Ciechanowicz [Ree84]. Circles [Wir90d]. Circuit [WL93, Wir95a, Wir96a]. Circuits [Wir98d, Wir05d, Wir98c]. City [Sof84]. Clarifications [SP96, SM79]. Class [Lev88, SOM94]. classes [Szy92a]. Classic [Bri01a]. Clean [CH79, KT90]. Clebsch [Kae95]. Cleverbyte [Wir76c, Wir77a]. Cliffs [Abr74, Edw77, Ros74]. Closing [Wir68a]. closure [Dij00]. cloth [Bis81]. cluster [Ebe00]. cm. [Abr74]. co [Wir03b]. co-design [Wir03b]. Code [Alp79, Amm76, Amm77, Dot90, Fra94b, Lid76, Lou90, Sch89a, Wir72e, Wir86c, NAJN81, VN24, GW77, GW78, Hit79]. Code-Generation [Fra94b]. Coding [Wir69b, Wir70a]. coefficients [Kae95]. Collection [Wir79a, EMR91]. Colloquium [Wir68a]. COM [Gut99, Gut00b]. Combining [Bud85, DR97]. come [Pou91]. Comm [Ran87a]. Command [DS84]. Comment [Wir76a]. Comments [Gre86, Hab73, LD75, Wir65a, vD89, Hab84, Jos80, LD84]. communication [Szy90]. Companion [FH98]. company [BOB^{+}00]. Comparing [Col84, FG84a, Kel21, Lad89]. Comparison [BdJ76, BdJ80, Bös92, Bös93, FG82a, Mös95a, Nel79, Wir86c, Eva84, FG84b, Mat84, Ten85, Wic84, Dot90]. Comparisons [Liu86]. Compatibility [Ten78]. Compilation [BBG^{+}81k, Fos86, Gut86, Gut87, HM82, PT92, Wir98d, Wir98c, Wir07f]. Compiler [Amm74, Amm76, Amm77, Ber78, BD76, CMH85, CRT80, Cre90b, Cre91, Dot90, Fos86, Fou75, Fra94a, GLN76, Gut87, Gut00a, Hit79, Kir95, Lou90, Mah80, Mös00, Mul24, NAJN74, NAJN76, NAJN81, Pow84a, Pow84b, Ree84, RS76, SFIK80, Som06, WQ72, Wir69a, Wir71a, Wir75d, Wir86c, WG92, Wir96c, Wir96d, Wir05a, Wir07d, Wir07e, Wir08f, Wir17b, WJ92a, WJ92b, Amm75, BB91, Des73, Geo74, GM00, MS83, Cre90a, VS15, WC83, Wir77f, Rob82, Cyt97]. Compilerbau [Wir77f]. Compilerbaus [Wir96d, Wir08f]. Compilers [Amm75, Pro92, Sha86a, WC81, And84, Por88a]. completenes [CD84]. complex [Sch00b]. Complexity [Wir08e]. component [Pou95, Pfi04, Ano06a, Ano06b, Pfi01]. Components [Gut00b, Szy00]. Composable [Hof00]. Composition [Wec98, Wir74a]. Compositional [VS15]. Compression [BMW91]. Comput [Ran87a]. computation [Gia89]. computations [Shi92]. Computer [Ano24a, AP84, BW71, Bis79b, Daw24, DS84, Gru24, Gut87, Haá82, ISO82, Kor24, LM24, MS87, Sev12, Wir69b, Wir70a, Wir72e, Wir79c, Wir79d, Wir81a, Wir85a, Wir90b, Wir95a, Wir98b, Wir07a, Wir08a, Wir10a, Wir10b, AI83, Eur85, IAJ83, IEE83, Kre89, PMR00, Rei84, Wir81b, Wir87b, Wir91b, Wir98a, Wir99b, Wir00b, ISO88]. Computers [Bal76, Fou75, WQ72, Wir65d, Wir66c, Wir67d, Wir68b, Wir68c, Wir15a, Wir08a]. Computing [BST89, Bri96f, Bri02d, GLL74, Wir02a, Wir02b, Wir15a]. Concept [Rad94, Sew84a]. Concepts [Gut85, Wir67b]. Concerns [CLL^{+}86, Gol85, Mof84, SS85]. Concurrency [BK87, LS94, BBS97]. Concurrent [Bri77, Bri02a, Bru87a, Bru87b, DS86, Gol86, Lee85, MFMN89, NPJ79, NW78, SH89, Sch89c, WJ92a, WJ92b, Hay80, Mat80, Bri75a, Bri75b, Bri75c, Bri76a, Bri93, Bri96a, Bri96b, Bri96c, Bri96e, Bri01b, Bri02b, Bri02c, BK87, CGHP79, DSW82, GKLM79, Ker82, Kru82, Rav82, Shr79b, Shr79a]. concurrente [Sch86]. Conference [Ano91, BBZ01, BBP96, BS03, BD01, GW00, LS06, Mos97, Sof84, USE85]. Configuration [Jor89a, Jor89b]. conquered [Emr21]. Considered [SS85]. Consistent [BJ86]. Constant [Coh91, Wir91a, Wir91c]. constrained [KS89]. Construction [Wir77f, Wir85a, Wir96c, Wir96d, Wir05a, Wir07a, Wir08f, Wir17b, Geo74, Gut00a, Mös00, VS15, Wir87b, Wir00b, Cyt97]. Constructive [Bow88]. Constructs [MS90]. context [LL70]. context-free [LL70]. contextual [CL82]. Contracts [Wec98]. Contribution [Wir66a, WH66a, WH66b]. Contributions [BD01, BD02, Wir00a, BD01]. Control [Gut97, Wir73a, Wir98b, Bri94a, DNR90, Wir98a, Wir99b]. Controller [KS84]. controversial [KPUG93]. Conversational [AN81, Rob83]. Conversion [MS90, For88]. Converting [GAK86a, GAK86b, Sch89a, LP83]. Conway [Fin77]. COPAS [AN81]. Copying [FM86]. core [Wir98a]. Coroutines [KS80, Pad90, Ger82, HT86]. Corrections [Wir76b]. Correctness [Mey99, Old84]. Correspondence [Coh91, Wir78c, Wir91c]. Corrigenda [HW74, Wir70a]. Corrigendum [Wir68b]. cost [Wir90b]. Costs [Com83]. Counters [Wir15d]. countess [For88]. Course [Fel86b, Fel86c, GLW94, MS87, Wir69a, Bös98]. craftsmanship [Pom93]. creating [Ano86a]. Creator [Gru24, Cha24, Pro24, VN24]. Criteria [Wir72b]. Critical [BdJ76, BdJ80, Hab73, Hab84]. critique [Wir67e]. Cross [Hit79, Wir80a]. Crux [Wir08e]. CS1 [Kof88a, Kof88b]. CS2 [Kof88a, Kof88b]. CSP [CMM87, Gol87]. Current [Wir10a]. Curriculum [GAK86a, GAK86b]. curves [Ano87a]. Cyber [Lid76]. cybernetics [Wir67e].

D [Ano79, Ano86d, Atk79b, Fin77]. Daily [Hop00]. Dallas [USE85]. Daniels [Rob82]. Data [CC87, Edw77, EO80, Fel88, GH93, KT90, Lea87, Moh89, SW86, SW87, WA86, WS85, Wir75e, Wir76d, Wir83a, Wir84c, Wir84d, Wir86a, Wir86b, Wir99a, Wir04a, Ano87a, Wal89, Wir07f]. database [Rei84]. Dataflow [DS84]. Datenstrukturen [Wir86a, Wir99a, Wir75e, Wir83a]. David [Atk83]. Davos [GLL74]. dead [Qua24]. December [Ano87b, Ano88]. Decimal [Gri90]. Declarations [BBG^{+}81b]. decompose [RWZ87]. decomposition [Ran87b]. DECSystem [GLN76, Nag76]. DECSystem-10 [GLN76, Nag76]. Defense [BBG^{+}81l]. defensive [Jos80]. defined [DNR90]. Definition [CDG^{+}92, HW72, HW73, HW74, WW65, WW66b, WW66c, Wir70b, Hoa74, WW66a]. Definitions [Wil82, Wir77k]. demand [Wir76c, Wir84e]. Denkplatz [Wir08k]. Department [BBG^{+}81l]. Dependence [MHT96]. dependencies [RLN03]. Dependent [BBG^{+}81m, Ran86, Ran87a, RWZ87]. derivative [Wir04b]. Descent [Amm78a]. Description [ABBH^{+}79, BBG^{+}69, Floxx, HLW63, Wir94b, HN91, JM84, Jen73]. Design [GLW94, Geo74, JM84, Ode89, Wir71a, Wir72b, Wir5 b, Wir77g, Wir85a, WG92, WL93, Wir95a, Wir96a, Wir07a, Wir10b, Wir15b, GKLM79, HE82, NW84, Pou93, Pro24, Wir74b, Wir87b, Wir00b, Wir03b]. Designed [Wir79d]. Designing [Ebe00, KM95, Szy92b, Wir89]. Detecting [GN87]. Detection [CC87]. determination [Arm88]. develop [PS94]. Developers [Kir95]. Development [Amm74, Ano87c, Com79, KC90, Mul86, Wir66a, WH66a, WH66b, Wir71b, Wir71c, Wir83b, Wir93b, Wir96b, Wir00a, Wir01b, Wir02d, Amm75, Ano87b, Kre89, Nie88, NW84, VN24, Wir93a, Bow88]. Devices [Wir08e]. Diagnostics [WB85a, WB85b]. Dialogs [Kna95]. died [Ano24a, Cha24, LM24, Mey24]. dienste [War03]. Dies [Gru24, Kor24, Ros24, Daw24]. Differences [Mös91, MW91b, Mul84, Wir08d]. Digital [GLW94, Wir95a, Wir96a, Wir05d]. Dijkstra [FvGGM90]. Dimensional [MTT83]. direct [RWZ87]. Discipline [BS84, Wir77b, Wir77c, Wir77d, Wir77e, Wir77i]. Discrete [SDK06]. displays [Ano86a]. Distributed [BST89, Bri01a, Kom98, Lal91]. Diversity [Wir77k]. Division [Wir08i]. Djavaheri [Ano86b]. DM21.40 [Atk79a]. Do [Gut97, Wir77k, Wir08k]. Doctorates [Wir13]. Document [BBG^{+}81l, GW84, Zel97]. Does [Dot90, BG88]. Domain [Sch01]. Dr [Ano24b]. Draft [ABBH^{+}79]. Dragon [Gut87]. Drawing [Wir90d]. Dynamic [CC87, FM86, Sch00a, Wir76a].

E12.50 [Bis81]. E7 [Fin77]. E7-95 [Fin77]. easily [PS94]. Easy [Wal86]. Economic [Wel78]. Edison [KS84]. Edition [Wir76b]. Editor [Szy92b, Gut85, JCL85, Szy91, Ano80, Hay80, Jos80, Wir66b]. Editorial [Wir72d, Wir77a]. eds [Ree84]. Edsger [FvGGM90]. Education [Wir02a, Wir02b, Wir91b]. Effective [Fer95a, Fer95b, Wir95b]. Efficient [Gut86]. Eignung [JF87]. Einführung [Wir75h, Wir77f]. elegant [VN24]. Elements [BBG^{+}81h]. elimination [CG94]. Ellipses [Wir90d]. Ellis [Atk82b, Rob82]. Embedded [Nie88, Wir01a]. Emil [TWB11a, TWB11b, TWB11c]. Emmy [Alp79]. Empirical [HH95a, HH95b, Coo89]. Emulating [Fra93]. End [Kna95]. Ended [Lea87]. Engelwood [Edw77]. Engineer [Wir81a, Lüt21, Wir81b]. Engineering [BD01, BD02, Pom85, Pom93, SOM94, Wir08b, Wir08c, Ano86d, Bud85, WS84]. Englewood [Abr74, Ros74]. Entwicklung [Amm75]. Enumerated [Cas91]. Enumerations [Lin90]. Environment [FM86, GC88, Hir83, Jor90a, Jor90b, MPT86, Moh77, MLKK97c, Lal91, Szy90]. environments [Pou95]. Erfinder [Qua24]. erklärbar [Wir08a]. Errata [WH66b, WW66a]. Erratum [Ran87a]. Error [Amm78a, Reh87, Shr79b, Shr79a]. errors [LF82]. Ershov [BBZ01, BBP96]. erstreben [Wir08k]. essay [Wir99c]. Essays [Bri96d]. Essence [Nau94, Wir03a]. Euclid [BET78, BK87]. EULER [WW66a, Web67, WW66b, WW66c, WW65]. Evaluating [GC00]. Evaluation [SIRB16, Mey99]. Event [Mar84]. Evolution [Pfi04, VN24]. Examples [Shr79b, SH89]. Exception [EBP93, Pad90]. Exceptions [BBG^{+}81c]. Execution [CBG01]. EXecutor [KE85]. exemples [Sch86]. exercise [Sal92]. Experience [Ber78, Bri77, Lou90, Sch88, Tag88, Lec73]. Experiences [Jor89a, Jor89b, NW78]. Experiment [PD81]. Experimental [WL93]. Experimentation [Wir08h]. Experiments [Wir10b, KS89]. explainable [Wir08a]. Explosion [Wir94a]. Expressions [BBG^{+}81i]. Extended [ISO91, LN98, PS94, ISO90a]. Extending [Hei03, Rov86, Tsi82]. Extensibility [Mös94]. Extensible [Jor90a, Jor90b, Mar96, PS94, Saw96, Szy92b, Szy91]. Extension [Coh91, KS80, Liu86, MTT81, MTT83, MTG89, Wir87a, Wir91c, WL93, BH95, Ger82, Sal92, Wir91a]. Extension-Board [WL93]. Extensions [CMH85, Lea87, Nig93, PSSW97, SM79, Wir88e]. external [Tsi82]. extract [Wir77a].

F [Pou95]. Facilities [Sew84b]. Facility [Hur84, Wir79b, CW82]. family [BCFT92, BCFT95]. Fast [ST86]. Father [Pos24]. Favorite [Ker81, Ker84, Bös98]. Features [BBG^{+}81m, Shr79b]. Federal [Ano85]. Fertigung [Pom93]. Few [Pes97]. FFF97 [Tem00]. figure [VN24]. Files [Gut86]. Find [Wir65c]. Findally [Atk79b]. finding [IH71]. Finite [Wir08e]. First [HH95a, HH95b, Li88, SIRB16, Ash87, Bös98]. first-course [Bös98]. Fish [Gut97]. Flaws [JM84]. Flexible [Wil82]. Flight [Wir98b, Wir98a, Wir99b]. Floating [Dot90, Dun84]. floating-point [Dun84].

Flow [CC87]. Fly [Fra94b]. forces [Sch89b]. Ford [Ano87c]. form [LL70]. FORMAC [Wir65b]. Formal [KP97b, WW65, WW66b, WW66c, BH95, WW66a]. Formatted [Woo86]. Formatting [BS84]. Formelsammlung [EMR91]. formulas [EMR91]. Fortran [CNW80, Ker82, Mös95a, CCN^{+}79, Fre81, MHT96, Moh77, PD81]. Fortran-to-Pcode [CCN^{+}79]. Fortran-to-Universal [CNW80]. Forum [FG82b, Pet82]. Foundations [Atk82b]. founded [Dij00]. four [Por88a]. Fourth [BBZ01]. FPGA [WL93, Wir15b]. FPGAs [GLW94, Wir96a]. Framework [Szy92b, KM95, Pou95]. Frameworks [GC03]. free [LL70, Ran86, Ran87a, RWZ87]. FreeDOS [War03]. Fresh [Fra97]. Front [BBG^{+}81d]. fully [BG88]. FUMBLR [McC83]. functional [Ode00]. Functions [BH86, Wir65c]. Fundamentals [Wir96d, Wir08f]. Future [GALW12, Gut94, Pou93].

G [Ano79, Ano87a, Bar77]. gadgets [Pou90]. Gary [Ano87c]. Gast [War03]. Gast-Systeme [War03]. Gates [KPUG93]. Gedanken [Wir94a]. Generalization [Wir63b, Wir63a, WW65, WW66b, WW66c, Wir75b, KM95, WW66a]. Generate [Dot90]. Generated [Dot90]. generating [Kae95]. Generation [Amm76, Amm77, Fra94b, Kor24, Lid76, Wir72e, Wir86c]. Generator [Sch89a, Wir80a, MS83]. Generic [BBG^{+}81e, CI85, Gol86, LS87, Wie84, WS85]. Generics [BJ86, ISO98a]. German [Amm75, EMR91, JF87, Mös93c, Mös98, Pom85, Pom93, Qua24, RW94, Wir72c, Wir74d, Wir75e, Wir75h, Wir76e, Wir77f, Wir83a, Wir85b, Wir86a, Wir90c, Wir94a, Wir96d, Wir99a, Wir08a, Wir08f, Wir08k]. GKS [ISO88]. glance [Wir18]. Glass [Wir05b, Wir06]. Glimpse [Pou93]. global [Old84]. GNU [Mul24]. Goals [Wir84a]. goes [Rib90]. Goldstein [Ano86d]. Good [Wir05b, Wir06]. Gordan [Kae95]. GPIB [Ano86d]. Grading [FW65a, FW65b]. grammar [KV15]. grammars [LL70]. Graphical [MTT81, MTT83, JCL85, ISO88]. Graphics [Det93, Mil89, ISO88, Yip84]. grew [Har90]. grid [Sch01]. Gries [Fin77]. Grogono [Bis79b]. Group [Sof84]. Growing [Wir08e]. Grundlagen [Wir96d, Wir08f]. Guest [Wir72d, Wir77a]. Guide [Chr86, FH98, Mas98, Mee87, Rei91, Tem95, Atk83, Gut90, Tem90]. Guidelines [Kir95].

H [Bis81]. Habermann [LD74]. Hall [Abr74, Edw77, Ros74, Wal89]. handbook [CK81]. Handbuch [War03]. Handling [EBP93, Pad90, Reh87]. Handwerk [Pom93]. Hardback [Atk82b, Bis82, Ano79, Bis84, Mee87, Ree84]. Hardware [HN91, Wir87c, Wir87d, Wir87e, Wir87f, Wir98c, Wir98d, Wir03b, Wir98a]. Hardware/software [Wir03b]. Hashing [BH86]. Heap [Mar79, Sch80]. Heap-based [Mar79]. heaven [Wir77a, Wir76c]. held [BS76, EN74]. Helicopter [Wir98b, Wir98a, Wir99b]. Hemel [Wal89]. Hempstead [Wal89]. Hermes [Hir83]. Heterogeneous [ER92]. Heyden [Bar77]. hierarchies [BS76]. High [Cra85, DSD97, Wir79b, Wir79c, Wir79d]. High-Level [Wir79b, Wir79c, Wir79d]. High-Performance [DSD97]. highly [BB91]. highly-portable [BB91]. Historic [Bös05, BD01]. Historical [Wir10a]. History [BG96, Bri96a, Ric74, Wir84a, Wir08b, Wir08c, Bri93, Bri96b]. Hobby [Wil80]. Honorary [Wir13]. Honours [Wir13]. Horwood [Atk82b, Rob82]. HotSpot^{T M} [GM00]. Human [Bös05].

IBM [Ano87b, Fou75, Mal71, Mal72, RS76, Sha86a, Web67]. ICL [WQ72]. Ideas [Wir05b, Wir06]. Identification [ST86, Ran86, Ran87a, RWZ87, Ran87b]. Identifiers [And86]. Idiomatic [FG82b, Pet82]. IEC [ISO91, ISO96, ISO98a, ISO98b, PS96, SP96]. IEEE [AI83, Ano86d, AI83, AC90, Ame83]. ihn [Wir08k]. II [BG96, Fou91, RWZ87, WW66c]. III [Ran87b]. Ilem [Wal86]. illustrated [SH89]. Illustré [Sch86]. imperative [Bai88a, Bai88b]. implement [GKLM79]. Implementation [AP84, Bal76, Bla89, BBG^{+}81m, Bri88, BDW88, CMM87, EBP93, Fei74, Gut00b, HS84, Mac79, Mar79, Mat80, NAJN74, NAJN76, NAJN81, OW83, PD02, Rei84, RS76, Sal81, Shr79a, WH86, Wir75d, Wir75c, Wir5 b, Wir77g, Wir81c, Wir07b, Wir14, Wir15b, Amm78b, Bar78, BBG68, DSW82, FH15, Fra90a, Gut99, HE82, Ker82, Par85, Rob82, Tem90, Web67, Hay80]. Implementations [BdJ76, BdJ80, Wir73a]. Implementing [Bis79c, CI85, Li88, Por88b, Sal79b, WS85, Ano86d]. Import [Szy92a]. Imported [And86]. Improving [You81]. Independent [LS87]. Index [Sch01]. Indicators [Atk79c]. Industrial [Ree00, Pom93]. industriellen [Pom93]. Informatics [BBZ01, BBP96, Wir08l]. Information [Ano85, ISO91, Wir67a, AC90, ISO88, ISO90b, ISO96, ISO98a, ISO98b, Bis84]. ing [Szy92b]. Inheritance [Wec98, Szy92a]. Initializing [McC83]. Input [BBG^{+}81f, Com83, Cra85, Lea87, Ode89]. Input-Output [BBG^{+}81f]. Input/Output [Cra85, Lea87, Ode89]. Insecurities [Spe82, Tor87a, WSH77]. Inspiration [Pos24]. Inspired [Kor24]. Installation [War03]. Instances [Bon87]. instruction [Coo89]. Integrated [GC88, Rov86]. Integrating [ADDM84]. Integration [Zel97, Wir64b]. intelligent [JCL85]. Interactions [MPT86]. Interactive [Com83, ORT81, NW84]. Interface [BHHM86, Bri76b, KT90, Kna95, WC81]. Interfaces [HKMN93, BS76]. Interfacing [Ano86a]. Interference [Bri94a]. Interlude [Wir78c]. Intermediate [GW77, GW78, Nel79]. Intermodular [Ste98]. International [Ano91, BBZ01, BBP96, Gia89, Kin89, Wal89, BS76, EN74, GLL74]. Interoperability [Gut00b]. Interpreter [Alp79, Hur84, JPPM96a, JPPM96b, JPPM96c, Sch89a]. Interpreter/Visualizer [JPPM96a, JPPM96b, JPPM96c]. Interrupts [Wir08g]. Interview [Hol21, TWB11a, TWB11b, TWB11c]. interviews [KW85]. intrinsically [Sch00b]. Introducing [Ano94, Wya94]. Introduction [BBG^{+}81g, MS84, MLKK97a, Nik97, Wir65b, Wir73c, Wir77f, Ano79, Atk79b, Bai88a, Bai88b, Bis79a, Har90, Abr74, Wir75h, Ros74, Bar77]. Introductory [Lee85, Wir95a]. Intuitionistic [Li88]. Inventor [Wir78b, Qua24]. IO [Lo85]. Irregular [Sch00a]. ISBN [Ano87c, Bow88, Mee87, Wal89]. isn’t [Knu90]. ISO [ISO90a, Ten85, AC90, Eur85, JW85, JW91, PS96, SP96]. ISO/IEC [ISO91, ISO96, ISO98a, ISO98b, PS96, SP96]. ISSAC [Gia89]. Issues [BBG^{+}81k]. IT-Welt [Qua24]. Italy [Gia89]. Ivory [Fel86a]. IX [Wir67e].

J [Ano86d, Bis81, Ree84]. J. [Ano86b, Ano86c, Ano86d, Ano86a, Ano87a, Ano87b, Ano88]. January [USE85]. JastAdd [FH15]. Java [Bös98, GC00, GM00, SIRB16]. Jerry [Pou90]. Jewel [Fra00a, Fra00b]. JKU [Ano24b]. JML [Mos97]. JMLC [BS03, GW00, LS06]. Job [Bri76b]. John [Atk83]. join [Sch89b]. Joint [BS03, GW00, LS06, Mos97]. Joyce [BA88, Bri88]. July [Ano86a, BS76, BBZ01, Gia89].

July/August [Ano86a]. June [Ano86b, Ano86c, Ano86d, BBP96, BD01, Sof84].

Kaare [Mee87]. Ken [KW85]. Kenneth [Atk79a]. Kermit [And89]. Kernel [GC88, ISO88, Ter86]. Key [Fra94b]. Kiama [SR15]. Kingdom [Wal89]. Klagenfurt [BS03]. Konfiguration [War03].

L [Ano86d, Ano88, Atk79a, Bis84]. L2 [Wir65c]. Laboratory [GLW94, Ano86d]. Lagoona [Fra97]. Lake [Sof84]. Language [Ame83, Ano06b, Ano18, AP84, Bal76, BBG^{+}69, BW71, BdJ80, BBG^{+}81l, Bri94b, Bri96c, Bri96f, Bri02b, Bri02d, CDG^{+}92, Cor88, ER92, Fes24, Floxx, Fra97, GW77, GW78, Gut00b, Hab73, Har94, HH95a, HH95b, HW72, HW73, HW74, Hur84, ISO82, ISO96, Ker81, Ker84, LD75, Mac79, MPT86, Mal71, Mal72, Mar79, MW91a, MW93, Nik97, Rad94, Rob83, SIRB16, Shr79b, Tem94, WB79, Wir65d, Wir66c, Wir67d, Wir68b, Wir68c, Wir70b, Wir70c, Wir71d, Wir72f, Wir72g, Wir72b, Wir73b, Wir75f, Wir75g, Wir77h, Wir79c, Wir79d, Wir85a, Wir88d, Wir90f, Wir94b, Wir96a, Wir98b, Wir01c, Wir02e, Wir05c, Wir07b, Wir07a, Wir16, AC90, Ano86a, BH95, Bar78, Bös98, Bri75a, Bri75b, Bri94a, Dub85, Eur85, Emr21, Gut99, Hab84, Har90, Hoa74, IAJ83, IEE83, ISO88, LD74]. language [Lec74, LD84, Mey99, Mös91, Pir19, Pou91, Rei84, Sal92, SOM94, Wir72a, Wir75a, Wir5 a, Wir76e, Wir84b, Wir87b, Wir90e, Wir90c, Wir00b, Woo94, BS76, AI83]. Language-Sensitive [Rob83]. Languages [BST89, BdJ76, BS03, FG82a, FG84a, GC00, GC03, GALW12, LS06, Nel79, SMB83, Wir79b, Wir87c, Wir87d, Wir87e, Wir87f, Wir00a, Wir03a, BG96, Eva84, FG84b, GW00, ISO90b, ISO91, ISO96, ISO98a, ISO98b, JF87, Mos97, Sou90, Wir67b, Wir74b, Wir76c, Wir84e]. Lara [Gut85]. Large [Com79, Rov86, You81]. latest [Pou91]. Laurence [Bis82]. Lawrie [Atk82b]. Lazy [Com83]. leader [KPUG93]. Lean [Sch00b, Wir95d, VN24]. Learning [Geh00]. Lectures [Ash87]. Less [CLL^{+}86]. Letter [Ano80, Wir66b]. Letters [Hay80, Jos80]. Level [Cra85, Wir79b, Wir79c, Wir79d, Tag88]. Leveraging [GC03]. Levi [Ran87b]. Levrat [Ano88]. Lexical [BBG^{+}81h]. Lexing [Spe83]. Library [BHHM86, Ode89, ADDM84]. Lie [Ran87a, Ran86, RWZ87, Ran87b]. life [Hop00]. Lightweight [Som06, Pou95]. like [Arm88, BW71, Mey99, Old84]. LILITH [Wir81b, Coo89, DS84, GW84, Ohr84, Pom85, Rei84, Wag00, Wir81a]. Lilith/Modula [DS84]. Limits [Gut87]. line [Mey99]. LineDrawing [Por88b]. Lines [Wir90d]. linguistic [JF87]. Link [Fer95a, Fer95b]. Link-Time [Fer95a, Fer95b]. Linked [Lev88]. Linux [War03]. Linz [Mos97]. LISP [JF87, JF87]. List [IH71, Lev88, Zim85]. Lists [BH86]. Local [Sha86b, SM95, Tag88]. Logic [Li88, Wir94b, Wir08e]. Lola [Wir94b, Wir95c, Wir96a, Wir14]. Lola-2 [Wir94b, Wir14]. London [Bar77]. Look [Fra97, Lid76, Ten78, JM84, KPUG93, Pou91]. Looking [Wir05b, Wir06]. looks [Pou90]. Loss [Ano24b]. Lotus [Gut00a]. Loughborough [Ano91]. Low [Tag88, Wir90b]. Low-cost [Wir90b]. Low-level [Tag88]. LSI [Hay80, Hit79, Mah80, Mat80]. LSI-11’ [Hay80, Hit79, Mah80, Mat80].

M [Ano79, Ano86b, Bis79a, Rob82]. Mac [Pou95]. Machine [Wir69b, Wir70a, Zam96, GM00, Jen73, GC00]. Machines [Lid76, Wir08e]. Macmillan [Bis79a]. MacOberon [Fra90a, Fra90b]. Macro [BO83, Com79]. made [BOB^{+}00]. mainframe [DSW82]. Making [AJ90]. Managed [GC03]. Management [Jor89a, Jor89b]. Manager [ORT81]. MANGLE [Sew87]. Manipulation [And85, Bis84, Wir64a]. Manual [Ano86e, BBG^{+}81l, JW74b, JW75, JW85, JW91, Rei91, Sit72, Fra90b, JW74a]. Many [Pos24]. MAP [Com79]. Mapping [Anoxxa]. March [Mos97]. Marktoberdorf [BS76]. Marshall [Ano86d, Bow88]. MAS [KP97a]. MASCOT [Bud85]. Mass [Fin77]. Masses [Ano18]. Master [BK87]. Master/Slave [BK87]. Matching [Liu86]. mathematics [EMR91]. Mathematik [EMR91]. Matter [BBG^{+}81a, BBG^{+}81d]. May [Ano86b, Ano86c, Ano86d]. May/June [Ano86b, Ano86c, Ano86d]. Measuring [DP85]. Mechanism [EBP93, Tsi82]. Medos [Knu00]. meets [And89, For88, Wag00]. Memo [Wir98a, Wir98b]. Memorial [BBP96, BBZ01]. Memoriam [GS24, Spi24]. Memory [McC83]. Message [Hof00]. Messages [FF00]. Messer [Bow88]. Method [Amm74, Amm75]. Methode [Amm75]. methodical [Atk79b]. Methodology [Mac79]. Methods [Sch01, KS89]. Metrowerks [Tay90]. Microcoded [CMH85]. Microcomputer [Bow77, OW83, Atk79a]. Microcomputers [Car85, Ano87a]. Microcontroller [Wir05c, Wir08h]. Microprocessor [Wir86c]. microprogrammed [Web67]. Microsoft [KPUG93]. Minicomputer [CRT80]. Minicomputers [BD76]. Minimal [BH86]. MINOS [Ode89]. MIRA [MTT83]. MIRA-3D [MTT83]. MIZAR [RD85, RD86]. Model [MS99a, MS99b, WH86, Wir98b, Web67, Wir98a, Wir99b]. Modern [ER92]. Modest [Zim85]. Modification [CRT80]. Modula [Ano86d, Ano86a, Ano87b, Ano91, Gol85, Gre86, Pom85, SS85, Tay90, Tor87b, Bud85, BOB^{+}00, CLL^{+}86, Col84, Gol86, PT92, Sou90, Wir5 a, Wir5 b, Wir78a, Wir80b, Wir90a, Aba94, And84, And85, And86, And89, AK92, Ano86b, Ano86c, Ano86d, Ano86a, Ano86e, Ano87a, Ano87b, Ano88, Ano94, AJ90, Bai88a, Bai88b, BET78, Bat91, BJ86, BG88, BHHM86, Bie85, Bis91, Bla89, Bon87, Bös92, Bös93, BB91, BH86, Bru86a, Bru86b, Bru87a, Bru87b, BDW88, BK87, CDJ^{+}89, CDG^{+}92, Car85, Cas91, Chr86, CMM87, Cor88, Cra85, CI85, DS84, Det93, DNR90, DO85, Dot90, DP85, Dun84, EBP93, EO80, ER92, Fel86a, Fel86b, Fel86c, Fel88, Fer95a, Fer95b, For88, Fos86, Fou91, Fre95, GAK86a, GAK86b, GH93, Gol87, GN87]. Modula [GC88, Gut84, Gut86, Gut87, Har90, Har92, Har94, HS84, Hem88, Hir83, HH95a, HH95b, Hop80, HKMN93, Hur84, ISO96, ISO98a, ISO98b, JPPM96a, JPPM96b, JPPM96c, JM84, Jor89a, Jor89b, Jor90a, Jor90b, Kin89, KC90, Kin91, Kin93, KT90, Kof88a, Kof88b, Kom98, KP97a, LG87, Lad89, Lea87, Lee85, LS87, LN98, Lev88, Li88, Lo85, Lou90, MPT86, Mah80, Mas98, MS87, MS84, Mof84, Moh89, MFMN89, Mul84, Mul86, Mul24, Nan89, Nel91, Nie88, Ode89, Ohr84, Pad90, Pas85, Pom85, Por88a, Por88c, Pou90, Pow84a, Pow84b, Pow84c, Pro92, PS96, PSSW97, Reh87, Rei84, RW92, Rib90, RL01, Rov86, RS95, Sal92, SIRB16, Sav87, Sch88, Sch86, SH89, SP96, Sch89b, Sch89c, ST86, Sed93, Sew87, Sew84a, Sew84b, Sew88, Sha86a, Sha86b, Sha87]. Modula [SW86, SMB83, Spe82, Spe83, SW87, SG82, SM95, Tag88, Ter86, Tho92a, Tho92b, Tor87a, Vix92, WA86, Wie84, WS84, WS85, Wie86, Wir77g, Wir77h, Wir77j, Wir82, Wir83c, Wir84a, Wir85c, Wir85b, Wir86a, Wir88b, Wir88c, Wir90e, Wir04b, Wir07c, WF84, Woo86, Woo93, Woo94, WJ92a, WJ92b, Wya94, Zim85, Ano86c, Mee87, Ano86b, Ano87c, Bow88]. Modula-2 [Ano86d, Ano86a, Ano87b, Ano88, Ano91, Gol85, Gre86, Pom85, SS85, Tay90, Tor87b, Bud85, BOB^{+}00, CLL^{+}86, Col84, Gol86, PT92, Sou90, Wir78a, Wir80b, Wir90a, And85, And86, And89, AK92, Ano86b, Ano86c, Ano86d, Ano86a, Ano86e, Ano87a, Ano87b, Ano88, AJ90, Bai88a, Bai88b, Bat91, BJ86, BG88, BHHM86, Bie85, Bis91, Bla89, Bon87, BB91, BH86, Bru86a, Bru86b, Bru87a, Bru87b, BDW88, Car85, Cas91, Chr86, CMM87, Cor88, Cra85, CI85, Det93, DNR90, DO85, Dot90, Dun84, EBP93, Fel86a, Fel86b, Fel86c, Fel88, For88, Fos86, GAK86a, GAK86b, GH93, Gol87, GN87, GC88, Gut84, Gut86, Gut87, Har90, HS84, Hem88, Hir83, HH95a, HH95b, Hop80, ISO96, ISO98a, ISO98b, JPPM96a, JPPM96b, JPPM96c, JM84]. Modula-2 [Jor89a, Jor89b, Kin89, KC90, Kin91, KT90, Kof88a, Kof88b, Kre89, KP97a, LG87, Lad89, Lea87, Lee85, LS87, Li88, Lo85, Lou90, MPT86, MS87, MS84, Mof84, Moh89, MFMN89, Mul84, Mul24, Nan89, Nie88, Ode89, Ohr84, Pad90, Pas85, Pom85, Por88a, Por88c, Pou90, Pow84a, Pow84b, Pow84c, Pro92, PS96, PSSW97, Reh87, Rib90, RL01, Rov86, Sal92, SIRB16, Sav87, Sch88, Sch86, SH89, SP96, Sch89b, Sch89c, ST86, Sew87, Sew84a, Sew84b, Sew88, Sha86a, Sha86b, Sha87, SW86, Spe82, Spe83, SW87, SG82, SM95, Tag88, Ter86, Tho92a, Tho92b, Tor87a, WA86, Wie84, WS84, WS85, Wie86, Wir82, Wir83c, Wir84a, Wir85c, Wir85b, Wir86a, Wir88c, Wir04b, Wir07c, WF84, Woo86, Woo93, Woo94, WJ92a, WJ92b, Ano86c, Mee87, Ano86b, Ano87c, Bow88]. Modula-2/86 [Ano86e]. Modula-2/MUPE-2 [MPT86]. Modula-3 [Aba94, Ano94, Bös92, Bös93, CDG^{+}92, ER92, Fer95a, Fer95b, Fre95, Har90, Har92, Har94, HKMN93, Jor90a, Jor90b, Kin93, Kom98, LN98, Lev88, Mas98, Nel91, RS95, Sed93, Vix92, Wya94]. Modula-Prolog [Mul86]. Modula/R [Rei84]. Modular [BS03, Bri77, GC03, GW00, LS06, Mos97, SM81, WB79, Wir77h, EF88, KV15, BS03, GW00, LS06, Mos97]. Modulas [Dub85]. Modulas-1 [Dub85]. Module [Por88b, Rad94, Wie84, Wir79b, CW82]. Modules [LS87, Sha86b, SM95, Szy92a, Szy00, ADDM84, PS94]. Molau [Ano87a]. Monitor [Sew84a]. Monitoring [Hur84]. Monitors [Bri93, Bri96a, Bri96b, Bri78b, Ter86]. Moore [Atk82b]. most [Arm88]. most-recent [Arm88]. Mourns [Ano24b, Qua24]. Move [AJ90]. MS [Dis93]. MS-DOS [Dis93]. Multi [Sch01, Kru82]. Multi-grid [Sch01]. multi-user [Kru82]. Multimedia [Pou90]. Multiple [Fel86b, Fel86c, Wir64b]. Multiprocessing [Wir96e]. Multiprocessor [AP84, Bri78a, BA88, Bri88]. Multiprogramming [WB79, Wir69b, Wir70a, Wir77h, Wir5 a]. Multitasking [Bis91, CBG01]. MUPE-2 [MPT86]. My [Ker81, Ker84, Bös98]. myself [Lüt21].

N [Edw77, LD74, Ros74, Wal89]. N. [Hop00, Mar14, Wir00c]. N.J [Abr74, Edw77, Ros74]. N.Wirth [Wir14]. nach [Amm75]. Names [BBG^{+}81i]. National [IEE83]. Need [Gut97, BG88, Szy92a]. needed [PS94]. neglected [Wir07f]. NELIAC [HLW63]. Nerima [Ano88]. Nested [Hei03, Pad90]. Net [Wir90b]. nets [Ode00]. NetWare [War03]. Network [ER92, Szy90, Tag88, Wir90b, Ebe00, Bri78a]. Netzwerke [War03]. Netzwerkkonfiguration [War03]. neue [RW94]. News [GS24]. Newton [KS89]. Newton-type [KS89]. Niklaus [Ano24b, Cyt97, Fes24, Abr74, Ano18, Ano24a, Ano24c, Bös00, Cha24, Daw24, GS24, Gru24, Hol21, Kor24, LM24, Mey24, Mös00, Pes97, PMR00, Pos24, Pou91, Pou95, Pro24, Qua24, Ros24, Sev12, Spi24, TWB11a, TWB11b, TWB11c, VN24]. nil [Ran86, Ran87a]. No [Ano79, Ano86b, Ano86c, Ano86d, Ano86a, Ano87a, Ano87b, Ano87c, Ano88, Atk79a, Atk79b, Atk82b, Atk83, Bar77, Bis79a, Bis79b, Bis81, Bis82, Bis84, Bow88, Fin77, Mee87, Ree84, Rob82, Ros74, Wir73a, Edw77, Ran87a]. noch [Wir08a]. nonlinearly [KS89]. Nook [Cyt97]. Notation [Wil82, Wir77k]. Note [Tor87b, Wir76a, Wir08i, Ger82, LL70, Wir66b]. Notes [NAJN74, NAJN76, NAJN81, PHT91, Wir66a, Wir75d, Wir95c]. Notes^{T M} [Gut00a]. notice [Ran87a]. Novell [War03]. November [Ano87b, Ano88]. November/December [Ano87b, Ano88]. Novosibirsk [BBZ01, BBP96, EN74]. Nr [Wir98a, Wir98b]. Nucleus [Hop80]. numerical [EMR91, KS89]. numerischen [EMR91].

O [Syd94]. Oakwood [Kir95]. Obeisance [Wir78b]. Oberon [RW94, War03, Dis93, Mil89, Pfi04, Anoxxb, Bös92, Bös93, BCFT92, BCFT95, Cre90b, Cre91, DR97, DG96, DSD97, DG98, FH98, Floxx, FH15, Fou91, Fra00a, Fra00b, Gut90, Gut94, Gut97, Gut99, Gut00b, Hof00, Kel21, KPUG93, Kir95, Kna95, Kna97, KP97b, Lal91, LS94, Lin90, Mar94, Mar96, MS99a, Mös95a, MTG89, Mös91, MW91b, MW91a, Mös93a, Mös93b, Mös93c, MW93, Mös94, Mös95b, Mös95c, Mös98, MS99b, MLKK97a, MLKK97b, Nau94, Nig93, Nik97, PHT91, Pir19, Pou91, Pou93, Pou95, Rad93, Rad94, Ree00, Cre90a, Rei91, RW92, RW94, RL01, Saw96, SR15, Som06, Szy90, Szy91, Tem90, Tem94, Tem95, Tem00, Wir88b, WG8 , Wir88a, Wir88d, WG89, Wir90e, Wir90f, Wir90c, WG92, Wir98b, Wir04a, Wir04b, Wir07c, Wir07d, Wir07e]. Oberon [Wir08d, Wir08g, Wir16, Wir18, Zam96, Zel97, vD89]. Oberon-0 [FH15, SR15]. Oberon-07 [Kel21, Wir07e, Wir18]. Oberon-2 [Pfi04, Bös92, Bös93, DSD97, Floxx, Kel21, Kir95, MS99a, Mös95a, Mös91, MW91b, MW91a, Mös93a, Mös93c, MW93, Mös95b, Mös95c, Mös98, MS99b, MLKK97a, MLKK97b, Nig93, Nik97, Pir19, RL01, Tem95]. Oberon-ARM [Wir08g]. Oberon-System [Kna97]. Oberon/F [Pou95]. Oberon0 [KV15, VS15]. Obituary [Ano24c]. Object [Bon87, ER92, Fra97, MTG89, Mös93b, Mös95c, Rad94, Tho92a, Tho92b, Wec98, BG88, Pou93, Pou95, Ste98, Wir90a, ISO98b, MTG89, Mös93c, Mös98, Tes85, Ano86c]. Object-Based [ER92]. Object-Orientation [Fra97]. Object-Oriented [MTG89, Mös93b, Mös95c, Rad94, Pou93, Pou95, Ste98, Wir90a, Ano86c]. Objects [Bla89, DR97, ER92, GN87, Gut97, Sav87, Tor87b, Aba94, BBS97, Gut99, Wie86]. Objektorientierte [Mös93c, Mös98]. occam [KS84]. October [Ano87a]. Ofront^{T M} [Tem95]. on-line [Mey99]. On-the-Fly [Fra94b]. One [CRT80, Gri90, Gut87]. One-address [CRT80]. One-pass [Gut87]. Online [Zel97]. Only [Lea87]. OOP [Har90]. OP [Cre90a]. OP2 [Cre90b, Cre91]. Open [Lea87]. Open-Ended [Lea87]. Operating [BHHM86, Bri76a, Bri76b, Bri96e, Bri01a, Bri01b, Bri02c, Bru86a, Bru86b, Fra93, Hir83, WG92, WF84, Kru82, Pou93]. Operation [Fel86b, Fel86c]. Operation-Systems [Fel86b, Fel86c]. Operations [LS79]. Optimization [Fer95a, Fer95b, SDK06]. Optimizations [Fra94a]. Optimizing [Atk82a, Lou90, Pow84a, Pow84b]. Order [Li88]. Organization [Amm78a, Wir69b, Wir70a]. Orientation [Fra97]. Oriented [ISO98b, LG87, MTG89, Mös93b, Mös93c, Mös95c, Mös98, Rad94, Tho92a, Tho92b, Bai88a, Bai88b, BG88, JCL85, Pou93, Pou95, Ste98, Wir90a, Ano86c]. Origin [Bri02a]. Original [BD01]. OS/2 [Rib90, Sch89b]. Osborne [Ano86b]. Other [GC00]. our [FvGGM90]. Output [BBG^{+}81f, Cra85, Lea87, Ode89]. Outreal [Wir65e]. Outward [Wal86]. Overlooked [Fra00a, Fra00b]. Overview [HS84]. Own [Ree00]. Oxford [LS06].

P [Ano88, Bow88, Sch89a, Wir75d, Alp79, Ber78, CRT80, GW77, GW78, Hur80, Jen73, Lou90, NAJN74, NAJN76, NAJN81]. P-Code [Sch89a, GW77, GW78, Hit79, Alp79, Lou90, NAJN81]. P-Compiler [Ber78]. P-machine [Jen73]. P4 [Rob82]. Package [LG87, Mar84]. Packages [BBG^{+}81j, Det93]. Page [Bis84]. Pages [Nou05, Ano79, Ano87c, Atk79a, Atk79b, Atk82b, Atk83, Bar77, Bis79a, Bis79b, Bis81, Bis82, Bis84, Bow88, Edw77, Fin77, Mee87, Ree84, Rob82, Ros74]. PAIL [GW78, GW77]. PAIL-3 [GW77]. PAIL-4 [GW78]. Paper [Wir65a, LD74, Woo94]. paperback [Ano87c, Atk82b, Bis82, Bis84, Bow88, Fin77, Wal89]. papers [BBZ01]. Paradigm [Wir96e]. Parallel [AP84, Bri96f, Bri96d, Bri02d, Bri94a, KS80, Wir66b]. parameter [Ran86, Ran87a, RWZ87]. parameter-dependent [Ran86, Ran87a, RWZ87]. parameter-free [Ran86, Ran87a, RWZ87]. Parameters [Sha86b]. parametric [HE82]. Parsers [Amm78a]. Parsing [Spe83]. Part [Fou91, Lüt21, Ran86, TWB11a, TWB11b, TWB11c, WW66a, ISO88, ISO96, ISO98a, ISO98b, WW66b, WW66c]. Partial [Fre95, Mey99]. partners [Szy00]. PASCAL [Amm75, JF87, Ran87a, AI83, Ano86b, Ano86c, Ano86d, Ano86a, Ano87a, Ano87b, Ano88, Hay80, ISO90a, JW85, Led84, Ten85, AC90, Amm75, Amm76, Amm77, Ano18, Bal76, Bar78, BD76, Col84, CGZ76, Eur85, Fei74, Geo74, GLN76, Haá82, HE82, HW72, Hoa74, JF87, JW74a, JW75, KS89, LP83, MS83, NW78, Nel79, NAJN74, NAJN76, Ran86, RWZ87, Ran87b, Shi92, Tsi82, WQ72, Wir70b, Wir70c, Wir71a, Wir72a, Wir72f, Wir72g, Wir73b, Wir75f, Wir75g, Wir75d, Wir75c, Wir76a, Wir76e, Wir80c, Ano80, Ame83, ABBH^{+}79, Alp79, ADDM84, Ano85, Pfi04, Ano06a, Ano06b, Ano21, AP84, Arm88, Atk79c, AN81, Atk82a, BS84, Ber78, Bis79c, Bis79d, Bow77, BFV84, Bri75a, Bri75b, Bri75c, Bri76a, Bri93, Bri96a, Bri96b, Bri96c, Bri96e, Bri01b, Bri02b]. Pascal [Bri02c, BWA82, BO83, Car85, Cha24, CC87, CD84, CK81, CGHP79, CH79, Com79, CW82, Com83, CL82, CMH85, CRT80, Des73, DS86, DSW82, EF88, Emr21, EMR91, Eva84, Fes24, FG82a, FG84a, FG84b, FM86, Fou75, Fre81, GAK86a, GAK86b, Ger82, GKLM79, Gru24, Hab73, Hab84, HL76, Har90, HM82, Hit79, HW73, HT86, Hur80, IAJ83, IEE83, ISO82, ISO88, ISO90b, ISO91, Jen73, JW74b, JW85, JW91, JCL85, JM84, Jos79, Jos80, Kae95, KE85, KP81, Ker81, Ker84, Ker82, KS84, KS80, Kru82, LF82, LD74, Lec74, LD75, LD84, LHN79, Lid76, Liu86, LS79, Mac79, MTT81, MTT83, MS90, Mar79, Mar75, Mar84, Mat84, MHT96, Mat80, McC83, MS84, Mey99, Moh77, Mul84, Nag76, NW84, NAJN81, ORT81, Old84, OW83]. Pascal [Par85, PD02, PV84, PD81, Pfi01, Pos24, Pro24, Qua24, Rav82, Ree91, RW92, RW94, Ric74, Rob83, RD85, RD86, RS76, Sal79b, Sal79c, Sal79a, Sal81, Sch80, Sch89a, Sha86b, SFIK80, Shr78, Shr79b, Shr79a, SMB83, SM79, SM81, SG82, SDK06, Ten78, Tes85, VN24, WC81, Wal86, WSH77, Wel78, WB79, WH86, WB85a, WB85b, WC83, Wic84, Wil80, Wir71d, Wir72e, Wir72b, Wir73a, Wir74c, Wir75a, Wir78b, Wir79a, Wir80a, Wir81c, Wir84b, Wir93a, Wir93b, Wir96b, Wir99a, Wir01c, Wir02e, Wir02c, Wir21, Yip84, You81, Ano79, Atk79a, Atk82b, Atk83, Bar77, Bis79a, Bis79b, Bis81, Bis84, Fin77, HW74, Ree84, Atk79b, Bis82, Rob82]. Pascal-Again [Sal79b]. PASCAL-Compiler [GLN76]. PASCAL-Compilers [Amm75].

Pascal-like [Arm88, Mey99, Old84]. Pascal-P [CRT80, Hur80]. Pascal-Plus [KS84, WB79]. PASCAL-program [KS89]. Pascal-Programmen [EMR91]. PASCAL-S [Wir75c, Wir80c, Wir81c]. Pascal-Version [Wir99a]. PASCAL-XSC [Shi92]. Pascal/Modula [Ano88]. Pascal/Modula-2 [Ano88]. Pascal/P [Hit79]. Pascal/P-Code [Hit79]. pass [Gut87]. Past [GALW12]. Patches [JM84, Lo85]. Path [DS86]. Pattern [Liu86]. Pay [Fra94a]. PC [Det93, Dis93, Sha86a, Ano87b]. PCFORT [CCN^{+}79]. PCODE [CNW80, CCN^{+}79]. PCs [Ano86a]. PDP [BD76, Fei74]. Pemberton [Rob82]. People [Bös05]. Perfect [BH86]. Performance [BA88, DSD97, Hur84, PV84, SIRB16]. Performed [Coh91, Wir91a, Wir91c]. Perlman [Ano79]. persistence [Kna97]. Personal [Bri96a, GW84, Wir79c, Wir79d, Wir81a, Wir00a, Wir15a, Bri93, Bri96b, Rei84, Wir81b]. Perspective [Wir10a, Wir15a, Wir74d]. Perspectives [BBZ01, BBP96, Wir91b, Wir00a]. Peter [Bis79b]. philosophy [KW85]. Phys [Ran87a]. PIC [Wir05c]. PICL [Wir05c, Wir07b]. pioneer [Ano24a, Daw24, LM24, PMR00, Pro24]. Pioneers [BD01, BD02, BD01]. Pitman [Atk79b, Atk83]. PL360 [Mal71, Mal72, Wir67c, Wir68b, Wir68c, WWS68]. place [Wir08k]. Planned [Wir72a]. PLDs [Wir96a]. Plea [Bös05, Wir95d]. Plotting [Ano87a]. Plus [KS84, WB79]. Pocket [Atk83, Atk83]. Point [Dot90, Dun84]. Pointer [LS79, Rav82, MHT96]. pointers [MHT96]. Pollution [DP85]. Polymorphic [Lea87]. polynomial [Arm88]. Portability [FK91, Lou90, NW78, PV84, Wir72d]. Portable [BHHM86, Cre90b, Cre91, Fra94b, Pow84a, Pow84b, Sch80, Sch89a, WF84, BB91, Pou95, Cre90a, Shi92]. Portal [Sch86, SH89]. Porting [DG96, Wir07e]. position [Sou90]. Possible [Rad94]. Poster [Baq99]. Postlude [GLN76]. POW [MLKK97c]. power [Tsi82]. pp [Abr74, Ano86b, Ano86c, Ano86d, Ano86a, Ano87a, Ano87b, Ano88]. practical [Bis79a]. Praxis [Gre86]. precedence [IH71, LL70, Wir65c]. precompiler [EF88]. Preemptive [CBG01]. Preferred [LS87]. Preliminary [HL76]. Prentice [Abr74, Edw77, Ros74, Wal89]. Prentice-Hall [Abr74, Edw77, Ros74, Wal89]. Preparation [GW84]. Preprocessor [Com79]. Press [Bis79a, Bis81, Bis84]. previews [Pou93]. PRICE [Atk83, Ano79, Ano87c, Atk79a, Atk79b, Atk82b, Bar77, Bis79a, Bis79b, Bis82, Bis84, Bow88, Mee87, Ree84, Rob82, Ros74, Wal89, Edw77]. primer [CGZ76, Fin77]. Principles [Wir69a]. Problem [Bow77, BFV84, Atk79a, Bai88a, Bai88b, Gol87, Par85]. problem-oriented [Bai88a, Bai88b]. Problems [Cor88, SG82, Ano79, JM84, KS89]. Procedural [Sha86b, Wir00a]. Procedure [Bri02a, Hei03, Lea87, Tsi82]. Procedures [Hei03, Sch89c, SM95, Wir75b, Old84]. Proceedings [Ano91, GLL74, USE85, BS03, Gia89, Sof84, BBP96, GW00, LS06, Mos97]. Process [BDW88, Hem88, LG87, PS96, Sew84b, Sew88]. Process-Oriented [LG87]. Processes [MS90]. Processing [Ano85, Bri01a, ISO88, IH71, Kru82, Wir66b]. Processor [BO83, Wir07d]. production [Pom93]. products [Pom93]. Prof [Ano24b]. Professor [Wir76c, Wir77a]. Program [BBG^{+}81e, BBG^{+}81k, Bri76a, Bri78a, Bri96e, Bri01b, Bri02c, Com79, Hur80, McC83, Wir66b, Wir71b, Wir71c, Wir80a, Wir80c, Wir83b, Wir01b, Wir02d, Bow88, BWA82, Kae95, KPUG93, KS89, Knu90, NW84, Ran86, Ran87a, RWZ87, Ree91]. Programmable [Wir08e]. Programmation [Sch86]. Programmen [EMR91]. Programmer [DNR90, Rei91].

Programmer-defined [DNR90]. Programmers [Kor24, MS84]. Programmieren [Wir75h, RW94, Wir72c, Wir85b]. Programmierens [Amm75]. Programmiersprache [Wir76e, Wir90c]. Programmiersprachen [JF87]. Programmierung [JF87, Mös93c, Mös98]. Programming [Ame83, Abr74, Amm74, Ano18, BST89, BdJ76, BdJ80, BS03, BBG^{+}81l, Bri77, Bri94b, Bri96c, Bri96d, Bri02a, Bri02b, Bru87a, Bru87b, DO85, EN74, ER92, Fes24, FG82a, FG82b, FG84a, FH98, Fra97, GALW12, Hab73, Har92, Har94, HH95a, HH95b, HW72, HW73, HW74, ISO82, ISO91, Jor90a, Jor90b, Ker81, Ker84, KW85, KS80, LD75, Lee85, LS06, Lin90, Mal71, Mal72, Mar79, MFMN89, Mös95a, MW91a, Mös93b, Mös93c, MW93, Mös95c, Mös98, MLKK97b, NPJ79, Nik97, Ode00, Pas85, Pet82, Por88b, Pow84c, RW92, SIRB16, SMB83, SM81, Tag88, Tem94, Wir65d, Wir66c, Wir67d, Wir68b, Wir68c, Wir70b, Wir70c, Wir71d, Wir72f, Wir72g, Wir72b, Wir72c, Wir73b, Wir73c, Wir75f, Wir75g, Wir76c, Wir77b, Wir77c, Wir77d, Wir77e, Wir77i, Wir79b, Wir82, Wir83c]. Programming [Wir84e, Wir85a, Wir85c, Wir87c, Wir87d, Wir87e, Wir87f, Wir88d, Wir88c, Wir90f, Wir98b, Wir00a, Wir01a, Wir01c, Wir02e, Wir03a, Wir04b, Wir05c, Wir07a, Wir08j, Wir15c, Wir16, Woo94, AI83, AC90, Amm75, Ano79, Ano86c, Atk82b, Atk83, Bai88a, Bai88b, BG88, BG96, Bri75a, Bri75b, Eur85, Emr21, Eva84, FG84b, GW00, Hab84, Hoa74, IAJ83, IEE83, JF87, KS89, Lal91, Lec73, LD74, Lec74, LD84, Mös91, Mos97, Nel91, Pir19, PS94, Rei84, Rib90, Ros74, SH89, Shi92, Sou90, SOM94, Wir67b, Wir72a, Wir74b, Wir74d, Wir75a, Wir75h, Wir76e, Wir84b, Wir87b, Wir90a, Wir90e, Wir90c, Wir99c, Wir00b, ISO90b, ISO96, ISO98a, ISO98b, RW94, Wir85b, Wir04b, Ano86b, Atk79b, Bis82, Bis79b, Bis81]. Programs [CC87, CH79, Fer95a, Fer95b, FW65a, FW65b, Mar75, RD85, RD86, SFIK80, SDK06, SM95, Wir74a, Wir76d, Wir79a, Wir80a, Wir98d, You81, ADDM84, Arm88, CL82, EMR91, LF82, Old84, Ran86, Ran87a, RWZ87, Ran87b, Ste98, Wir98c, Edw77]. Project [Wal86, WG92]. Projects [Fel86b, Fel86c, Wir00c]. PROLOG [JF87, JF87, Mul86]. proof [Knu90]. Properties [RD85, RD86]. property [Arm88]. Proposal [Gut97, Tho92a, Tho92b, Wir64a]. Proposed [BBG^{+}81l]. Pros [Bat91]. Protecting [Tor87b, Wie86]. Proving [RD85, RD86]. PRTDS [WB85a]. PSE [Tay90]. PSI [BBZ01]. PT [MS83]. Publication [Ano85, Bri96f, Bri02d, Bis79d]. Publications [Bow88]. Publishers [Fin77]. purpose [Mac79].

Quality [Dot90]. Quasi [KS80]. Quasi-parallel [KS80]. Queens [Gol87].

R [Ano86d, Ano86a, Bis79a, Fin77, Rei84]. R1990 [AC90]. R65 [Wir65b]. R65-24 [Wir65b]. Racehorses [Dot90]. Racing [Fel86b, Fel86c]. RADICAL [Ran87a, Ran86, Ran86]. radically [Pou93]. Range [Wel78]. rapid [Ano86a]. Raster [Wir90d]. Read [Wir75b]. Real [BDW88, Wir77b, Wir77c, Wir77d, Wir77e, Wir77i, Wir01a, Bud85, Dub85, Tem00, HL76]. Real-Time [BDW88, Wir77b, Wir77c, Wir77d, Wir77e, Wir77i, Wir01a, Bud85, Dub85, HL76]. Really [Gut97]. Realm [Nik97]. Recollections [Wir93a, Wir93b, Wir96b]. Recommended [SM79]. Record [LS79, Wir87a]. Recovery [Amm78a, Shr78, Shr79b, Shr79a]. Recursive [Amm78a]. Redefined [FK91]. redesignation [AC90]. Reference [BBG^{+}81l, Sit72, Wir80a, Fra90b]. references [MHT96]. Refinement [Wir71b, Wir71c, Wir83b, Wir01b, Wir02d]. Reflection [MS99a, MS99b]. regelbasierter [JF87]. relation [Dij00]. relations [IH71]. release [Gut90]. Remarks [Hay80, Ano80]. Remembering [Pos24, VN24]. Remote [Bri02a, Gut97]. Reply [LD74, Wir91a, Wir91c]. Report [Ano06b, JW74b, JW75, JW85, JW91, Tes85, Wir72g, Wir73b, JW74a, Wir76e]. Representation [Bis84, BBG^{+}81m]. Reprint [Wir83b, Wir02d, Wir02e]. Reproducible [Bri78b]. Reserved [Sal79c, ST86]. Response [Fel86a, Gol85, Wir78c]. Results [HH95a, HH95b]. retrospect [Knu00]. Revelation [Fre95]. Revidierter [Wir76e]. Review [Abr74, Ano79, Ano86b, Ano86c, Ano86d, Ano86a, Ano87a, Ano87b, Ano87c, Ano88, Atk79a, Atk79b, Atk82b, Atk83, Bar77, Bis79a, Bis79b, Bis81, Bis82, Bis84, Bow88, Edw77, Fin77, Mee87, Ree84, Rob82, Ros74, Wal89, Por88a, Ano87b]. reviewed [And84]. Revised [Mal71, Mal72, Wir72g, Wir73b, Wir18, BBZ01, Pir19, Wir76e, Wir08d]. Revision [AC90]. Revisited [Sav87, Wir78c]. Richard [Ano87c]. RIP [Fes24, Pro24]. RISC [Wir10c, Wir15b, Wir17a]. RISC-Architecture [Wir17a]. rivals [Szy00]. Road [Wir02a, Wir02b]. Robot [KS84]. Rohl [Bis81]. Rome [Gia89]. RPC [ER92]. rule [JF87]. rule-based [JF87]. Rules [BBG^{+}81q]. Run [Amm78a, WB85a, WB85b, LF82]. Run-Time [Amm78a, WB85a, WB85b, LF82]. Runtime [Som06]. Russia [BBZ01, BBP96].

S [Ano79, Ano86b, Ano86a, Ano87b, Ano87c, Bis81, Bis84, Rob82, Wir75c, Wir80c, Wir81c]. SA [Wir98b]. SAC-2 [Kre89]. Safe [Har92]. Safely [Hei03]. Salt [Sof84].

Salute [FvGGM90]. SAM2S [WF84]. Sather [SOM94]. Savage [Dun84]. saw [Lüt21]. Scalars [Atk79c]. scene [Woo94]. Scheduling [Hem88, Sch89c, Sew88]. Schneider [Ano79]. School [Wil80, BS76, Bös00]. Science [Bis79b, Gru24, MS87, Sev12, Wir95a, Wir02a, Wir02b, Wir10a, Daw24, PMR00, Wir91b]. Scientific [Bow88, Bri96f, Bri02d, Mös95a, Shi92]. Scope [And86]. Scratch [Wir89]. Script [Som06]. sd&m [BD01]. Seamless [Zel97]. Search [Bri96d]. Second [MS87, Ano91, BBP96]. Selected [Wir13]. Self [Gut97]. Self-Active [Gut97]. Semantically [BS84]. Semantically-based [BS84]. Semantics [Hof00, Zam96, BH95]. Semaphores [Bri02a, Wir78c]. Sensitive [Rob83]. sentence [MS83]. Separate [Fos86, Gut86]. September [Ano87a, Ano91, GLL74, GW00, LS06]. September/October [Ano87a]. Sequence [Sal79a]. Sequential [Shr78]. Serendipity [Jen00]. Series [Bal76, Bis79b, WQ72]. Services [Zel97]. Session [Baq99]. set [Coo89, Wir07f]. Seven [And84, JM84]. Several [BdJ76, BdJ80]. sharing [GKLM79]. Shell [DS84]. Shelley [Atk83]. shop [LP83]. Should [Fra94a]. Sicht [Wir74d]. SIGMA [Des73]. Signals [Wir78c]. Silver [KV15]. Simple [Fer95a, Fer95b, Hop80, Lo85, Rad94, Knu90, LP83]. Simplicity [Bri96d, Bös00, Geh00]. Simulation [KS80, LG87, MS90, Mar84, Haá82, RL01, Sal92]. Single [CRT80]. Single-accumulator [CRT80]. Slave [BK87]. slicing [Ste98]. Slim [FK97]. Small [BW71, GKLM79]. Smaller [Rad94]. SNOBOL4 [Liu86]. Software [Ano87c, Bri95a, Bri95b, BD01, BD02, DP85, Fra94b, Gut94, KP81, Kre89, Mar96, Mul86, ORT81, Pom85, Pom93, PS94, Pro24, Ros24, Sof84, Saw96, WS84, Wir72d, Wir81a, Wir94a, Wir95d, Wir08b, Wir08c, Sew87, VN24, Wir81b, Wir03b, Pom93, Wir94a, Ano86d]. Software-Explosion [Wir94a]. Softwareprodukten [Pom93]. Softwaretechnik [Pom85]. Solaris [War03]. Solo [Bri76a, Bri76b, Bri96e, Bri01b, Bri02c]. Solution [SG82]. Solutions [JM84, KS84]. Solving [Bow77, Gol87, Ano79, Atk79a, BFV84]. Some [Ano80, Gol85, HH95a, HH95b, HKMN93, Jos80, Liu86, Mof84, NPJ79, SS85]. Sorting [Wie84]. Sorts [Fel86b, Fel86c]. Source [Hur84]. Soviet [Wir67e]. SPARC [Tem90]. SPARC-Oberon [Tem90]. Spatial [Sch00a]. Special [Mac79]. Special-purpose [Mac79]. Specification [Eur85, EO80, HL76, ISO82, KP97b, KV15]. Specifications [BBG^{+}81m, Par85]. SPLIT [RWZ87]. sprachlichen [JF87]. Spreadsheet [Kom98]. Springer [Atk79a, Mee87]. Springer-Verlag [Mee87]. Stages [Wal86]. Stand [FF00]. Stand-Alone [FF00]. Standard [AI83, BBG^{+}81l, IEE83, JW91, PS96, SP96, IAJ83, Mar84, AC90, IAJ83, IEE83, JW85, Led84, WH86, Woo93]. Standardization [Kin89]. Standardized [PSSW97]. Standards [Ano85, Woo94, Ten85]. Stanford [Alp79]. State [Atk79c, Por88a, Wir08e, Atk82a]. statement [HM82]. Statements [BBG^{+}81n, Sal81, Atk82a]. States [BBG^{+}81l]. static [BWA82, LN98]. Statistical [Saw96]. Steps [RW92]. Stepwise [Wir71b, Wir71c, Wir83b, Wir01b, Wir02d, KM95]. still [Wir08a]. Storage [Amm78a]. strange [Gut00a]. Stress [Pro92]. String [Liu86, Wir64a]. Strings [Bis79c, Sal79b, Sal79a]. strive [Wir08k]. Structure [BBG^{+}81k, Hur80, You81]. Structured [Amm74, Lec74, Por88b, Wir74a, Amm75, Bri94a, Lec73, Sch00a]. Structured-Universe [Sch00a]. Structures [Edw77, Fel88, FM86, Sch00a, SW86, SW87, WS85, Wir75e, Wir76d, Wir83a, Wir84c, Wir84d, Wir86a, Wir86b, Wir99a, Wir04a, BWA82, Wir66b, Wal89]. Structuring [Wir79b]. Strukturierten [Amm75]. Student [SIRB16]. Students [Wir95a]. Studie [JF87]. Study [DS86, Gut00b, Hop80, Ree00, Ano86a, JF87, LF82]. Style [LHN79]. Styles [Wir08j]. Stylish [Mas98]. Stylistics [Sal79c]. subjective [Mat84]. Subprograms [BBG^{+}81o]. Subset [Wir75c, Wir81c, MS83]. subtle [JM84]. successor [RL01]. Successors [Wir02c]. Suitability [BK87, JF87]. Suitable [Mil89]. Suite [KC90, CD84]. Summary [Wir00c, Wir67e]. Summer [Sof84, BS76]. sums [RWZ87]. SUPERMAC [BO83]. SuperPascal [Bri94a, Bri94b, Bri95a, Bri95b, Bri96f, Bri02d]. support [BG88]. supporting [Lal91, Ter86]. Surveyor [FG82b, Pet82]. Swinging [KPUG93]. Switzerland [GLL74, GW00]. Symbol [Gut86]. Symbolic [Gia89, KE85]. symbols [IH71]. Symposium [EN74, Gia89, GLL74]. synchronization [BBS97]. SYNCRO [DS84]. Syntactic [HLW63, Mar14, Wil82, Wir77k]. SYNTAX [Mar14]. SYNTAX-Technology [Mar14]. Syntaxes [Woo86]. Synthesis [WC81]. System [Anoxxb, AN81, BHHM86, BBZ01, BBP96, Bri76a, Bri76b, Bri96e, Bri01b, Bri02c, CDJ^{+}89, CC87, DO85, DP85, FH98, Fra93, GW84, Gut90, Gut94, Hir83, LS94, Mar94, Pas85, Pou95, Pow84c, Rob83, Som06, Web67, WB85a, Wir67c, WWS68, Wir79b, WG8 , Wir89, WG89, WG92, Wir98b, Wir08h, Wir10b, WF84, Zel97, Ano86b, BCFT92, BCFT95, Haá82, KV15, Kru82, Mös94, NW84, Sew87, Shi92, Szy91, SOM94, Wir88a, Wir98a, Wir99b, Yip84, DG96, ISO88, Kna97, KP97a, Rei91, Wir74d, Wir95c]. System/360 [Web67]. Systematic [Wir73c, Abr74, Wir72c, Wir75h, Ros74]. Systematisches [Wir72c, Wir75h].

Systeme [War03, JF87].

Systemprogramming [Wir74d]. Systems [BST89, Bri01a, Bru86a, Bru86b, CBG01, ER92, Fel86b, Fel86c, Mar96, Nel91, PS94, Rov86, Wir73a, Wir01a, Ano87b, Bud85, GKLM79, ISO88, JF87, Nie88, Pou93, Sch00b].

Table [BMW91]. tables [Kae95]. Takara [KW85]. Taken [Wir02a, Wir02b]. Target [GC00, Wir17a, Zam96]. Tasks [BBG^{+}81p, Wir96e]. Taste [Woo93]. Teach [Moh89, MFMN89]. Teaching [Bös05, Bru86a, Bru86b, Sev12, Wir96a, Lec74]. Tech [Nou05]. Technical [Coh91, PHT91, Wir78c, Wir91c, Wir98a, Wir98b]. Techniken [Wir96d, Wir08f]. technique [LP83]. Techniques [Wir96d, Wir08f]. Technology [Ano91, Gut94, Mar14, Wir95b, AC90, ISO90b, ISO91, ISO96, ISO98a, ISO98b, Wir67e]. ten [Woo94]. Termination [CH79]. test [CD84, CG94]. Testing [Pro92, Bri78b]. Tests [Coh91, Wir91a, Wir91c]. Texas [USE85]. Text [Lo85, Szy92b, Gut85, Szy91]. Textbook [Wir95a]. tha [PS94]. Their [BD01]. them [JM84, Wir76c, Wir84e]. Themselves [Fra94a]. Theoretical [EN74]. theory [Aba94]. thought [Wir08k]. Thoughts [Wir94a]. Threads [Wir96e]. Threat [Kin93]. Three [KS84, MTT83, Wir15d, Ano87b]. Three-Dimensional [MTT83]. TI [Nou05]. TI-99 [Nou05]. TI-99/4A [Nou05]. Time [Amm78a, BDW88, Coh91, CBG01, Fer95a, Fer95b, Gri90, WB85a, Wir77b, Wir77c, Wir77d, Wir77e, Wir77i, Wir91c, Wir01a, Bud85, Dub85, GKLM79, LF82, WB85b, HL76, Wir91a]. time-sharing [GKLM79]. Titan [Gru24]. Tool [MFMN89, Mul86]. Tools [KP81, Pom85, Ree00, PS94]. Top [Fra93]. TopSpeed [Nan89]. tot [Qua24]. Tower [Fel86a]. Trace [Sch80]. transaction [Kru82]. Transfer [Wir95b]. transferability [Lec73]. transforming [LL70]. transitive [Dij00]. Translating [SM95, Wir98d, Wir14]. translation [Wir98c]. Translator [CCN^{+}79, CNW80, Fre81, Tem95, Ano88]. transportable [EF88]. Traps [Wir08g]. trauert [Qua24]. Tree [BMW91]. Trenches [Bie85, Gre86]. Tribute [VN24]. Trinity [Ano21]. Triptychon [Wir05d]. Troubleshooting [War03]. True [Sha87]. Tscharmer [Ano88]. Turbo [EMR91, EMR91, Ree91, Sha86b]. Turing [Ash87]. Tutorial [Gut84, Wir15c]. twenty [Ash87]. Two [WS85, Atk82a]. two-state [Atk82a]. Type [CDJ^{+}89, Coh91, CG94, Lea87, LS87, Nig93, Rad94, Ten78, Wir88e, BH95, IH71, KS89, Par85, SOM94, Wir07f, Wir91a, Wir91c]. type-ambiguity [Par85]. Type-Extension [Coh91, Wir91a, Wir91c]. typeflow [CG94]. Types [BBG^{+}81b, Bri75c, Cas91, CI85, EO80, Gol86, Hei03, KT90, MTT81, SW87, Wir87a, HE82].

U.S [Atk79a]. UCSD [Ano88, BFV84, CK81, PV84]. UFORT [CNW80]. UK [LS06, Ano91]. Undergraduate [GAK86a, GAK86b]. Unfair [Hem88]. Uninitialized [Bon87, GN87, Sav87, Tor87b, Wie86]. UNISEX [KE85]. United [BBG^{+}81l, Wal89]. Units [BBG^{+}81e]. Univ. [Ano24b]. Univac [Bal76]. Universal [Bri75c, CNW80]. Universe [Sch00a]. Universität [Wir74d]. University [Ano91, Bis81, Bis84, Wir74d]. UNIX [KE85, Pow84c]. UNIX-based [KE85]. Unnecessary [Wir77k]. Unwarranted [SS85]. Update [Car85]. Ural [Ano87b]. USA [Sof84, USE85]. Use [Gol86, Wil80, Wir5 b, Wir77j, Wir79d]. Useful [HKMN93]. USENIX [Sof84]. User [Ano86e, JW74b, JW75, JW85, JW91, Kna95, Rei91, Tem90, Tem95, JW74a, Kru82].

Users [Kna95]. Using [Bis84, Bow77, FH98, GLW94, HT86, KS84, Lea87, Li88, Moh77, Moh89, Pow84c, Reh87, Sch89c, SW86, Wec98, Amm75, Ano86c, Atk79a, BFV84, CG94, GKLM79, KV15, Mac79, Nie88]. Utah [Sof84]. Utility [BHHM86].

V4 [Anoxxb]. Validation

[KC90, WC83, Ree84]. value [Geh00]. Values [Hei03]. Variables [Rav82, Old84]. VAX [PV84]. Verification [LS79, Mar75]. Verilog [Wir14]. Verlag [Atk79a, Mee87]. Version [KP97a, Tem95, Wir99a, Wir04a, Wir99a]. Versus [Car85, Gut00a, HH95a, HH95b, Mat84, SIRB16, Wir96e]. via [Bis81]. Video [Pou90]. View [Bie85, Bös05, Gre86]. Viewpoint [Wir21]. Virtual [GC00, GM00]. virtuelle [War03]. Visibility [BBG^{+}81q]. Vision [Gut94]. Visionary [Ros24]. visit [Wir76c, Wir77a]. Visualizer [JPPM96a, JPPM96b, JPPM96c]. vital [VN24]. VMware [War03]. Vol [Ano86b, Ano86c, Ano86d, Ano86a, Ano87a, Ano87b, Ano88]. vom [Pom93]. Voyage [Rad93, Saw96]. vs [Bie85, Fel86a, Gre86].

W [Ano79, Atk79b, FvGGM90, BBG68, BBG^{+}69, Sit72]. waren [Wir08a]. warum [Wir08k]. Was [Wir08k]. Watt [Atk79b, Atk83]. way [Pom93]. WEB [Kre89, Som06]. WEB/Modula [Kre89]. WEB/Modula-2 [Kre89]. Weber [IH71, LL70]. Webster [Bar77]. Wege [Pom93]. Wegmann [Ano86c]. Weingart [Ano79]. Well [Wir74a, Dij00]. well-founded [Dij00]. Well-Structured [Wir74a]. Welt [Qua24]. were [Wir08a]. Werkzeuge [Pom85]. Wesley [Bis79b]. Who [Kor24]. whose [Knu90]. Wichmann [Ree84]. Wiener [Ano86a, Ano87c]. Wijngaarden [Wir65a]. Wiley [Ano79, Ano87c, Bis82, Ree84]. Wilson [Bis79a, Bis84]. Windows [Fel86b, Fel86c, War03, KPUG93, MLKK97b, Pou95]. Winter [USE85]. Winthrop [Fin77]. wir [Wir08k]. Wirth [Abr74, Edw77, Qua24, Ros74, Wal89, Ano18, Ano21, Ano24a, Ano24b, Ano24c, Bös00, Cha24, Cyt97, Daw24, Fes24, For88, GS24, Gru24, Hol21, Hop00, IH71, Kor24, LL70, LM24, Mar14, Mey24, Mös00, Pes97, PMR00, Por88b, Pos24, Pou91, Pou95, Pro24, Qua24, Ros24, Sev12, Spi24, TWB11a, TWB11b, TWB11c, VN24, Wir00c]. Wish [Zim85]. within [Wir64a]. without [Lin90, Old84]. Word [BH86, Wir68a]. Words [Pes97, Sal79c, ST86]. Working [MLKK97c]. workshop [Pou91]. Workstation [GW84, War03, HN91, JCL85]. workstations [Wir88a]. world [Emr21, Qua24, Sch00b, Tem00, Wag00]. Worst [CBG01]. Worst-Case [CBG01]. Write [Wir75b, Szy91, Szy92b]. Write-ing [Szy92b]. Writing [DS86, SFIK80]. Written [Hop80, Kru82].

X3.97 [AI83]. X3.97-1983 [AI83]. Xerox [Des73, Gut87]. xix [Mee87]. XNet [Lal91]. Xref [Wir80a]. XSC [Shi92]. xvi [Ano87c].

Year [Wir08l]. years [Ash87, Wir21, Woo94]. York [Ano79]. Young [CW82].

Z [Ree84]. Zimmerman [Fin77]. zur [EMR91, JF87, Pom93, Wir94a]. Zürich [GW00, Amm78b].

References

Abadi:1994:BMT

[Aba94] Martin Abadi. Baby Modula-3 and a theory of objects. Journal of Functional Programming, 4(2):249–283, April 1994. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (elec-

                                                   tronic). URL https://www.
                                                   cambridge.org/core/product/
                                                   7091DC3F4FC43DC753EE1D5372D8C38C.

Addyman:1979:DDP

[ABBH^{+}79] A. M. Addyman, R. Brewer, D. G. Burnett-Hall, R. M. de Morgan, W. Findlay, M. I. Jackson, D. A. Joslin, M. J. Rees, D. A. Watt, J. Welsh, and B. A. Wichmann. A draft description of Pascal. Software—Practice and Experience, 9(5):381–424, May 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Abrahams:1974:SPI

[Abr74]  Paul Abrahams.  Book re-

view: Niklaus Wirth, Systematic Programming: An Introduction, Prentice-Hall, Inc., Englewood Cliffs, N.J., 1973, 167 pp., 24 cm. Mathematics of Computation, 28(127):881–883, July 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

ANSI:1990:ANSe

[AC90] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard for information technology: programming language PASCAL: ANSI/ISO 7185-1990: Revision and redesignation of ANSI/IEEE 770.X3.97-1983 (R1990). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, revised edition, 1990. v + 78 (printed), vi + 92 (online) pp. Available in PostScript for personal use only at http://pascal.miningco.com/msub1.htm.

Ancona:1984:ILM

[ADDM84] M. Ancona, Leila De Floriani, G. Dodero, and S. Mancosu. Integrating library modules into Pascal programs. Software—Practice and Experience, 14(5):401–412, May 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

ANSI:1983:AIX

[AI83] American National Standards Institute and Institute of Electrical and Electronics Engineers. ANSI/IEEE 770 X3.97-1983: An American National Standard: IEEE Standard Pascal computer programming language. IEEE, New York, NY, USA, 1983. ISBN 0-471-88944-X. 128 pp. LCCN QA76.73.P2 A43 1983.

Auping:1990:MMM

[AJ90] Judy V. Auping and J. C. Johnston. Making the move to Modula-2. Dr. Dobb’s Journal of Software Tools, 15(9):16– ??, September 1990. CODEN DDJOEB. ISSN 1044-789X.

Angele:1992:MAC

[AK92] Jürgen Angele and Detlef Küpper. Modula-2 an alternative to C? ACM SIGPLAN Notices, 27(4):17–26, April 1992. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Alpert:1979:PPC

[Alp79] Donald Alpert. A Pascal Pcode interpreter for the Stanford Emmy. Technical Note 164, Stanford University, Department of Computer Science, Stanford, CA, USA, 1979. iii + 36 pp.

ANSI:pascal

[Ame83] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. The Pascal Programming Language. ANSI/IEEE 770X3.97-1983, 1983. See also [JW74a, JW74b, JW85].

Ammann:1974:MSP

[Amm74] Urs Ammann. Method of structured programming applied to the development of a compiler. In Günther et al. [GLL74], pages 93–99. ISBN 0-444-10596-4, 0-7204-2097-0. LCCN QA76 .I567 1973.

Ammann:1975:EPC

[Amm75] Urs Ammann. Die Entwicklung eines PASCAL-Compilers nach der Methode des Strukturierten Programmierens. (German) [The development of a PASCAL compiler using the structured programming method]. Ph.D. thesis, ETH Zürich, Zürich, Switzerland, 1975. 166 pp.

Ammann:1976:CGP

[Amm76] Urs Ammann. On code generation in a PASCAL compiler. Report 13, ETH Zürich, Zürich, Switzerland, 1976. i +

41 pp. URL http://pascal.
hansotten.com/uploads/wirth/
Pascal-On-Code-Generation-
in-a-Pascal-Compiler-eth-
3056-01.pdf.

Ammann:1977:CGP

[Amm77] Urs Ammann. On code generation in a PASCAL compiler. Software—Practice and Experience, 7(3):391–423, May/ June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ammann:1978:ERR

[Amm78a] Urs Ammann. Error recovery in recursive descent parsers: Runtime storage organization. Report 25, ETH Zürich, Zürich, Switzerland, May 1978. iii +

46 pp. URL http://pascal.
hansotten.com/uploads/wirth/
Error%20Recovery%20in%20Descent%
20Parsers%20Ammann%20025.
pdf.

Ammann:1978:ZI

[Amm78b] Urs Ammann. The Zurich implementation. In Barron [Bar78], pages 63–82. ISBN 0-471-27835-1 (hardcover). LCCN QA76.73.P2

P2.  URL http://pascal.
hansotten.com/uploads/wirth/
The%20Zurich%20Implementation%
20Amman.pdf.

Atkinson:1981:CCP

[AN81] L. V. Atkinson and S. D. North.

COPAS: a conversational Pascal system. Software—Practice and Experience, 11(8):819–829, August 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Andersen:1984:SMC

[And84] T. L. Andersen. Seven Modula compilers reviewed. Journal of Pascal, Ada and Modula-2, ?? (??):??, March/April 1984. CO-DEN JOPAD5. ISSN 0747-1351.

Anderson:1985:BMM

[And85] Brian R. Anderson. Bit manipulation in Modula-2. Dr. Dobb’s Journal of Software Tools, 10 (11):38–??, November 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Anderson:1986:SII

[And86] Terry L. Anderson. The scope of imported identifiers in Modula-2. ACM SIGPLAN Notices, 21 (9):17–21, September 1986. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anderson:1989:KMM

[And89] Brian R. Anderson. Kermit meets Modula-2. Dr. Dobb’s Journal of Software Tools, 14(5): 22–36, 83–93, May 1989. CO-DEN DDJOEB. ISSN 0888-3076.

Anonymous:1979:BRB

[Ano79] Anonymous. Book review: An introduction to programming and problems solving with Pascal, G. M. Schneider, S. W. Weingart and D. M. Perlman, Wiley, New York, 1978. No. of pages: 394.

Price: £9.20 (hardback). Software—Practice and Experience, 9(9):792–794, September 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1980:LES

[Ano80] Anonymous. Letter to the Editor: Some remarks on ‘A case for acquiring Pascal’. Software —Practice and Experience, 10 (3):247–248, March 1980. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See [Jos79].

Anonymous:1985:FIP

[Ano85] Anonymous. Federal Information Processing Standards Publication: Pascal. U.S. National Bureau of Standards, Gaithersburg, MD, USA, 1985. 8 pp.

URL https://nvlpubs.nist.
gov/nistpubs/Legacy/FIPS/
fipspub109-1985.pdf. FIPS

PUB 109.

Anonymous:1986:ARBf

[Ano86a] Anonymous. Article review: Interfacing assembly language to Modula-2 — a case study: creating rapid displays for PCs: Wiener, R. S. J. Pascal, Ada & Modula-2 Vol 5 No 4 (July/ August 1986) pp 21–26. Microprocessors and Microsystems, 10(8):452, October 1986. CO-DEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0141933186902383.

Anonymous:1986:ARBb

[Ano86b] Anonymous. Article review: Modula-2: an alternative to C for system programming: Djavaheri, M. and Osborne, S. J. Pascal, Ada, Modula-2 Vol 5 No 3 (May/June 1986) pp 47–52. Microprocessors and Microsystems, 10(7):399, September 1986. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0141933186903704.

Anonymous:1986:ARBc

[Ano86c] Anonymous. Article review: Object-oriented programming using Modula-2: Wegmann, A. J. Pascal, Ada, Modula-2 Vol 5 No 3 (May/June 1986) pp 5–17. Microprocessors and Microsystems, 10(7):399, September 1986. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0141933186903753.

Anonymous:1986:ARBd

[Ano86d] Anonymous. Article review: Software engineering in Modula-2: implementing the GPIB (IEEE 488) in a laboratory: Marshall, J. L. and Goldstein, R. D., J. Pascal, Ada, Modula-2 Vol 5 No 3 (May/ June 1986) pp 28–46. Microprocessors and Microsystems, 10 (7):399, September 1986. CO-DEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0141933186903728.

Anonymous:1986:MUM

[Ano86e] Anonymous. Modula-2/86 User’s Manual. LOGITECH, Inc., 806 Veterans Blvd., Redwood City, CA 94063, USA, third edition, 1986. xviii +

467 pp. URL http://www.
bitsavers.org/pdf/logitech/
modula-2/Modula-2_86_Ver_2.
00_Users_Manual_Apr86.pdf.

Anonymous:1987:ARBa

[Ano87a] Anonymous. Article review:

Plotting curves and data with microcomputers: Molau, G. E. J. Pascal, Ada, Modula-2 Vol 5 No 5 (September/October 1986) pp 5–45. Microprocessors and Microsystems, 11(2):112, March 1987. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0141933187902286.

Anonymous:1987:ARBc

[Ano87b] Anonymous. Article review: Review of three Modula-2 development systems for the IBM PC: Ural, S. J. Pascal, Ada and Modula-2 Vol 5 No 6 (November/December 1986) pp 48–55. Microprocessors and Microsystems, 11(3):168, April 1987. CO-DEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/014193318790161X.

Anonymous:1987:BRB

[Ano87c] Anonymous. Book review: Modula-2: A Software Development Approach, Gary A. Ford and Richard S. Wiener, Wiley, 1985. ISBN 0-471-87834-0. No. of pages: 400 + xvi. Price: £21.45 (paperback). Software —Practice and Experience, 17 (2):167, February 1987. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1988:ARB

[Ano88] Anonymous. Article review: A UCSD Pascal/Modula-2 translator: Levrat, B., Nerima, L. and Tscharmer, P., J. Pascal, Ada, Modula-2 Vol 6 No 6 (November/December 1987) pp 17–24. Microprocessors and Microsystems, 12(2):114, March 1988. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0141933188900610.

Anonymous:1991:PSI

[Ano91] Anonymous, editor. Proceedings of the Second International Modula-2 Conference, Modula-2 and Beyond: 11–13 September 1991, Loughborough University of Technology, UK. Centre for Extension Studies, Loughborough University of Technology, Loughborough, Leicestershire, 1991. LCCN QA76.73.M63 I57 1991.

Anonymous:1994:IM

[Ano94] Anonymous. Introducing Modula-3. Linux Journal, 7:??, November 1994. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:2006:CP

[Ano06a] Anonymous. Component Pascal. Web site., August 24, 2006.

URL https://en.wikipedia.
org/wiki/Component_Pascal.

This site has undergone numerous updates from 2006 to 2024.

Anonymous:2006:CPL

[Ano06b] Anonymous. Component Pascal language report. Report, Oberon Microsystems, Inc., Zürich, Switzerland, October 2006. 32 pp. URL https:

//www.astrobe.com/CPIde/CP-
Lang.pdf.

Anonymous:2018:NWP

[Ano18] Anonymous. Niklaus Wirth and PASCAL — programming language for the masses. Web site, February 15, 2018. URL http:

//scihi.org/nikolaus-wirth-
pascal/.

Anonymous:2021:WTP

[Ano21] Anonymous. The Wirth trinity – Pascal. Web site, April 28, 2021.

URL https://craftofcoding.
wordpress.com/2021/04/28/
the-wirth-trinity-pascal/.

Anonymous:2024:CPN

[Ano24a] Anonymous. Computer pioneer Niklaus Wirth has died. Web site, January 4, 2024. URL

https://www.publicnow.com/
view/7A7BEAF82FAC7EB495CD524FBF368290C34A0F11?
1704415819.

Anonymous:2024:JML

[Ano24b] Anonymous. The JKU mourns the loss of Univ. Prof. Dr. Niklaus Wirth. Johannes Kepler University Linz Web site, January 2024. URL

https://www.jku.at/en/2024/
obituary-for-univ-prof-niklaus-
wirth/.

Anonymous:2024:NWO

[Ano24c] Anonymous. Niklaus Wirth obituary. Rebellion Research Web site, January 4, 2024. URL

https://www.rebellionresearch.
com/niklaus-wirth-obituary.

Anonymous:20xx:MC

[Anoxxa] Anonymous. Mapping to C. Web site, 20xx. URL

https://freepages.modula2.
org/xdsdoc/xm011.html.

Anonymous:20xx:OSV

[Anoxxb] Anonymous. The Oberon System (V4). Web site,

20xx.  URL https://ssw.
jku.at/Research/Projects/
Oberon.html.

Ardo:1984:IPB

[AP84] Anders Ardö and Lars Philipson. Implementation of a Pascal based parallel language for a multiprocessor computer. Software—Practice and Experience, 14(7):643–657, July 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Armbruster:1988:PDM

[Arm88] Dieter Armbruster. A polynomial determination of the mostrecent property in Pascal-like programs. Theoretical Computer Science, 56(1):3–15, 1988. CO-DEN TCSCDI. ISSN 0304-3975,1879-2294.

Ashenhurst:1987:ATA

[Ash87] Robert L. Ashenhurst, editor.

ACM Turing Award Lectures: the first twenty years, 1966– 1985. ACM Press anthology series. ACM Press and Addison-Wesley, New York, NY 10036, USA and Reading, MA, USA, 1987. ISBN 0-201-07794-9, 0-89791-183-0. xviii + 483 pp. LCCN QA76.24 .A33 1987.

Atkinson:1979:BRBa

[Atk79a]  L. V. Atkinson.  Book re-

view: Microcomputer problem solving using Pascal, Kenneth L. Bowles, Springer Verlag, Berlin, 1977. No. of pages: 563. Price: DM21.40, U.S. $9.80. Software—Practice and Experience, 9(3):250–252, March 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkinson:1979:BRBb

[Atk79b] L. V. Atkinson. Book review:

Pascal: An introduction to methodical programming, W. Findally and D. W. Watt, Pitman, 1978. No. of pages: 306. Price: £4.95. Software—Practice and Experience, 9(6):511–512, June 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkinson:1979:PSS

[Atk79c] L. V. Atkinson. Pascal scalars as state indicators. Software —Practice and Experience, 9 (6):427–431, June 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkinson:1982:OTS

[Atk82a] L. V. Atkinson. Optimizing twostate case statements in Pascal. Software—Practice and Experience, 12(6):571–581, June 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkinson:1982:BRB

[Atk82b] Laurence V. Atkinson. Book review: Foundations of programming with Pascal, Lawrie Moore, Ellis Horwood. No. of pages:

238. Price: £14.50 (Hardback) £5.00 (paperback). Software— Practice and Experience, 12(1): 101–102, January 1982. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Atkinson:1983:BRB

[Atk83] L. V. Atkinson. Book review: Pocket guide to Pascal, David Watt, Pitman. No. of pages: 64. PRICE: £2.25. Pocket guide to programming, John Shelley, Pitman. No. of pages: 64. Price: £2.25. Software—Practice and Experience, 13(10):981, October 1983. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

BrinchHansen:1988:JPM

[BA88] Per Brinch Hansen and Rangachari Anand. Joyce performance on a multiprocessor. Electrical Engineering and Computer Science Technical Report 28, School of Computer and Information Science, Syracuse University, Syracuse, NY, USA, September 1988. ii + 19 pp.

URL https://surface.syr.
edu/eecs_techreports/28.

Bailes:1988:API

[Bai88a] P. A. Bailes. An advanced problem-oriented introduction to imperative programming with Modula-2. The Computer Journal, 31(6):531–539, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.

oup.co.uk/computer_journal/
hdb/Volume_31/Issue_06/tiff/
531.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_31/Issue_06/tiff/532.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/533.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/534.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/535.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/536.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_31/Issue_
06/tiff/537.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/538.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/539.tif.

Bailes:1988:APO

[Bai88b] P. A. Bailes. An advanced problem-oriented introduction to imperative programming with Modula-2. The Computer Journal, 31(6):531–539, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.

oup.co.uk/computer_journal/
hdb/Volume_31/Issue_06/tiff/
531.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_31/Issue_06/tiff/532.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/533.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/534.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/535.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/536.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/537.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/538.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_
06/tiff/539.tif.

Ball:1976:PIP

[Bal76] M. S. Ball. PASCAL 1100: an implementation of the PASCAL language for Univac 1100 series computers. Technical Report 271, Naval Ocean Systems Center, San Diego, California 92152, USA, July 1, 1976. 24 pp.

URL https://apps.dtic.mil/
sti/tr/pdf/ADA059861.pdf.

Baquero:1999:PS

[Baq99] Carlos Baquero. Poster session.

In Object-Oriented Technology ECOOP’99 Workshop Reader, pages 371–383. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-46589-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/3-540-46589-8_22.

Barron:1977:BRBd

[Bar77] D. W. Barron. Book review: Introduction to Pascal, C. A. G. Webster, Heyden, London, 1976. No. of pages: 129. Price: £5-50. Software—Practice and Experience, 7(4):539–541, July 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barron:1978:PLI

[Bar78] D. W. (David William) Barron, editor. PASCAL: the language and its implementation. Wiley series in computing. Wiley, New York, NY, USA, 1978. ISBN 0-471-27835-1 (hardcover). ix + 301 pp. LCCN QA76.73.P2 P2.

Bates:1991:MP

[Bat91] Rodney M. Bates. Modula-2: The pros. C Users Journal, 9 (1):137–??, January 1991. ISSN 0898-9788.

Bowen:1991:HMC

[BB91] D. E. Bowen and A. C. Beers. A highly-portable Modula-2 compiler. In Anonymous [Ano91], pages 68–76. LCCN QA76.73.M63 I57 1991.

Bauer:1968:AWI

[BBG68] Henry R. Bauer, Sheldon Becker, and Susan L. Graham. ALGOL W implementation. Report CS-98, Stanford University, Department of Computer Science, Stanford, CA, USA, May 20, 1968. vi + 147 pp. URL http://

                                        infolab.stanford.edu/TR/CS-
                                        TR-68-98.html. From the in-

troduction: “The project was initiated and directed by Professor Niklaus Wirth, who proposed many of the ideas incorporated in the compiler and suggested ways to bring them about. Joseph W. Wells, Jr. and Edwin H. Satterthwaite, Jr. wrote the PL360 System in which the compiler is embedded, the linkages to the compiler, and the loader.”.

Bauer:1969:AWI

[BBG^{+}69] Henry Bauer, Sheldon Becker, Susan L. Graham, Edwin Satterthwaite, George E. Forsythe, and Robert W. Floyd. Algol W language description. Report CS 110, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1969. ii + 103 pp. URL https://

bitsavers.org/pdf/stanford/
cs_techReports/CS110_ALGOL_
W_Revised_Sep69.pdf.

Brauer:1981:BM

[BBG^{+}81a] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Back matter. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 207–243. ISBN 3-540-10693-6. LCCN QA76.73.A35 P76.

URL https://link.springer.
com/content/pdf/bbm:978-3-
540-38620-9/1.

Brauer:1981:DT

[BBG^{+}81b] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Declarations and types. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 15–45. ISBN 3-540-10693-6. LCCN QA76.73.A35 P76.

URL https://link.springer.
com/chapter/10.1007/3-540-
10693-6_3.

Brauer:1981:E

[BBG^{+}81c] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Exceptions. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 143–153. ISBN 3-540-10693-6. LCCN QA76.73.A35 P76.

URL https://link.springer.
com/chapter/10.1007/3-540-
10693-6_11.

Brauer:1981:FM

[BBG^{+}81d] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Front matter. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages i–x. ISBN 3-540-10693-6. LCCN QA76.73.A35 P76.

URL https://link.springer.
com/content/pdf/bfm:978-3-
540-38620-9/1.

Brauer:1981:GPU

[BBG^{+}81e] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Generic program units. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 155–167. ISBN 3-540-10693-6. LCCN QA76.73.A35 P76.

URL https://link.springer.
com/chapter/10.1007/3-540-
10693-6_12.

Brauer:1981:IO

[BBG^{+}81f] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Inputoutput. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 183–206. ISBN 3-540-10693-6. LCCN QA76.73.A35 P76.

URL https://link.springer.
com/chapter/10.1007/3-540-
10693-6_14.

Brauer:1981:I

[BBG^{+}81g] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Introduction. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 1–6. ISBN 3-540-10693-6. LCCN QA76.73.A35 P76.

URL https://link.springer.
com/chapter/10.1007/3-540-
10693-6_1.

Brauer:1981:LE

[BBG^{+}81h] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Lexical elements. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 7–13. ISBN 3-540-10693-6. LCCN QA76.73.A35 P76.

                                        URL https://link.springer.
                                        com/chapter/10.1007/3-540-
                                        10693-6_2.

Brauer:1981:NE

[BBG^{+}81i] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Names and expressions. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 47–71. ISBN 3-540-10693-6. LCCN QA76.73.A35 P76.

                                        URL https://link.springer.
                                        com/chapter/10.1007/3-540-
                                        10693-6_4.

Brauer:1981:P

[BBG^{+}81j] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Packages. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 93–103. ISBN 3-540-10693-6. LCCN QA76.73.A35 P76.

                                        URL https://link.springer.
                                        com/chapter/10.1007/3-540-
                                        10693-6_7.

Brauer:1981:PSC

[BBG^{+}81k] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Program structure and compilation issues. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 131–141. ISBN 3-540-10693-6. LCCN QA76.73.A35 P76.

URL https://link.springer.
com/chapter/10.1007/3-540-
10693-6_10.

Brauer:1981:PLA

[BBG^{+}81l] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth, editors. The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense, volume 106 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. ISBN 3-540-10693-6. LCCN QA76.73.A35 P76.

Brauer:1981:RSI

[BBG^{+}81m] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Representation specifications and implementation dependent features. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 169–181. ISBN 3-540-10693-6. LCCN QA76.73.A35 P76.

URL https://link.springer.
com/chapter/10.1007/3-540-
10693-6_13.

Brauer:1981:Sa

[BBG^{+}81n] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Statements. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 73–82. ISBN 3-540-10693-6. LCCN QA76.73.A35 P76.

                                        URL https://link.springer.
                                        com/chapter/10.1007/3-540-
                                        10693-6_5.

Brauer:1981:Sb

[BBG^{+}81o] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Subprograms. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 83–92. ISBN 3-540-10693-6. LCCN QA76.73.A35 P76.

                                        URL https://link.springer.
                                        com/chapter/10.1007/3-540-
                                        10693-6_6.

Brauer:1981:T

[BBG^{+}81p] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Tasks. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 115–130. ISBN 3-540-10693-6. LCCN QA76.73.A35 P76.

URL https://link.springer.
com/chapter/10.1007/3-540-
10693-6_9.

Brauer:1981:VR

[BBG^{+}81q] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Visibility rules. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 105–114. ISBN 3-540-10693-6. LCCN QA76.73.A35 P76.

URL https://link.springer.
com/chapter/10.1007/3-540-
10693-6_8.

Bjorner:1996:PSI

[BBP96] Dines Bjørner, Manfred Broy, and Igor V. Pottosin, editors. Perspectives of System Informatics: Second International Andrei Ershov Memorial Conference, Akademgorodok, Novosibirsk, Russia, June 25–28, 1996, Proceedings, volume 1181 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CO-DEN LNCSD9. ISBN 3-540-62064-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I5443 1996.

Back:1997:ABC

[BBS97] Ralph Back, Martin Büchi, and Emil Sekerinski. Action-based concurrency and synchronization for objects. In Transformation-Based Reactive Systems Development, pages 248–262. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-69058-1. ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        URL http://link.springer.
                                        com/chapter/10.1007/3-540-
                                        63010-4_17.

Bjarner:2001:PSI

[BBZ01] Dines Bjärner, Manfred Broy, and Alexandre V. Zamulin, editors. Perspectives of System Informatics: Fourth International Andrei Ershov memorial Conference, PSI 2001, Akademgorodok, Novosibirsk, Russia, July 2001: revised papers, volume 2244 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-43075-X (softcover), 3-540-45575-2 (e-book). LCCN QA75.5 .I52 2001.

Brandis:1992:OSF

[BCFT92] Marc Brandis, Régis Crelier, Michael Franz, and Josef Templ. The Oberon system family. Technical report 174, ETH Zürich, Zürich, Switzerland, April 1992. ii + 32 pp.

                                        URL http://hdl.handle.net/
                                        20.500.11850/68908.

Brandis:1995:OSF

[BCFT95] Marc M. Brandis, Régis Crelier, Michael Franz, and Josef Templ. The Oberon system family. Software—Practice and Experience, 25(12):1331–1366, December 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bron:1976:PCP

[BD76] C. Bron and W. De Vries. A PASCAL compiler for PDP 11 minicomputers. Software—Practice and Experience, 6(1):109– 116, January/March 1976. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Broy:2001:PTC

[BD01] Manfred Broy and Ernst Denert, editors. Pioneers and Their Contributions to Software Engineering: sd&m Conference on Software Pioneers, Bonn, June 28/29, 2001, Original Historic Contributions. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN QA76.758.

Broy:2002:SPC

[BD02] Manfred Broy and Ernst Denert, editors. Software Pioneers: Contributions to Software Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-43081-4, 3-642-59412-3 (ebook), 3-642-59413-1, 3-642-63970-4 (print). 728 pp. LCCN

QA76.758. URL http://link.
springer.com/10.1007/978-3-
642-59412-0.

Boom:1976:CCS

[BdJ76] H. J. Boom and E. de Jong. A critical comparison of several implementations of programming languages. Report, Mathematisch Centrum, Amsterdam, The Netherlands, 1976.

Boom:1980:CCS

[BdJ80] H. J. Boom and E. de Jong. A critical comparison of several programming language implementations. Software—Practice and Experience, 10(6):435–473, June 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Burns:1988:MIR

[BDW88] A. Burns, G. L. Davies, and A. J. Wellings. A Modula-2 implementation of a real-time process abstraction. ACM SIGPLAN Notices, 23(10):49–58, October 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berry:1978:EPP

[Ber78] R. E. Berry. Experience with the Pascal P-compiler. Software —Practice and Experience, 8 (5):617–627, September/October 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barnard:1978:EM

[BET78] David T. Barnard, W. David Elliott, and David H. Thompson. Euclid and Modula. ACM SIGPLAN Notices, 13(3):70–84, March 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bowles:1984:PSU

[BFV84] Kenneth L. Bowles, Stephen D.

Franklin, and Dennis J. Volper. Problem solving using UCSD Pascal. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1984. ISBN 0-387-90822-6 (New York), 3-540-90822-6 (Berlin). xi + 340 + 106 pp. LCCN QA76.73.U25 B69 1984.

Bergin:1988:WDM

[BG88] Joseph Bergin and Stuart Greenfield. What does Modula-2 need to fully support object oriented programming? ACM SIGPLAN Notices, 23(3):73–82, March 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergin:1996:HPL

[BG96] T. J. Bergin, Jr. and R. G. Gibson, Jr., editors. History of programming languages: II. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-201-89502-1. xvi + 864 pp. LCCN QA76.7 .H57 1996.

Bruckner:1986:MPH

[BH86] Jared A. Bruckner and James Harp. Minimal perfect hashing functions for Modula-2 word lists. Journal of Pascal, Ada and Modula-2, 5(6):39–40, November–December 1986. CO-DEN JPAME8. ISSN 0747-1351, 0735-1232.

Bancroft:1995:FSL

[BH95] Peter Bancroft and Ian Hayes. A formal semantics for a language with type extension. In ZUM ’95: The Z Formal Specification Notation, pages 298–314. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-44782-2. ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        URL http://link.springer.
                                        com/chapter/10.1007/3-540-
                                        60271-2_127.

Biagioni:1986:POS

[BHHM86] E. Biagioni, K. Hinrichs, G. Heiser, and C. Muller. Portable operating system interface and utility library for Modula-2. IEEE Software, 3(6):18–26, November 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bielak:1985:AVM

[Bie85] Richard Bielak. ADA(*) vs. Modula-2: A view from the trenches. ACM SIGPLAN Notices, 20(12):13–17, December 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See comments [Gre86].

Bishop:1979:BRBa

[Bis79a] J. M. Bishop. Book review: A practical introduction to Pascal, I. R. Wilson and A. M. Addyman, Macmillan Press, 1978. No. of pages: 148. Price: £3.50. Software—Practice and Experience, 9(3):252–253, March 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:1979:BRBb

[Bis79b] Judy M. Bishop. Book review:

Programming in Pascal, Peter Grogono, Addison-Wesley Series in Computer Science, 1978. No. of pages: 359. Price: £7.50. Software—Practice and Experience, 9(5):425–426, May 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:1979:ISP

[Bis79c] Judy M. Bishop. Implementing strings in Pascal. Software —Practice and Experience, 9(9): 779–788, September 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:1979:PP

[Bis79d] Judy M. Bishop. On Publication Pascal. Software—Practice and Experience, 9(9):711– 717, September 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:1981:BRBb

[Bis81] Judy M. Bishop. Book review:

Programming via Pascal, J. S. Rohl and H. J. Barrett, Cambridge University Press, 1980. No. of pages: 327, E12.50 (cloth). Software—Practice and Experience, 11(8):880–881, August 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:1982:BRB

[Bis82] J. M. Bishop. Book review: Pascal programming, Laurence Atkinson, Wiley, Chichester, 1980. No. of pages: 428. Price £16.50 (Hardback); £6.95 (paperback). Software—Practice and Experience, 12(1):99, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:1984:BRI

[Bis84] J. M. Bishop. Book review: Information Representation and Manipulation Using Pascal, E. S. Page and L. B. Wilson. Cambridge University Press, 1983. No. of pages: 272. Price: hardback, £15.00; paperback, £6.95. Software—Practice and Experience, 14(2):195, February 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bisignani:1991:MM

[Bis91] Michael Bisignani. Multitasking in Modula-2. Computer Language Magazine, 8(2):53–??, February 1991. CODEN COM-LEF. ISSN 0749-2839.

Beidler:1986:CGM

[BJ86] J. Beidler and P. Jackowitz. Consistent generics in Modula-2. ACM SIGPLAN Notices, 21(4):32–41, April 1986. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Butler:1987:SMS

[BK87] Gregory Butler and Matthew J.

Kendall. The suitability for master/slave concurrency of Concurrent Euclid, Ada and Modula. Software—Practice and Experience, 17(2):117–134, February 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Blaschek:1989:IOM

[Bla89] Günther Blaschek. Implementation of objects in Modula-2. Structured Programming, 10 (3):147–155, 1989. CODEN STPGEM. ISSN 0935-1183.

Borstler:1991:TCT

[BMW91] Jürgen Börstler, Ulrich Möncke, and Reinhard Wilhelm. Table compression for tree automata. ACM Transactions on Programming Languages and Systems, 13 (3):295–314, July 1991. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org/pubs/toc/Abstracts/
0164-0925/117013.html.

Brown:1983:SMP

[BO83] P. J. Brown and J. A. Ogden.

The SUPERMAC macro processor in Pascal. Software—Practice and Experience, 13(4):295–304, April 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Burton:2000:CCM

[BOB^{+}00] Robert Burton, Farrell Ostler, Thorn Boyer, Fon Brown, and Matt Morrise. The chip company that made $100M with MODULA-2. In Böszörményi [Bös00], pages 219–226. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1 (dpunktVerl.). LCCN QA76.756 .S36 2000. URL http:/

/digitale-objekte.hbz-nrw.
de/storage2/2016/09/08/file_
8/6866144.pdf.

Bondy:1987:UMA

[Bon87] J. Bondy. Uninitialized Modula-2 abstract object instances, yet again. ACM SIGPLAN Notices, 22(5):58–63, May 1987. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boszormenyi:1992:CMO

[Bös92] Laszlo Böszörmenyi. A comparison of Modula-3 and Oberon-2. In Shifting Paradigms in Software Engineering, pages 126– 137. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-7091-9258-7.

URL http://link.springer.
com/chapter/10.1007/978-3-
7091-9258-0_13.

Boszormenyi:1993:CMO

[Bös93] László Böszörményi. A comparison of Modula-3 and Oberon-2. Structured Programming, 14(1): 15–22, 1993. CODEN STPGEM. ISSN 0935-1183.

Boszormenyi:1998:WJM

[Bös98] László Böszörményi. Why Java is not my favorite first-course language. Software — Concepts and Tools, 19(3):141–145, March 1998. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Boszormenyi:2000:SNW

[Bös00] László Böszörményi, editor. The school of Niklaus Wirth: the art of simplicity. dpunkt-Verlag, Heidelberg, Germany, 2000. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1 (dpunktVerl.). viii + 260 pp. LCCN QA76.756 .S36 2000. URL http:/

/digitale-objekte.hbz-nrw.
de/storage2/2016/09/08/file_
8/6866144.pdf.

Boszormenyi:2005:TPP

[Bös05] Laszlo Böszörmenyi. Teaching:

People to people — about people a plea for the historic and human view. In From Computer Literacy to Informatics Fundamentals, pages 93–103. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-31958-1. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.
com/chapter/10.1007/978-3-
540-31958-0_13.

Bowles:1977:MPS

[Bow77] Kenneth L. Bowles. Microcomputer Problem Solving Using Pascal. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. ISBN 3-540-90286-4, 1-4615-9998-9 (e-book), 3-662-38578-3 (e-book). 563 pp. LCCN QA76.73.P2 .B68. URL http:/

/link.springer.com/10.1007/
978-3-662-38578-4.

Bowen:1988:BRB

[Bow88] J. P. Bowen. Book review: Modula-2: Constructive program development, P. Messer and I. Marshall, Blackwell Scientific Publications, 1986. ISBN 0-632-01609-4. No. of pages: 544. Price: £35.00 (paperback £14.95). Software—Practice and Experience, 18(2):167–168, February 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

BrinchHansen:1975:PLCa

[Bri75a] Per Brinch Hansen. The programming language Concurrent Pascal. IEEE Transactions on Software Engineering, SE-1(2):199–207, June 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/brinch-hansen.net/papers/
1975a.pdf;   http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=6312840.

BrinchHansen:1975:PLCb

[Bri75b] Per Brinch Hansen. The programming language Concurrent Pascal. In Friedrich L. Bauer and Klaus Samelson, editors, Language Hierarchies and Interfaces, International Summer School, Marktoberdorf, Germany, July 23–August 2, 1975, volume 46 of Lecture Notes in Computer Science, pages 82–110. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975.

BrinchHansen:1975:UTC

[Bri75c] Per Brinch Hansen. Universal types in Concurrent Pascal. Information Processing Letters, 3 (6):165–166, July ??, 1975. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

BrinchHansen:1976:SOSa

[Bri76a] Per Brinch Hansen. The Solo operating system: a Concurrent Pascal program. Software —Practice and Experience, 6(2): 141–149, April/June 1976. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

URL http://brinch-hansen.
net/papers/1976b.pdf.

BrinchHansen:1976:SOSb

[Bri76b] Per Brinch Hansen. The Solo operating system: Job interface. Software—Practice and Experience, 6(2):151–164, April/ June 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

BrinchHansen:1977:EMC

[Bri77] Per Brinch Hansen. Experience with modular concurrent programming. IEEE Transactions on Software Engineering, SE-3 (2):156–159, March 1977. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://brinch-hansen.
net/papers/1977a.pdf.

BrinchHansen:1978:NMP

[Bri78a]  Per Brinch Hansen.  Net-

work: a multiprocessor program. IEEE Transactions on Software Engineering, SE-4(3):194–199, May 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//brinch-hansen.net/papers/
1977d.pdf.

BrinchHansen:1978:RTM

[Bri78b] Per Brinch Hansen. Reproducible testing of monitors. Software—Practice and Experience, 8(6):721–729, November/December 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

URL http://brinch-hansen.
net/papers/1978b.pdf.

BrinchHansen:1988:MIJ

[Bri88] Per Brinch Hansen. A multiprocessor implementation of Joyce. Electrical Engineering and Computer Science Technical Report 29, School of Computer and Information Science, Syracuse University, Syracuse, NY, USA, September 1988. ii + 19 pp.

URL https://surface.syr.
edu/eecs_techreports/29.

BrinchHansen:1993:MCP

[Bri93]  Per Brinch Hansen.  Moni-

tors and Concurrent Pascal: a personal history. ACM SIGPLAN Notices, 28(3):1–35, March 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/brinch-hansen.net/papers/
1993a.pdfa;   http://www.
acm.org:80/pubs/citations/
proceedings/plan/154766/p1-
hansen/.

BrinchHansen:1994:ICS

[Bri94a]  Per Brinch Hansen.  Inter-

ference control in SuperPascal — a block-structured parallel language. The Computer Journal, 37(5):399–406, ???? 1994. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://brinch-hansen.net/
papers/1994d.pdf;   http:/
/comjnl.oxfordjournals.org/
content/37/5/399.full.pdf+
html;   http://www3.oup.co.
uk/computer_journal/Volume_
37/Issue_05/Vol37_05.body.
html#AbstractHansen.

BrinchHansen:1994:PLS

[Bri94b] Per Brinch Hansen. The programming language SuperPascal. Software—Practice and Experience, 24(5):467–483, May 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

BrinchHansen:1995:SSa

[Bri95a] Per Brinch Hansen. The Super-Pascal software. ACM SIGNUM Newsletter, 30(3):27, July 1995.

CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

BrinchHansen:1995:SSb

[Bri95b] Per Brinch Hansen. The Super-Pascal software. ACM SIGSOFT Software Engineering Notes, 20 (5):91, December 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

BrinchHansen:1996:MCPa

[Bri96a] Per Brinch Hansen. Monitors and Concurrent Pascal: a personal history. In The Search for Simplicity: Essays in Parallel Programming [Bri96d], chapter 21, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996. BrinchHansen:1996:MCPb

[Bri96b] Per Brinch Hansen. Monitors and Concurrent Pascal: a personal history. In Bergin, Jr. and Gibson, Jr. [BG96], pages 121– 172. ISBN 0-201-89502-1. LCCN QA76.7 .H57 1996.

BrinchHansen:1996:PLC

[Bri96c] Per Brinch Hansen. Programming language Concurrent Pascal. In The Search for Simplicity: Essays in Parallel Programming [Bri96d], chapter 7, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:SSE

[Bri96d] Per Brinch Hansen. The Search for Simplicity: Essays in Parallel Programming. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7566-7. xiv + 528 pp. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:SOSa

[Bri96e] Per Brinch Hansen. Solo operating system: a Concurrent Pascal program. In The Search for Simplicity: Essays in Parallel Programming [Bri96d], chapter 8, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:SPL

[Bri96f] Per Brinch Hansen. SuperPascal:

a publication language for parallel scientific computing. In The Search for Simplicity: Essays in Parallel Programming [Bri96d], chapter 24, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:2001:COS

[Bri01a] Per Brinch Hansen. Classic Operating Systems: from Batch Processing to Distributed Systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 0-387-95113-X. x + 597 pp. LCCN QA76.76.O63 B7425 2001. URL

http://www.loc.gov/catdir/
enhancements/fy0816/00045036-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0816/
00045036-t.html;   https:/
/link.springer.com/book/10.
1007/978-1-4757-3510-9.

BrinchHansen:2001:SOSa

[Bri01b] Per Brinch Hansen. The Solo operating system: A Concurrent Pascal program. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 324–336. ISBN 0-387-95113-X. LCCN QA76.76.O63 B7425 2001. URL

http://www.loc.gov/catdir/
enhancements/fy0816/00045036-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0816/
00045036-t.html;   https:/
/link.springer.com/book/10.
1007/978-1-4757-3510-9.

BrinchHansen:2002:OCP

[Bri02a] Per Brinch Hansen, editor. The Origin of Concurrent Programming: From Semaphores to Remote Procedure Calls. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. x + 534 pp. LCCN QA76.642

.O75 2002. URL http://link.
springer.com/10.1007/978-1-
4757-3472-0.

BrinchHansen:2002:PLC

[Bri02b] Per Brinch Hansen. The programming language Concurrent Pascal. In The Origin of Concurrent Programming: From Semaphores to Remote Procedure Calls [Bri02a], pages 297–

318. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN QA76.642

.O75 2002. URL http://link.

springer.com/10.1007/978-1-
4757-3472-0.

BrinchHansen:2002:SOSa

[Bri02c] Per Brinch Hansen. The Solo operating system: a Concurrent Pascal program. In The Origin of Concurrent Programming: From Semaphores to Remote Procedure Calls [Bri02a], pages 321– 333. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN QA76.642

.O75 2002. URL http://link.
springer.com/10.1007/978-1-
4757-3472-0.

BrinchHansen:2002:SPL

[Bri02d] Per Brinch Hansen. Super-Pascal: a publication language for parallel scientific computing. In The Origin of Concurrent Programming: From Semaphores to Remote Procedure Calls [Bri02a], pages 495– 524. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN QA76.642

.O75 2002. URL http://link.
springer.com/10.1007/978-1-
4757-3472-0.

Brumfield:1986:TOSa

[Bru86a] Jeffrey A. Brumfield. Teaching operating systems with Modula-2. In Proceedings of the Seventeenth SIGCSE Technical Symposium on Computer Science Education, SIGCSE ’86, pages 273– 282. ACM Press, New York, NY 10036, USA, 1986. ISBN 0-89791-178-4.

Brumfield:1986:TOSb

[Bru86b] Jeffrey A. Brumfield. Teaching operating systems with Modula-2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 18(1): 273–282, February 1986. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brumfield:1987:CPMa

[Bru87a]  Jeffrey Brumfield.  Concur-

rent programming in Modula-2. In Proceedings of the Eighteenth SIGCSE Technical Symposium on Computer Science Education, SIGCSE ’87, pages 191– 200. ACM Press, New York, NY 10036, USA, 1987. ISBN 0-89791-217-9.

Brumfield:1987:CPMb

[Bru87b] Jeffrey Brumfield. Concurrent programming in Modula-2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 19(1): 191–200, February 1987. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bauer:1976:LHI

[BS76] Friedrich Ludwig Bauer and Klaus Samelson, editors. Language hierarchies and interfaces: international summer school [held July 23 to August 2, 1975 in Marktoberdorf], volume 46 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. ISBN 0-387-07994-7 (New York), 3-540-07994-7 (Berlin), 3-540-37972-X (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .L335. URL http:/

/link.springer-ny.com/link/
service/series/0558/tocs/
t0046.htm;   http://www.springerlink.
com/content/978-0-387-07994-
3;   http://www.springerlink.
com/openurl.asp?genre=book&
isbn=978-3-540-07994-1;
http://www.springerlink.com/
openurl.asp?genre=issue&issn=
0302-9743&volume=46.

Bailes:1984:SBF

[BS84] Paul A. Bailes and Antonio Salvadori. A semantically-based formatting discipline for Pascal. Software—Practice and Experience, 14(3):235–251, March 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Boszormenyi:2003:MPL

[BS03] László Böszörményi and Peter Schojer, editors. Modular Programming Languages: Joint Modular Languages Conference, JMLC 2003, Klagenfurt, Austria, August 25–27, 2003: proceedings: 2003, volume 2789 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNCSD9. ISBN 3-540-40796-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .J6578 2003. URL http://

                                        link.springer-ny.com/link/
                                        service/series/0558/tocs/
                                        t2789.htm;   http://www.springerlink.
                                        com/content/978-3-540-40796-
                                        6;   http://www.springerlink.
                                        com/openurl.asp?genre=issue&
                                        issn=0302-9743&volume=2789;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=volume&
                                        id=doi:10.1007/b12023.

Bal:1989:PLD

[BST89] Henri E. Bal, Jennifer G. Steiner, and Andrew S. Tanenbaum. Programming languages for distributed computing systems. ACM Computing Surveys, 21(3): 261–322, September 1989. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (elec-

                                        tronic).  URL http://www.
                                        acm.org/pubs/toc/Abstracts/
                                        0360-0300/72552.html.

Budgen:1985:CMM

[Bud85] David Budgen. Combining MAS-COT with MODULA-2 to aid the engineering of real-time systems. Software—Practice and Experience, 15(8):767–793, August 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bell:1971:ALA

[BW71] D. A. Bell and B. A. Wichmann. An ALGOL-like assembly language for a small computer. Software—Practice and Experience, 1(1):61–72, January/ March 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brookes:1982:SAP

[BWA82] G. R. Brookes, I. R. Wilson, and A. M. Addyman. A static analysis of Pascal program structures. Software—Practice and Experience, 12(10):959–963, October 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Carroll:1985:MVP

[Car85] David W. Carroll. Modula-2 versus Pascal for microcomputers: an update. Dr. Dobb’s Journal of Software Tools, 10(11):28– 34, November 1985. CODEN DDJOEB. ISSN 1044-789X.

Cashman:1991:BET

[Cas91] Mark Cashman. The benefits of enumerated types in Modula-2. ACM SIGPLAN Notices, 26 (2):35–39, January 1991. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Corti:2001:AWC

[CBG01] Matteo Corti, Roberto Brega, and Thomas Gross. Approximation of worst-case execution time for preemptive multitasking systems. In Languages, Compilers, and Tools for Embedded Systems, pages 178–198. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-45245-1. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.
com/chapter/10.1007/3-540-
45245-1_12.

Chan:1987:ADD

[CC87] Fun Ting Chan and Tsong Yueh Chen. AIDA: a dynamic data flow anomaly detection system for Pascal programs. Software —Practice and Experience, 17 (3):227–239, March 1987. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Castaneda:1979:PTP

[CCN^{+}79] Fernando Castaneda, Frederick Chow, Peter Nye, Dan Sleator, and Gio Wiederhold. PCFORT — a Fortran-to-Pcode translator. Technical Report STAN-CS-79-714, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1979. vi + 85 pp. URL http:

                                        //i.stanford.edu/pub/cstr/
                                        reports/cs/tr/79/714/CS-TR-
                                        79-714.pdf.

Ciechanowicz:1984:CPT

[CD84] Z. J. Ciechanowicz and A. C. De Weever. The ‘completeness’ of the Pascal test suite. Software—Practice and Experience, 14(5):463–471, May 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cardelli:1992:MLD

[CDG^{+}92] Luca Cardelli, James Donahue, Lucille Glassman, Mick Jordan, Bill Kalsow, and Greg Nelson. Modula-3 language definition. ACM SIGPLAN Notices, 27(8):15–42, August 1992. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cardelli:1989:MTS

[CDJ^{+}89] L. Cardelli, J. Donahue, M. Jordan, B. Kalsow, and G. Nelson. The Modula 3 type system. In Proceedings of the 16th ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages, POPL ’89, pages 202–212. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-294-2.

Corney:1994:TTE

[CG94] Diane Corney and John Gough.

Type test elimination using typeflow analysis. In Programming Languages and System Architectures, pages 137–150. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-48356-X. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.
com/chapter/10.1007/3-540-
57840-4_29.

Coleman:1979:ACP

[CGHP79] Derek Coleman, Robin M. Gallimore, J. W. Hughes, and M. S. Powell. An assessment of Concurrent Pascal. Software— Practice and Experience, 9(10): 827–837, October 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Conway:1976:PP

[CGZ76] Richard Walter Conway, David Gries, and E. C. Zimmerman. A primer on PASCAL. Winthrop computer systems series. Winthrop Publishers, Cambridge, MA, USA, 1976. ISBN 0-87626-694-4. xii + 433 pp. LCCN QA76.73.P2 C66.

Coleman:1979:CTP

[CH79] D. Coleman and J. W. Hughes. The clean termination of Pascal programs. Acta Informatica, 11(3):195–210, March 22, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Chaik:2024:PCN

[Cha24] Andy Chaik. Pascal creator Niklaus Wirth has died. PC Gamer Web site, January 5,

                                        2024.  URL https://www.
                                        pcgamer.com/pascal-creator-
                                        niklaus-wirth-has-died/.

Christian:1986:GM

[Chr86]  Kaare Christian.  A Guide

to Modula-2. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 0-387-96242-5. xix + 436 pp. LCCN QA76.73.M63 C494 1986. US$28.00.

Czyzowicz:1985:IGT

[CI85] Jurek Czyzowicz and Michal Iglewski. Implementing generic types in Modula-2. ACM SIG-PLAN Notices, 20(12):26–32, December 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:1981:UPH

[CK81] Randy Clark and Stephen Koehler. The UCSD Pascal handbook. Prentice-Hall software series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1981. ISBN 0-13-935536-7 (paperback), 0-13-935544-8 (hardcover). xvi + 356 pp. LCCN QA76.73.P2 C56.

Cook:1982:CAP

[CL82] Robert P. Cook and Insup Lee.

A contextual analysis of Pascal programs. Software—Practice and Experience, 12(2):195– 203, February 1982. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Christiansen:1986:MLC

[CLL^{+}86] H. P. Christiansen, A. Lacroix, K. Lundberg, A. McKeeman, and P. D. V. van der Stok. More or less concerns about MODULA-2. ACM SIGPLAN Notices, 21(9):27–31, September 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cooper:1985:EMP

[CMH85] R. E. M. Cooper, Bruce J.

McKenzie, and R. Harries. Extensions to a microcoded Pascal compiler. Software—Practice and Experience, 15(5):519–522, May 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Collado:1987:MIC

[CMM87] M. Collado, R. Morales, and J. J.

Moreno. A Modula-2 implementation of CSP. ACM SIGPLAN Notices, 22(6):25–38, June 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chow:1980:UFU

[CNW80] Frederick Chow, Peter Nye, and Gio Wiederhold. UFORT: a fortran-to-universal PCODE translator. Technical Report CSL-TR-79-168, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1980. iv + 88 pp.

URL http://i.stanford.edu/
pub/cstr/reports/csl/tr/79/
168/CSL-TR-79-168.pdf.

Cohen:1991:TCT

[Coh91]  Norman H. Cohen.  Tech-

nical correspondence: Typeextension type tests can be performed in constant time. ACM Transactions on Programming Languages and Systems, 13(4): 626–629, October 1991. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org/pubs/toc/Abstracts/
0164-0925/115297.html. See

[Wir88e, Wir91c].

Collins:1984:CMP

[Col84]  Steve Collins.  Comparing

MODULA-2 with PASCAL and ADA. Data Processing, 26(10): 32–34, December 1984. CO-DEN ???? ISSN 0011-684X (print), 1878-3058 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0011684X84901795.

Comer:1979:MPM

[Com79] Douglas Comer. MAP: A Pascal macro preprocessor for large program development. Software —Practice and Experience, 9(3): 203–209, March 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Comer:1983:CBL

[Com83] Douglas Comer. The costs and benefits of lazy interactive input in Pascal. Software—Practice and Experience, 13(3):287–293, March 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cook:1989:EAL

[Coo89] R. P. Cook. An empirical analysis of the Lilith instruction set. IEEE Transactions on Computers, 38(1):156–158, January 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=8740.

Cornelius:1988:PLM

[Cor88] B. J. Cornelius. Problems with the language Modula-2. Software—Practice and Experience, 18(6):529–543, June 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Crawford:1985:HLI

[Cra85]  Albert L. Crawford.  High

level input/output in Modula-2. ACM SIGPLAN Notices, 20 (12):18–25, December 1985. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Regis:1990:OPO

[Cre90a]  Régis Crelier.  OP 2: A

portable Oberon compiler. Technical report 125, ETH Zürich, Zürich, Switzerland, February 1990. i + 46 pp. URL

http://hdl.handle.net/20.
500.11850/68707.

Crelier:1990:OPO

[Cre90b] Régis Crelier. OP2: a portable Oberon compiler. Report 125, ETH Zürich, Zürich, Switzerland, February 1990. 46 pp. URL

https://github.com/Oleg-N-
Cher/OfrontPlus/.

Crelier:1991:OPO

[Cre91] Régis Crelier. OP2: a portable Oberon compiler. In Anonymous [Ano91], pages 1–10. LCCN QA76.73.M63 I57 1991. URL

https://github.com/Oleg-N-
Cher/OfrontPlus/.

Cornelius:1980:MPP

[CRT80] B. J. Cornelius, D. J. Robson, and M. I. Thomas. Modification of the Pascal-P compiler for a single-accumulator one-address minicomputer. Software—Practice and Experience, 10(3):241–246, March 1980. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Comer:1982:AYM

[CW82] Douglas Comer and Keith Williamson. An alternative to Young’s module facility for Pascal. Software—Practice and Experience, 12(10):907–913, October 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cytron:1997:BNC

[Cyt97] Ron K. Cytron. Book nook:

Compiler Construction by Niklaus Wirth. ACM SIGPLAN Notices, 32(2):20–21, February 1997. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Daws:2024:CSP

[Daw24] Ryan Daws. Computer science pioneer Niklaus Wirth dies aged 89. Developer Tech Web site, January 4, 2024.

URL https://www.developer-
tech.com/news/2024/jan/04/
computer-science-pioneer-
niklaus-wirth-dies-aged-89/

.

Desjardins:1973:PCX

[Des73] P. Desjardins. A Pascal compiler for the Xerox SIGMA 6. ACM SIGPLAN Notices, 8(6):34–36, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Detmer:1993:PGP

[Det93] Richard C. Detmer. PC graphics packages for Modula-2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(3):52–57, September 1993. CODEN SIGSD3.

ISSN 0097-8418 (print), 2331-3927 (electronic).

Dotzel:1996:POS

[DG96] Günter Dotzel and Hartmut Goebel. Porting the Oberon System to AlphaAXP. ACM SIGPLAN Notices, 31(11):66– 73, November 1996. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dotzel:1998:BO

[DG98] Günter Dotzel and Hartmut Goebel. 64 bit Oberon. ACM SIGPLAN Notices, 33(2):56–58, February 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dijkstra:2000:TCW

[Dij00] Edsger W. Dijkstra. On the transitive closure of a well-founded relation. In Böszörményi [Bös00], pages 31–40. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1 (dpunktVerl.). LCCN QA76.756 .S36 2000. URL http:/

                                        /digitale-objekte.hbz-nrw.
                                        de/storage2/2016/09/08/file_
                                        8/6866144.pdf.

Disteli:1993:OPM

[Dis93] Andreas R. Disteli. OBERON for PC on an MS-DOS base. Technical report 203, ETH Zürich, Zürich, Switzerland, November 1993. i + 55 pp.

                                        URL http://hdl.handle.net/
                                        20.500.11850/68911.

DiSanto:1990:PCA

[DNR90] Michele Di Santo, Libero Nigro, and Wilma Russo. Programmerdefined control abstractions in Modula-2. Computer Languages, 15(3):141–152, ???? 1990. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Djavaheri:1985:MAC

[DO85] Morris Djavaheri and Stan Osborne. Modula-2 — an alternative to C for system programming. In USENIX Association [USE85], pages 34–42.

Dotzel:1990:DMG

[Dot90] Günter Dotzel. Does Modula-2 generate racehorses? Comparison of compiler generated code quality for floating point arithmetic. ACM SIGPLAN Notices, 25(12):85–88, December 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dubery:1985:SAP

[DP85] J. M. Dubery and A. J. Pinches.

Software for an air pollution measuring system: an application of Modula. Software —Practice and Experience, 15 (4):413–422, April 1985. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Disteli:1997:COA

[DR97] A. R. Disteli and P. Reali. Combining Oberon with active objects. Lecture Notes in Computer Science, 1204:221–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeMarco:1984:SDC

[DS84] Tom DeMarco and Aurel Soceneantu. SYNCRO: A dataflow command shell for the Lilith/ Modula computer. In Proceedings of the 7th International Conference on Software Engineering, ICSE ’84, pages 207– 213. IEEE Press, 1984. ISBN 0-8186-0528-6.

Dowsing:1986:WCA

[DS86] R. D. Dowsing and M. T. Sanderson. Writing concurrent assemblers — a case study in Path Pascal. Software—Practice and Experience, 16(12):1117–1135, December 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dotzel:1997:OHP

[DSD97] Günter Dotzel, Wojtek Skulski, and Paul F. Dubois. Oberon-2, a high-performance alternative to C++. Computers in Physics, 11 (1):81–??, January 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.
org/doi/10.1063/1.4822520.

Dunman:1982:MIC

[DSW82] B. R. Dunman, Stephen R. Schach, and Peter T. Wood. A mainframe implementation of Concurrent Pascal. Software —Practice and Experience, 12 (1):85–89, January 1982. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dubery:1985:MRR

[Dub85] J. M. Dubery. Modulas-1, -2, and -3(?) or a real real-time language. Annual Review in Automatic Programming, 13 (part 1)(??):81–89, ???? 1985. CO-DEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0066413885904483.

Dunford:1984:SFPb

[Dun84] Christopher J. Dunford. Savage floating-point benchmark in Modula-2 in 16BST. Dr. Dobb’s Journal of Software Tools, 9(8): 106–??, August 1984. CODEN DDJOEB. ISSN 1044-789X.

Eberle:2000:DCN

[Ebe00] Hans Eberle. Designing a cluster network. In Böszörményi [Bös00], pages 153–172. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1 (dpunktVerl.). LCCN QA76.756 .S36 2000. URL http:/

/digitale-objekte.hbz-nrw.
de/storage2/2016/09/08/file_
8/6866144.pdf.

Egorov:1993:YAA

[EBP93] Igor I. Egorov, Ruslan P. Bogatyrev, and Dmitry L. Petrovichev. Yet another approach to Modula-2 implementation of exception handling mechanism. Structured Programming, 14(1): 23–36, 1993. CODEN STPGEM. ISSN 0935-1183.

Edwards:1977:BRB

[Edw77] R. Edwards. Book review: Algorithms + Data Structures = Programs, N. Wirth, Prentice-Hall Inc., Engelwood Cliffs, N.J., 1976. no. of pages: 366. price: £9.70. Software—Practice and Experience, 7(2):303–304, March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Egenhofer:1988:PMT

[EF88] Max J. Egenhofer and Andrew U. Frank. A precompiler for modular, transportable Pascal. ACM SIGPLAN Notices, 23(3):22–32, March 1988. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Engeln-Mullges:1991:FNM

[EMR91] Gisela Engeln-Müllges and Fritz Reutter. Formelsammlung zur numerischen Mathematik mit Turbo Pascal-Programmen. (German) [Collection of formulas for numerical mathematics with Turbo Pascal programs]. Bibliographisches Institut, Mannheim, Germany, third edition, 1991. ISBN 3-411-15003-3. xxiv + 912 pp. LCCN QA297 .E55 1991. With an appendix by Albert Becker and Jürgen Dietel.

Emry:2021:PPL

[Emr21] Katharina Emry. Pascal: a programming language that conquered the world. Web site, April 20, 2021. URL

https://inf.ethz.ch/news-
and-events/spotlights/infk-
news-channel/2021/04/niklaus-
wirth-pascal-conquers-the-
world.html.

Ershov:1974:IST

[EN74] Andrei Ershov and Valery Aleksandrovich Nepomniaschy, editors. International Symposium on Theoretical Programming [held in Novosibirsk, Aug. 7–11, 1972], volume 5 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. ISBN 0-387-06720-5 (New York paperback), 3-540-06720-5 (Berlin paperback). LCCN QA76 .I5798 1972.

Ernst:1980:SAD

[EO80] George W. Ernst and William F.

Ogden. Specification of abstract data types in Modula. ACM Transactions on Programming Languages and Systems, 2 (4):522–543, October 1980. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Evers:1992:MNO

[ER92] David Evers and Peter Robinson. Modula-3 network objects over ANSA: Heterogeneous object-based RPC in a modern systems programming language. In Proceedings of the 5th Workshop on ACM SIGOPS European Workshop: Models and Paradigms for Distributed Systems Structuring, EW 5, pages 1–5. ACM Press, New York, NY 10036, USA, 1992. ISBN 1-4503-7340-2.

ECS:1985:SCP

[Eur85] European Committee for Standardization. Specification for computer programming language PASCAL: ISO 7185-1983. British Standards Institute, London, UK, 1985. ISBN 0-580-12531-9. 73 pp. LCCN

                                        ???? URL http://www.iso.ch/
                                        cate/d13802.html. This Euro-

pean standard was accepted by CEN on 1985-08-09 … and it … exists in three official versions (English, French, German) (BS 6192: 1982, ISO 7185-1983, EN 27 185: 1985).

Evans:1984:CPL

[Eva84] Arthur Evans, Jr. A comparison of programming languages Ada, Pascal, C. In Feuer and Gehani [FG84a], page ?? ISBN 0-13-154840-9 (paperback), 0-13-154857-3 (hardcover). LCCN QA76.73.A35 C66 1984.

Feiereisen:1974:IPP

[Fei74] L. Feiereisen. Implementation of PASCAL on the PDP 11/45. In ????, editor, DECUS Conference, Zürich, September 1974, page 259. ????, ????, 1974.

Feldman:1986:AVM

[Fel86a] M. B. Feldman. Ada vs. Modula-2: a response from the Ivory Tower. ACM SIGPLAN Notices, 21(5):39–44, May 1986. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feldman:1986:MPOa

[Fel86b] Michael B. Feldman. Modula-2 projects for an operationsystems course: Racing sorts and multiple windows. In Proceedings of the Seventeenth SIGCSE Technical Symposium on Computer Science Education, SIGCSE ’86, pages 283– 288. ACM Press, New York, NY 10036, USA, 1986. ISBN 0-89791-178-4.

Feldman:1986:MPOb

[Fel86c] Michael B. Feldman. Modula-2 projects for an operationsystems course: Racing sorts and multiple windows. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 18(1):283–288, February 1986. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Feldman:1988:DSM

[Fel88] Michael B. Feldman. Data Structures with Modula-2. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1988. ISBN 0-13-197344-4. xvii + 301 pp. LCCN QA76.73.M63 F45 1988. US$40.00. Includes chapter on hash tables.

Fernandez:1995:SELa

[Fer95a] Mary F. Fernández. Simple and effective link-time optimization of Modula-3 programs. In Proceedings of the ACM SIGPLAN 1995 Conference on Programming Language Design and Implementation, PLDI ’95, pages 103–115. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-697-2.

Fernandez:1995:SELb

[Fer95b]  Mary F. Fernández.  Sim-

ple and effective link-time optimization of Modula-3 programs. ACM SIGPLAN Notices, 30(6): 103–115, June 1995. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fessi:2024:PPL

[Fes24]  Ahmed Fessi.  Pascal pro-

gramming language — RIP Niklaus Wirth. Web site, January 7, 2024. URL https:/

                                        /medium.com/@AhmedF/pascal-
                                        programming-language-rip-
                                        niklaus-wirth-d230a3d9ea88.

Frohlich:2000:SAM

[FF00] Peter H. Fröhlich and Michael Franz. Stand-alone messages. In Modular Programming Languages, pages 90–103. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-44519-6. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.com/
                                        chapter/10.1007/10722581_9.

Feuer:1982:CPL

[FG82a] Alan R. Feuer and Narain H. Gehani. Comparison of the programming languages C and Pascal. ACM Computing Surveys, 14 (1):73–92, March 1982. CODEN CMSVAN. ISSN 0010-4892. See also [Pet82, FG82b].

Feuer:1982:SFI

[FG82b] Alan R. Feuer and Narain H.

Gehani. Surveyor’s forum: Idiomatic programming. ACM Computing Surveys, 14(4):625– 626, December 1982. CODEN CMSVAN. ISSN 0010-4892. See [FG82a, Pet82].

Feuer:1984:CAP

[FG84a] Alan R. Feuer and Narain Gehani, editors. Comparing and Assessing Programming Languages: Ada, C, and Pascal. Prentice-Hall software series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-154840-9 (paperback), 0-13-154857-3 (hardcover). xiv + 271 pp. LCCN QA76.73.A35 C66 1984.

Feuer:1984:CPL

[FG84b] Alan R. Feuer and Narain H.

Gehani. A comparison of the programming languages C and Pascal. In Feuer and Gehani [FG84a], page ?? ISBN 0-13-154840-9 (paperback), 0-13-154857-3 (hardcover). LCCN QA76.73.A35 C66 1984.

Fischer:1998:OCG

[FH98] A. L. Fischer and M. Hannes.

The Oberon Companion: a Guide to Using and Programming Oberon System 3. Hochsvhulverlag AG, Zürich, Switzerland, 1998. ISBN 3-7281-2493-1. xi + 322 pp. LCCN QA76.76.O63 F566 1998.

Fors:2015:JIO

[FH15] Niklas Fors and Görel Hedin. A JastAdd implementation of Oberon-0. Science of Computer Programming, 114(??):74– 84, December 15, 2015. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S016764231500026X.

Findlay:1977:BRB

[Fin77] W. Findlay. Book review: A primer on Pascal, R. Conway, D. Gries and E. C. Zimmerman, Winthrop Publishers Inc., Cambridge, Mass., 1976. No. of pages: 433. E7-95 (paperback). Software—Practice and Experience, 7(3):432, June 1977. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Franz:1991:PR

[FK91] Michael Franz and Thomas Kistler. Portability redefined. In Anonymous [Ano91], pages 216– 224. LCCN QA76.73.M63 I57 1991.

Franz:1997:SB

[FK97] Michael Franz and Thomas Kistler. Slim binaries. Communications of the ACM, 40(12): 87–94, December 1997. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1997-40-12/
p87-franz/.

Floyd:20xx:DOL

[Floxx] Paul Floyd. A description of the Oberon-2 language. Web site,

20xx. URL http://www.edm2.
com/0608/oberon2.html.

Forsythe:1986:CDS

[FM86] Kenneth Forsythe and Stanislaw Matwin. Copying of dynamic structures in a Pascal environment. Software—Practice and Experience, 16(4):335–340, April 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Forrester:1988:WMC

[For88] J. M. Forrester. Wirth meets the countess (Modula-2 to Ada conversion). Ada User, 9(2):83–86, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

Foster:1986:SCM

[Fos86] David G. Foster. Separate compilation in a Modula-2 compiler. Software—Practice and Experience, 16(2):101–106, February 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Foulkes:1975:PCI

[Fou75]  William Bruce Foulkes.  A

Pascal Compiler For IBM 360/370 Computers. Ph.D. thesis, Department of Computer Science, Universal of Manitoba, Winnipeg, MB, Canada, 1975. 267 pp. URL

https://www.proquest.com/
pqdtglobal/docview/302819505.

Fountain:1991:MCP

[Fou91] ?. Fountain. Modula’s children, part II: Oberon. Byte Magazine, 16(??):??, ?? 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Franz:1990:IM

[Fra90a] Michael Franz. The implementation of MacOberon. Technical report 141, ETH Zürich, Zürich, Switzerland, October 1990. i + 28 pp. URL

                                        http://hdl.handle.net/20.
                                        500.11850/68699.

Franz:1990:MRM

[Fra90b]  Michael Franz.  MacOberon

reference manual. Technical report 142, ETH Zürich, Zürich, Switzerland, November 1990. i + 32 pp.

                                        URL http://hdl.handle.net/
                                        20.500.11850/68900.

Franz:1993:EOS

[Fra93] Michael Franz. Emulating an operating system on top of another. Software—Practice and Experience, 23(6):677–692, June 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Franz:1994:COS

[Fra94a] Michael Franz. Compiler optimizations should pay for themselves. In Peter Schulthess, editor, Advances in modular languages: Proceedings of the Joint Modular Languages Conference, University of Ulm, Germany, 28–30 September 1994, pages 111–121. Universitätsverlag, Ulm, Germany, 1994. ISBN 3-89559-220-X. LCCN QA76.6 .J6578 1994.

Franz:1994:CGF

[Fra94b] Michael Steffen Oliver Franz.

Code-Generation On-the-Fly: a Key to Portable Software. Dr.sc.tech. dissertation (number 10497), ETH Zürich, Zürich, Switzerland, 1994. 97 pp. URL

https://www.proquest.com/
pqdtglobal/docview/304133041.

Franz:1997:PLL

[Fra97] Michael Franz. The programming language Lagoona: a fresh look at object-orientation. Software — Concepts and Tools, 18 (1):14–26, ???? 1997. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Franz:2000:OOJa

[Fra00a]  Michael Franz.  Oberon —

the overlooked jewel. Report, University of California, Irvine, Irvine, CA, USA, June

15, 2000.  13 pp.  URL

https://people.cis.ksu.edu/
~danielwang/Investigation/
System_Security/download.
pdf.

Franz:2000:OOJb

[Fra00b] Michael Franz. Oberon: the overlooked jewel. In Böszörményi [Bös00], pages 41–54. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1 (dpunktVerl.). LCCN QA76.756 .S36 2000. URL http:/

/digitale-objekte.hbz-nrw.
de/storage2/2016/09/08/file_
8/6866144.pdf.

Freak:1981:FPT

[Fre81] R. A. Freak. A Fortran to Pascal translator. Software—Practice and Experience, 11(7):717–732, July 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Freeman:1995:PRM

[Fre95] Steve Freeman. Partial revelation and Modula-3. Dr. Dobb’s Journal of Software Tools, 20 (10):36, 38, 40, 42, 110, 112, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Feijen:1990:BOB

[FvGGM90] W. H. J. Feijen, A. J. M. van Gasteren, David Gries, and J. Misra, editors. Beauty is our Business: a Birthday Salute to Edsger W. Dijkstra. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. xix + 453 pp. LCCN QA76 .B326 1990. URL http://

www.zentralblatt-math.org/
zmath/en/search/?an=0718.

68004. Contains important treatment of accurate binaryto-decimal conversion [Gri90, Knu90].

Forsythe:1965:AGPa

[FW65a] George E. Forsythe and Niklaus Wirth. Automatic grading programs. Technical Report CS-TR-65-17, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1965. ii + 17 pp. URL http:

//i.stanford.edu/pub/cstr/
reports/cs/tr/65/17/CS-TR-
65-17.pdf.

Forsythe:1965:AGPb

[FW65b] George E. Forsythe and Niklaus Wirth. Automatic grading programs. Communications of the ACM, 8(5):275–278, May 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gabrini:1986:CPMa

[GAK86a] Philippe J. Gabrini, J. Mack Adams, and Barry L. Kurtz. Converting from Pascal to Modula-2 in the undergraduate curriculum. In Proceedings of the Seventeenth SIGCSE Technical Symposium on Computer Science Education, SIGCSE ’86, pages 50–52. ACM Press, New York, NY 10036, USA, 1986. ISBN 0-89791-178-4.

Gabrini:1986:CPMb

[GAK86b] Philippe J. Gabrini, J. Mack Adams, and Barry L. Kurtz. Converting from Pascal to Modula-2 in the undergraduate curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 18(1):50–52, February 1986. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Graham:2012:PLP

[GALW12] Susan Graham, Frances E. Allen, Barbara Liskov, and Niklaus Wirth. Programming languages past achievements and future challenges. In ACM Turing Centenary Celebration, ACM-TURING ’12. ACM Press, New York, NY 10036, USA, 2012. ISBN 1-4503-3919-0.

Guoliang:1988:KMI

[GC88] Z. Guoliang and Z. Chengxiang. The kernel of Modula-2 integrated environment. ACM SIGSOFT Software Engineering Notes, 13(4):57–60, October 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gough:2000:EJV

[GC00] K. John Gough and Diane Corney. Evaluating the Java Virtual Machine as a target for languages other than Java. In Modular Programming Languages, pages 278–290. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-44519-6. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/10722581_22.

Gough:2003:LMF

[GC03] K. John Gough and Diane Corney. Leveraging managed frameworks from modular languages.

In Modular Programming Languages, pages 150–162. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-45213-3. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.
com/chapter/10.1007/978-3-
540-45213-3_20.

Gehring:2000:LVS

[Geh00] Stephan W. Gehring. Learning the value of simplicity. In Böszörményi [Bös00], pages 95–100. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1 (dpunktVerl.). LCCN QA76.756 .S36 2000. URL http:

//digitale-objekte.hbz-nrw.
de/storage2/2016/09/08/file_
8/6866144.pdf.

Georgiadis:1975:DCP

[Geo74] Panagiotis I. Georgiadis. Design and construction of a PASCAL compiler. M.Sc. thesis, University of Warwick, Coventry CV4 7AL, UK, 1974. 65 pp.

Gerritse:1982:NEP

[Ger82] J. J. Gerritse. A note on extension of Pascal by coroutines. Software—Practice and Experience, 12(12):1163–1164, December 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Glaser:1993:DAM

[GH93] H. Glaser and R. Harrison.

Data abstraction in Modula-2. Information and Software Technology, 35(11–12):619–626, November/December 1993. CO-DEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/095058499390076F.

Gianni:1989:SAC

[Gia89] P. (Patrizia) Gianni, editor. Symbolic and algebraic computation: International Symposium ISSAC ’88, Rome, Italy, July 4– 8, 1988: proceedings, volume 358 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Graef:1979:HDI

[GKLM79] Norwin Graef, Horst Kretschmar, Klaus-Peter Löhr, and Bernt Morawetz. How to design and implement small time-sharing systems using Concurrent Pascal. Software—Practice and Experience, 9(1):17–24, January 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gunther:1974:ICS

[GLL74] A. Günther, B. Levrat, and H. Lipps, editors. International Computing Symposium 1973: Proceedings of the International Computing Symposium 1973, Davos, Switzerland, 4–7 September 1973. North-Holland, Amsterdam, The Netherlands, 1974. ISBN 0-444-10596-4, 0-7204-2097-0. LCCN QA76 .I567 1973.

Grosse-Lindemann:1976:PPC

[GLN76] C. O. Grosse-Lindemann and H. H. Nagel. Postlude to a PASCAL-compiler bootstrap on a DECSystem-10. Software —Practice and Experience, 6 (1):29–42, January/March 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gehring:1994:LDD

[GLW94] S. Gehring, S. Ludwig, and N. Wirth. A laboratory for a digital design course using FP-GAs. Lecture Notes in Computer Science, 849:385–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Griesemer:2000:CJH

[GM00] Robert Griesemer and Srdjan Mitrovic. A compiler for the Java HotSpot^{T M} virtual machine. In Böszörményi [Bös00], pages 133–152. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1 (dpunktVerl.). LCCN QA76.756 .S36 2000. URL http:

//digitale-objekte.hbz-nrw.
de/storage2/2016/09/08/file_
8/6866144.pdf.

Greenfield:1987:DUM

[GN87] S. Greenfield and R. Norton.

Detecting uninitialized Modula-2 abstract objects. ACM SIGPLAN Notices, 22(6):52–58, June 1987. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goldberg:1985:RSC

[Gol85] Morton Goldberg. A response to “Some Concerns about Modula-2”. ACM SIGPLAN Notices, 20 (8):71–72, August 1985. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [Mof84].

Goldsby:1986:CUG

[Gol86] M. E. Goldsby. Concurrent use of generic types in MODULA-2. ACM SIGPLAN Notices, 21(6):28–34, June 1986. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goldsby:1987:SNQ

[Gol87] Michael E. Goldsby. Solving the “N ≤ 8 Queens” problem with CSP and Modula-2. ACM SIGPLAN Notices, 22(2): 43–52, February 1987. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Greenwood:1986:CVT

[Gre86] J. R. Greenwood. Comments on “A View from the Trenches”. Ada vs. Modula-2 vs. Praxis. ACM SIGPLAN Notices, 21 (5):45–49, May 1986. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [Bie85].

Gries:1990:BDO

[Gri90] David Gries. Binary to decimal, one more time. In Feijen et al. [FvGGM90], chapter 16, pages 141–148. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326 1990. URL http:/

/www.zentralblatt-math.org/
zmath/en/search/?an=0718.

68004. This paper presents an alternate proof of Knuth’s algorithm [Knu90] for conversion between decimal and fixed-point binary numbers.

Gruber:2024:NWT

[Gru24] John Gruber. Niklaus Wirth, titan of computer science and creator of Pascal, dies at 89. Web site, January 2024. URL

https://daringfireball.net/
linked/2024/01/11/niklaus-
wirth-rip.

Garfinkel:2024:NMN

[GS24] Simson Garfinkel and Eugene H.

Spafford. News: In memoriam: Niklaus Wirth. Communications of the ACM, 67 (3):20, March 2024. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3641309.

Gutknecht:1984:TM

[Gut84] Jürg Gutknecht. Tutorial on Modula-2. Byte Magazine, 9(8): 157–176, August 1984.

Gutknecht:1985:CTE

[Gut85] J. Gutknecht. Concepts of the text editor Lara. Communications of the ACM, 28(9):942– 960, September 1985. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic).  URL http://www.
                                        acm.org/pubs/toc/Abstracts/
                                        0001-0782/4288.html.

Gutknecht:1986:SCM

[Gut86] J. Gutknecht. Separate compilation in Modula-2: an approach to efficient symbol files. IEEE Software, 3(6):29–38, November 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Gutknecht:1987:OPC

[Gut87] Jürg Gutknecht. One-pass compilation at its limits — a Modula-2 compiler for the Xerox Dragon computer. Software—Practice and Experience, 17(7):469–484, July 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gutknecht:1990:OGS

[Gut90] Jürg Gutknecht. The Oberon guide: System release 1.2. Technical report 138, ETH Zürich, Zürich, Switzerland, October 1990. i + 60 pp.

                                        URL http://hdl.handle.net/
                                        20.500.11850/68700.

Gutknecht:1994:OSV

[Gut94] Jürg Gutknecht. Oberon System 3: Vision of a future software technology. Software — Concepts and Tools, 15(1):26–33, ???? 1994. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Gutknecht:1997:DFR

[Gut97] Jürg Gutknecht. Do the fish really need remote control? A proposal for self-active objects in Oberon. In Modular Programming Languages, volume 1204 of Lecture Notes in Computer Science, pages 207–220. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNCSD9. ISBN 3-540-68328-3. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/3-540-62599-2_41.

Gutknecht:1999:OIL

[Gut99] Jürg Gutknecht. Oberon as an implementation language for COM objects. Technical report 323, Institute for Scientific Computing, ETH Zürich, Zürich, Switzerland, 1999. 22

pp. URL http://hdl.handle.
net/20.500.11850/68819.

Gutknecht:2000:CCV

[Gut00a] Jürg Gutknecht. Compiler construction versus Lotus Notes^{T M}: a strange battle. In Böszörményi [Bös00], pages 101–120. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1 (dpunktVerl.). LCCN QA76.756 .S36 2000. URL http:/

/digitale-objekte.hbz-nrw.
de/storage2/2016/09/08/file_
8/6866144.pdf.

Gutknecht:2000:OIL

[Gut00b] Jürg Gutknecht. Oberon as an implementation language for COM components: a case study in language interoperability. In Gutknecht and Weck [GW00], pages 119–133. CODEN LNCSD9. ISBN 3-540-67958-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .J6578 2000. URL http://

                                        link.springer-ny.com/link/
                                        service/series/0558/tocs/
                                        t1897.htm;   http://www.springerlink.
                                        com/content/978-3-540-67958-
                                        5;   http://www.springerlink.
                                        com/openurl.asp?genre=issue&
                                        issn=0302-9743&volume=1897.

Gilbert:1977:PCI

[GW77] Erik J. Gilbert and David W. Wall. P-Code intermediate assembly language (PAIL-3). S-l project document PAIL-3, Stanford University, Department of Computer Science, Stanford, CA, USA, July 18, 1977. ???? pp. URL ????.

Gilbert:1978:PCI

[GW78] Erik J. Gilbert and David W. Wall. P-Code Intermediate Assembler Language (PAIL-4). Technical Note 148, Stanford University, Department of Computer Science, Stanford, CA, USA, March 18, 1978. ii +

                                        28 pp.  URL http://www.
                                        bitsavers.org/pdf/stanford/
                                        sel_techReports/TN148_P-Code_
                                        AsmLang_PAIL-4_Mar78.pdf.

Gutknecht:1984:ADP

[GW84] Jürg Gutknecht and Werner Winiger. Andra: The document preparation system of the personal workstation Lilith. Software—Practice and Experience, 14(1):73–100, January 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gutknecht:2000:MPL

[GW00] Jurg Gutknecht and Wolfgang Weck, editors. Modular programming languages: Joint Modular Languages Conference, JMLC 2000, Zürich, Switzerland, September 6–8, 2000: Proceedings, volume 1897 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNCSD9. ISBN 3-540-67958-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .J6578 2000. URL http://

link.springer-ny.com/link/
service/series/0558/tocs/
t1897.htm;   http://www.springerlink.
com/content/978-3-540-67958-
5;   http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=1897.

Haa:1982:CSS

[Haá82] Anna Haá. Computer system simulation in PASCAL. Software —Practice and Experience, 12 (8):777–784, August 1982. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Habermann:1973:CCP

[Hab73] A. Nico Habermann. Critical comments on the programming language Pascal. Acta Informatica, 3(1):47–57, December 31, 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [Wir71d, LD75].

Habermann:1984:CCP

[Hab84] A. N. Habermann. Critical comments on the programming language Pascal. In Feuer and Gehani [FG84a], page ?? ISBN 0-13-154840-9 (paperback), 0-13-154857-3 (hardcover). LCCN QA76.73.A35 C66 1984.

Harbison:1990:MIO

[Har90] S. Harbison. Modula-3: An introduction to the OOP language that grew from Pascal and Modula-2. Byte Magazine, 15 (12):385–388, 390, 392, November 1990. CODEN BYTEDJ. ISSN 0360-5280.

Harbison:1992:SPM

[Har92] Sam Harbison. Safe programming with Modula-3. Dr. Dobb’s Journal of Software Tools, 17 (10):88–96, 126–127, October 1992. CODEN DDJOEB. ISSN 1044-789X.

Harbison:1994:MPL

[Har94] Sam Harbison. The Modula-3 programming language. Dr. Dobb’s Journal of Software Tools, 19(??):24, ???? 1994. CO-DEN DDJOEB. ISSN 1044-789X. Repeat of Oct92 article.

Hayden:1980:LER

[Hay80] Charles Hayden. Letters to the Editor: Remarks on ‘Implementation of concurrent Pascal on LSI-11’. Software—Practice and Experience, 10(10):849–850, October 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See [Mat80].

Hennessy:1982:DIP

[HE82] John L. Hennessy and Hilding Elmquist. The design and implementation of parametric types in PASCAL. Software—Practice and Experience, 12(2):169– 184, February 1982. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Heinlein:2003:SEP

[Hei03] Christian Heinlein. Safely extending procedure types to allow nested procedures as values. In Modular Programming Languages, pages 144–149. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-45213-3. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.
com/chapter/10.1007/978-3-
540-45213-3_19.

Hemmendinger:1988:UPS

[Hem88] David Hemmendinger. Unfair process scheduling in Modula-2. ACM SIGPLAN Notices, 23(3):7–12, March 1988. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hitz:1995:MVCa

[HH95a] Martin Hitz and Marcus Hudec. Modula-2 versus C++ as a first programming language — some empirical results. In Proceedings of the Twenty-Sixth SIGCSE Technical Symposium on Computer Science Education, SIGCSE ’95, pages 317– 321. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-693-X.

Hitz:1995:MVCb

[HH95b] Martin Hitz and Marcus Hudec. Modula-2 versus C++ as a first programming language — some empirical results. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):317–321, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hirschy:1983:HOS

[Hir83] Eric Hirschy. Hermes: an operating system for a Modula-2 environment. In Proceedings of the 1983 ACM SIGS-MALL Symposium on Personal and Small Computers, SIGS-MALL ’83, pages 163–167. ACM Press, New York, NY 10036, USA, 1983. ISBN 0-89791-123-7.

Hitson:1979:PPC

[Hit79] Bruce L. Hitson. Pascal/P-Code cross compiler for the LSI-11. Technical Report SLAC-PUB-2246, Stanford Linear Accelerator Center, Stanford, CA, USA, January 1979.

4 pp. URL https://www.slac.
stanford.edu/pubs/slacpubs/
2000/slac-pub-2246.pdf.

Horning:1993:SUM

[HKMN93] Jim Horning, Bill Kalsow, Paul McJones, and Greg Nelson. Some useful Modula-3 interfaces. Memo 113, Digital Equipment Corporation, Systems Research Center, Palo Alto, CA, USA, December 1993.

Hansen:1976:PSR

[HL76] Gilbert J. Hansen and Charles E.

Lindahl. Preliminary specification of Real-Time Pascal. Technical Report 76-C-0017-l, Department of Computer and Information Science, University of Florida, Gainesville, FL 32611, USA, July 1976. iv + 86 pp.

URL https://apps.dtic.mil/
sti/tr/pdf/ADA031451.pdf.

Huskey:1963:SDB

[HLW63] H. D. Huskey, Ralph Love, and Niklaus Wirth. A syntactic description of BC NELIAC. Communications of the ACM, 6(7): 367–375, July 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hennessy:1982:CPC

[HM82] John L. Hennessy and Noah Mendelsohn. Compilation of the Pascal case statement. Software—Practice and Experience, 12(9):879–882, September 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Heeb:1991:HDW

[HN91] Beat Heeb and Immo Noack. Hardware description of the workstation Ceres-3. Technical report 168, ETH Zürich, Zürich, Switzerland, October 1991. 58

                                        pp. URL http://hdl.handle.
                                        net/20.500.11850/69237.

Hoare:1974:ADP

[Hoa74] C. A. R. Hoare. An axiomatic definition of the programming language PASCAL. In Ershov and Nepomniaschy [EN74], pages 1–16. ISBN 0-387-06720-5 (New York paperback), 3-540-06720-5 (Berlin paperback). LCCN QA76 .I5798 1972.

Hof:2000:CMS

[Hof00] Markus Hof. Composable message semantics in Oberon. In Gutknecht and Weck [GW00], pages 11–25. CODEN LNCSD9. ISBN 3-540-67958-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .J6578 2000. URL http://

                                        link.springer-ny.com/link/
                                        service/series/0558/tocs/
                                        t1897.htm;   http://www.springerlink.
                                        com/content/978-3-540-67958-
                                        5;   http://www.springerlink.
                                        com/openurl.asp?genre=issue&
                                        issn=0302-9743&volume=1897.

Holzheu:2021:INW

[Hol21]  Michael Holzheu.  Interview

mit Niklaus Wirth. Web video (1h), 2021. URL https:

                                        //www.youtube.com/watch?v=
                                        OEmMx55SF8U.

Hoppe:1980:SNW

[Hop80]  Jiri Hoppe.  A simple nu-

cleus written in Modula-2: a case study. Software—Practice and Experience, 10(9):697– 706, September 1980. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hoppe:2000:DLW

[Hop00] Jirka Hoppe. Daily life with N. Wirth. In Böszörményi [Bös00], pages 247–252. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1 (dpunktVerl.). LCCN QA76.756 .S36 2000. URL http:/

/digitale-objekte.hbz-nrw.
de/storage2/2016/09/08/file_
8/6866144.pdf.

Hartel:1984:MIO

[HS84] P. H. Hartel and D. Starreveld. Modula-2 implementation overview. Journal of Pascal, Ada and Modula-2, ??(??):9–23, July/ August 1984. CODEN JOPAD5. ISSN 0735-1232.

Horton:1986:UCP

[HT86] I. A. Horton and S. J. Turner.

Using coroutines in Pascal. Software—Practice and Experience, 16(1):45–61, January 1986. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hurst:1980:PPP

[Hur80] A. John Hurst. Pascal-P, program structure and program behaviour. Software—Practice and Experience, 10(12):1029– 1036, December 1980. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hurst:1984:SLP

[Hur84] A. J. Hurst. A source language performance monitoring facility for the B1800 Modula interpreter. ACM SIG Micro Newsletter, 15(3):28–36, September 1984. CODEN SIGMDJ. ISSN 1050-916X.

Hoare:1972:ADP

[HW72] C. A. R. Hoare and Niklaus Wirth. An axiomatic definition of the programming language PASCAL. Report, ETH Zürich, Zürich, Switzerland, November 1972. v +

29 pp. URL https://www.
research-collection.ethz.
ch/bitstream/handle/20.500.
11850/68663/eth-3028-01.pdf.

Hoare:1973:ADP

[HW73] C. A. R. Hoare and Niklaus Wirth. An axiomatic definition of the programming language Pascal. Acta Informatica, 2(4):335–355, December 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See addenda and corrigenda [HW74].

Hoare:1974:ACA

[HW74] C. A. R. Hoare and N. Wirth. Addenda and corrigenda to An Axiomatic Definition of the Programming Language Pascal. Acta Informatica, 3(3):296, July 22, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [HW73].

IEEE-PSC:1983:ISP

[IAJ83] IEEE Pascal Standards Committee, American National Standards Institute Committee X3J9, and Joint ANSI/X3J9-IEEE Pascal Standards Committee. IEEE standard Pascal computer programming language: an American National Standard. ANSI/IEEE; 770 X3.97 ANSI/IEEE Std. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1983. various pp.

ISB:1983:ISP

[IEE83] IEEE Standards Board, editor. IEEE Standard Pascal computer programming language: an American National Standard. IEEE., New York, NY, USA, 1983. ISBN 0-471-88944-X. 128 pp. LCCN QA76.73.P2 A43 1983.

Iimori:1971:LPF

[IH71] Sueo Iimori and Seiiti Huzino.

List processing for finding Wirth–Weber type precedence relations between symbols. Memoirs of the Faculty of Science. Kyushu University. Series A. Mathematics, 25:152–166, 1971. ISSN 0373-6385 (print), 1883-2172 (electronic).

ISO:1982:SCP

[ISO82] ISO. Specification for Computer Programming Language Pascal, ISO 7185-1982. International Organization for Standardization, Geneva, Switzerland, 1982. ISBN ???? LCCN

???? URL http://www.iso.ch/
cate/d13802.html.

ISO:1988:IIPb

[ISO88] ISO. ISO 8651-2:1988: Information processing systems — Computer graphics — Graphical Kernel System (GKS) language bindings — Part 2: Pascal. International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? 168

pp.  LCCN ????  CHF

212. URL http://www.iso.ch/
cate/d16025.html.

ISO:1990:EPI

[ISO90a]  ISO.  Extended Pascal ISO

10206:1990. International Organization for Standardization, Geneva, Switzerland, 1990. ISBN ???? xii + 218 pp. LCCN

???? URL http://www.iso.ch/
cate/d18237.html. Available in

PostScript for personal use only at http://pascal.miningco.com/msub1.htm.

ISO:1990:IITa

[ISO90b] ISO. ISO 7185:1990: Information technology — Programming languages — Pascal. International Organization for Standardization, Geneva, Switzerland, 1990. ISBN ???? 78 pp. LCCN ???? CHF 164; US$112.00.

URL http://webstore.ansi.
org/ansidocstore/product.
asp?sku=ISO+7185%3A1990;

http://www.iso.ch/cate/d13802.
html.

ISO:1991:IIIa

[ISO91] ISO. ISO/IEC 10206:1991 Information technology — Programming languages — Extended Pascal. International Organization for Standardization, Geneva, Switzerland, 1991. ISBN ???? 214 pp. LCCN ???? CHF

236. URL http://www.iso.ch/
cate/d18237.html. Available in

English only.

ISO:1996:IIIc

[ISO96] ISO. ISO/IEC 10514-1:1996:

Information technology — Programming languages — Part 1: Modula-2, Base Language. International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 707 pp. LCCN ???? CHF 360; US$245.00. URL http://

webstore.ansi.org/ansidocstore/
product.asp?sku=ISO%2FIEC+
10514%2D1%3A1996;   http:
//www.iso.ch/cate/d18583.
html.

ISO:1998:IIIc

[ISO98a] ISO. ISO/IEC 10514-2:1998:

Information technology — Programming languages — Part 2: Generics Modula-2. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 45 pp. LCCN ???? CHF 128; US$88.00.

URL http://webstore.ansi.
org/ansidocstore/product.
asp?sku=ISO%2FIEC+10514%2D2%

                                        3A1998;   http://www.iso.
                                        ch/cate/d20792.html.

ISO:1998:IIId

[ISO98b] ISO. ISO/IEC 10514-3:1998: Information technology — Programming languages — Part 3: Object Oriented Modula-

2. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 48 pp. LCCN ???? CHF 136; US$92.00. URL

                                        http://webstore.ansi.org/
                                        ansidocstore/product.asp?
                                        sku=ISO%2FIEC+10514%2D3%3A1998;
                                         http://www.iso.ch/cate/d20793.
                                        html.

Jesshope:1985:IPE

[JCL85] C. R. Jesshope, M. J. Crawley, and G. L. Lovegrove. An intelligent Pascal editor for a graphical oriented workstation. Software— Practice and Experience, 15(11): 1103–1119, November 1985. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jensen:1973:DPP

[Jen73]  Kathleen Jensen.  [Descrip-

tion of the Pascal P-machine]. Letter, ETH Zürich, Zürich, Switzerland, January 1973. 5

                                        pp.  URL http://pascal.
                                        hansotten.com/uploads/wirth/
                                        P-Machine_Description_Jan73.
                                        pdf.

Jensen:2000:S

[Jen00] Kathleen Jensen. Serendipity. In Böszörményi [Bös00], pages 245–246. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1 (dpunktVerl.). LCCN QA76.756 .S36 2000. URL http:

//digitale-objekte.hbz-nrw.
de/storage2/2016/09/08/file_
8/6866144.pdf.

Janko:1987:SBS

[JF87] Wolfgang H. Janko and Reinhard Feurer. Eine Studie zur Beurteilung der sprachlichen Eignung der Programmiersprachen BASIC, PASCAL, APL, APL2, LISP und PRO-LOG zur Programmierung regelbasierter Systeme. (German) [A study to assess the linguistic suitability of the programming languages BASIC, PASCAL, APL, APL2, LISP and PROLOG for programming rule-based systems]. In Ökonomie und Mathematik: Rudolf Henn zum 65. Geburtstag. (German) [Economics and Mathematics: for Rudolf Henn on his 65th birthday], pages 355–364. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 3-540-17819-8.

Johnson:1984:PDF

[JM84] Marc C. Johnson and Allen Munro. Pascal’s design flaws: Modula-2 solutions and Pascal patches: A description of seven subtle problems with Pascal, and a look at how Modula-2 avoids them. Byte Magazine, 9(3):371– 372, 374, 376, 378, 380, 382, 384, 387–388, March 1984. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

                                        URL https://ia800308.us.
                                        archive.org/20/items/byte-
                                        magazine-1984-03/1984_03_
                                        BYTE_09-03_Simulation.pdf.

Jordan:1989:ECMa

[Jor89a]  M. Jordan.  Experiences in

configuration management for Modula-2. In Proceedings of the 2nd International Workshop on Software Configuration Management, SCM ’89, pages 126– 128. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-334-5.

Jordan:1989:ECMb

[Jor89b]  M. Jordan.  Experiences in

configuration management for Modula-2. ACM SIGSOFT Software Engineering Notes, 14(7): 126–128, October 1989. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jordan:1990:EPEa

[Jor90a] Mick Jordan. An extensible programming environment for Modula-3. In Proceedings of the Fourth ACM SIGSOFT Symposium on Software Development Environments, SDE 4, pages 66– 76. ACM Press, New York, NY 10036, USA, 1990. ISBN 0-89791-418-X.

Jordan:1990:EPEb

[Jor90b] Mick Jordan. An extensible programming environment for Modula-3. ACM SIGSOFT Software Engineering Notes, 15(6): 66–76, October 1990. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Joslin:1979:CAP

[Jos79] D. A. Joslin. A case for acquiring Pascal. Software—Practice and Experience, 9(8):691–692, August 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See comment [Ano80].

Joslin:1980:LES

[Jos80] D. A. Joslin. Letters to the Editor: Some defensive comments on ‘a case for acquiring Pascal’. Software—Practice and Experience, 10(7):590, July 1980. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jimenez-Peris:1996:MIVa

[JPPM96a] Ricardo Jiménez-Peris and Marta Patiño-Mart´ ınez. A Modula-2 interpreter/visualizer. In Proceedings of the 1st Conference on Integrating Technology into Computer Science Education, ITiCSE ’96, page 232. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-844-4.

Jimenez-Peris:1996:MIVb

[JPPM96b] Ricardo Jiménez-Peris and Marta Patiño-Mart´ ınez. A Modula-2 interpreter/visualizer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):232, January 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jimenez-Peris:1996:MIVc

[JPPM96c] Ricardo Jiménez-Peris and Marta Patiño-Mart´ ınez. A Modula-2 interpreter/visualizer. SIGCUE Outlook, 24(1–3):232, January 1996. ISSN 0163-5735.

Jensen:1974:PUMa

[JW74a] Kathleen Jensen and Niklaus Wirth. PASCAL: user manual and report, volume 18 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNCSD9. ISBN 0-387-06950-X. ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 170 pp. LCCN QA267.A1 L43 no.18. URL http://

link.springer-ny.com/link/
service/series/0558/tocs/
t0018.htm;   http://www.springerlink.
com/content/978-0-387-06950-
0;   http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=18.

Jensen:1974:PUMb

[JW74b] Kathleen Jensen and Niklaus Wirth. Pascal User Manual and Report. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1974. ISBN 0-387-90144-2, 3-540-90144-2, 3-540-07167-9 (print), 3-540-37500-7 (e-book). viii + 167 pp. LCCN QA76.73.P35 J461

1975.  URL http://link.
springer.com/10.1007/978-3-

540-37500-5. See also [JW85, Ame83].

Jensen:1975:PUM

[JW75] Kathleen Jensen and Niklaus Wirth. PASCAL, User Manual and Report, volume 18 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1975. CODEN LNCSD9. ISBN 0-387-07167-9 (New York), 3-540-07167-9 (Berlin), 0-387-90144-2. ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 167 pp. LCCN QA76.73.P35 .J461 1974. US$5.90. URL http:/

/link.springer-ny.com/link/
service/series/0558/tocs/
t0018.htm;   http://www.springerlink.
com/content/978-0-387-07167-
1;   http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=18.

Jensen:1985:PUM

[JW85] Kathleen Jensen and Niklaus Wirth. Pascal User Manual and Report: ISO Pascal Standard. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1985. ISBN 0-387-96048-1, 3-540-96048-1. xvi + 266 pp. LCCN QA76.73.P2 J46 1985. Revised by Andrew B. Mickel and James F. Miner. See [Ame83, JW74a, JW74b].

Jensen:1991:PUM

[JW91] Kathleen Jensen and Niklaus Wirth. Pascal User Manual and Report: ISO Pascal Standard. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 1991. ISBN 0-387-97649-3, 3-540-97649-3. xvi + 266 pp. LCCN QA76.73.P2 J46 1991. Revised by Andrew B. Mickel and James F. Miner.

Kaeding:1995:PPG

[Kae95] Thomas A. Kaeding. Pascal program for generating tables of SU(3) Clebsch–Gordan coefficients. Computer Physics Communications, 85(1):82–88, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

King:1990:DMV

[KC90] K. N. King and David A. Crick. The development of a Modula-2 validation suite. In Proceedings of the 28th Annual Southeast Regional Conference, ACM-SE 28, pages 95–100. ACM Press, New York, NY 10036, USA, 1990. ISBN 0-89791-356-6.

Kemmerer:1985:UUB

[KE85] Richard A. Kemmerer and Steven T. Eckmann. UNISEX: A UNIX-based symbolic EXecutor for Pascal. Software—Practice and Experience, 15(5):439–458, May 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Keller:2021:COO

[Kel21]  Rochus Keller.  Comparing

Oberon+ with Oberon-2 and Oberon-07. Web site, July 16,

                                        2021. URL https://oberon-
                                        lang.github.io/2021/07/16/

comparing-oberon+-with-oberon-
2-and-07.html.

Kernighan:1981:WPM

[Ker81] Brian W. Kernighan. Why Pascal is not my favorite programming language. Computing Science Technical Report 100, AT&T Bell Laboratories, Murray Hill, NJ 07974, USA, April

2, 1981. URL https://github.
com/arnoldrobbins/cstr100.

Kerridge:1982:FIC

[Ker82] Jon M. Kerridge. A FORTRAN implementation of Concurrent Pascal. Software—Practice and Experience, 12(1):45–55, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kernighan:1984:WPM

[Ker84] Brian W. Kernighan. Why Pascal is not my favorite programming language. In Feuer and Gehani [FG84a], pages 170– 186. ISBN 0-13-154840-9 (paperback), 0-13-154857-3 (hardcover). LCCN QA76.73.A35 C66 1984. See also [WSH77, KP81, Ker81].

King:1989:ISM

[Kin89] K. N. King. The international standardization of Modula-2. In Proceedings of the 17th Conference on ACM Annual Computer Science Conference, CSC ’89, pages 313–319. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-299-3.

King:1991:WNM

[Kin91] K. N. King. What’s new with Modula-2? Dr. Dobb’s Journal of Software Tools, 16(6):42, 44, 46, 48–49, June 1991. CODEN DDJOEB. ISSN 1044-789X.

King:1993:MTA

[Kin93] K. N. King. Modula-3: a threat to Ada? In Proceedings of the Tenth Annual Washington Ada Symposium on Ada: Ada’s Role in Software Engineering, WADAS ’93, pages 147– 154. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-609-3.

Kirk:1995:OGO

[Kir95] Brian Kirk. The Oakwood guidelines for Oberon-2 compiler developers. Report, Robinson Associates, Red Lion House, St Mary’s Street Painswick GLOS, GL6 6QR, UK, October 20, 1995.

49 pp. URL https://github.
com/Oleg-N-Cher/OfrontPlus/

.

Koskimies:1995:DFS

[KM95] Kai Koskimies and Hanspeter Mössenböck. Designing a framework by stepwise generalization. In Software Engineering — ESEC ’95, pages 479–498. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-45552-3. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.
com/chapter/10.1007/3-540-
60406-5_31.

Knasmuller:1995:ODU

[Kna95] Markus Knasmüller. Oberon dialogs: a user interface for end users. In GISI 95, pages 749–757. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-642-79958-2. ISSN 1431-472X.

URL http://link.springer.
com/chapter/10.1007/978-3-
642-79958-7_104.

Knasmuller:1997:APO

[Kna97] Markus Knasmüller. Adding persistence to the Oberon-System. In Modular Programming Languages, pages 97–111. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-68328-3. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.
com/chapter/10.1007/3-540-
62599-2_33.

Knuth:1990:SPW

[Knu90] Donald E. Knuth. A simple program whose proof isn’t. In Feijen et al. [FvGGM90], chapter 27, pages 233–242. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326 1990. URL http://

www.zentralblatt-math.org/
zmath/en/search/?an=0718.

68004. Contains important treatment of accurate binaryto-decimal conversion [Gri90, Knu90].

Knudsen:2000:MR

[Knu00] Svend Erik Knudsen. Medos in retrospect. In Böszörményi [Bös00], pages 69–86. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1 (dpunktVerl.). LCCN QA76.756 .S36 2000. URL http:/

                                        /digitale-objekte.hbz-nrw.
                                        de/storage2/2016/09/08/file_
                                        8/6866144.pdf.

Koffman:1988:CMCa

[Kof88a] Elliot B. Koffman. The case for Modula-2 in CS1 and CS2. In Proceedings of the Nineteenth SIGCSE Technical Symposium on Computer Science Education, SIGCSE ’88, pages 49–53. ACM Press, New York, NY 10036, USA, 1988. ISBN 0-89791-256-X.

Koffman:1988:CMCb

[Kof88b] Elliot B. Koffman. The case for Modula-2 in CS1 and CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 20(1): 49–53, February 1988. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kominek:1998:BDS

[Kom98]  John Kominek.  Building

a distributed spreadsheet in Modula-3. Linux Journal, 49: ??, May 1998. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL

                                        ftp://ftp.ssc.com/pub/lj/
                                        listings/issue49/2690.tgz.

Kornelis:2024:NWW

[Kor24] Chris Kornelis. Niklaus Wirth, who inspired a generation of computer programmers, dies at 89. Wall Street Journal Web site, February 24, 2024.

URL https://www.wsj.com/
tech/personal-tech/niklaus-
wirth-computer-programmers-
dies-at-89-6817b134.

Kernighan:1981:STP

[KP81] Brian W. Kernighan and P. J.

Plauger. Software Tools in Pascal. Addison-Wesley, Reading, MA, USA, 1981. ISBN 0-201-10342-7. ix + 366 pp. LCCN QA76.6 .K493. US$13.95.

Kredel:1997:MAS

[KP97a] Heinz Kredel and Michael Pesch.

MAS, the Modula-2 Algebra System, version 1.00. SIGSAM Bulletin, 31(3):52–54, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Kutter:1997:FSO

[KP97b] Ph. W. Kutter and A. Pierantonio. The formal specification of Oberon. J.UCS: Journal of Universal Computer Science, 3(5):443–503, May 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (elec-

tronic). URL http://medoc.
springer.de:8000/jucs_3_5/
specification_oberon;   http:
//www.jucs.org/jucs_3_5/specification_
oberon; internal&sk=05460486.

Kenner:1993:ASG

[KPUG93] Hugh Kenner, Dick Pountain, Jon Udell, and Raymond Ga Cote. Again the swinging gates: Another look at Microsoft’s leader, a controversial Windows book, how to program in Oberon, and more. Byte Magazine, 18(2):222–??, February 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kredel:1989:SDC

[Kre89] H. Kredel. Software development for computer algebra or from ALDES/SAC-2 to WEB/ Modula-2. In Gianni [Gia89], pages 447–455. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Kruijer:1982:MUO

[Kru82] H. S. M. Kruijer. A multi-user operating system for transaction processing, written in Concurrent Pascal. Software—Practice and Experience, 12(5):445–454, May 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kriz:1980:EPC

[KS80] J. Kriz and H. Sandmayr. Extension of Pascal by coroutines and its application to quasiparallel programming and simulation. Software—Practice and Experience, 10(10):773–789, October 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kerridge:1984:TSR

[KS84] Jon M. Kerridge and Dan Simpson. Three solutions for a robot arm controller using Pascal-Plus, occam and Edison. Software —Practice and Experience, 14 (1):3–15, January 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kleinmichel:1989:NTM

[KS89] Helmut Kleinmichel and Klaus Schönefeld. Newton-type methods for nonlinearly constrained programming problems. PASCALprogram and numerical experiments. Wissenschaftliche Zeitschrift der Technischen Universität Dresden, 38(3):219–224, 1989. ISSN 0043-6925.

Klaeren:1990:CMI

[KT90] Herbert Klaeren and Peter Thiemann. A clean Modula-2 interface to abstract data types. Structured Programming, 11(2): 69–78, 1990. CODEN STPGEM. ISSN 0935-1183.

Kaminski:2015:MSO

[KV15] Ted Kaminski and Eric Van Wyk. A modular specification of Oberon0 using the Silver attribute grammar system. Science of Computer Programming, 114 (??):33–44, December 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0167642315003020.

Knuth:1985:PPI

[KW85] Donald E. Knuth and Niklaus Wirth. Programming philosophy (interviews by Ken Takara). Computer Language Magazine, 2 (5):25–35, May 1985. CODEN COMLEF. ISSN 0749-2839.

Ladd:1989:CMC

[Lad89] Scott Robert Ladd. Comparing Modula-2 and C++. Dr. Dobb’s Journal of Software Tools, 14(1): 62–68, 102–104, January 1989. CODEN DDJOEB. ISSN 0888-3076.

Lalis:1991:XSD

[Lal91] Spiros Lalis. XNet: supporting distributed programming in the Oberon environment. Technical report 161, ETH Zürich, Zürich, Switzerland, May 1991. i + 31

                                        pp. URL http://hdl.handle.
                                        net/20.500.11850/68905.

Lecarme:1974:RPH

[LD74] Olivier Lecarme and Pierre Desjardins. Reply to a paper by A. N. Habermann on the programming language Pascal. ACM SIGPLAN Notices, 9(10): 21–27, October 1974. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lecarme:1975:MCP

[LD75] Olivier Lecarme and Pierre Desjardins. More comments on the programming language Pascal. Acta Informatica, 4(3):231– 243, September 31, 1975. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [Wir71d, Hab73]. An earlier version appeared in ACM SIG-PLAN Notices, October, 1975.

Lecarme:1984:MCP

[LD84] O. Lecarme and P. Desjardins.

More comments on the programming language Pascal. In Feuer and Gehani [FG84a], page ?? ISBN 0-13-154840-9 (paperback), 0-13-154857-3 (hardcover). LCCN QA76.73.A35 C66 1984.

Leap:1987:MIO

[Lea87] Thomas R. Leap. Modula-2 input/output procedure using polymorphic and open-ended data type extensions (abstract only). In Proceedings of the 15th Annual Conference on Computer Science, CSC ’87, page 381. ACM Press, New York, NY 10036, USA, 1987. ISBN 0-89791-218-7.

Lecarme:1973:ESP

[Lec73] Olivier Lecarme. An experience in structured programming and transferability. ACM SIGPLAN Notices, 8(9):95–96, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lecarme:1974:SPP

[Lec74] Olivier Lecarme. Structured programming, programming teaching and the language Pascal. ACM SIGPLAN Notices, 9(7): 15–21, July 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ledgard:1984:APSc

[Led84] Henry F. Ledgard. The American Pascal Standard: with annotations. Springer books on professional computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 0-387-91248-7, 3-540-91248-7. 97 pp. LCCN ????

Lees:1985:ICP

[Lee85] B. Lees. Introductory concurrent programming with Modula-2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 17(3): 34–41, September 1985. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Levy:1988:LLC

[Lev88] E. Levy. The linked list class of Modula-3. ACM SIGPLAN Notices, 23(8):93–102, August 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leblanc:1982:CSR

[LF82] Richard J. Leblanc and Charles N. Fischer. A case study of run-time errors in Pascal programs. Software—Practice and Experience, 12(9):825–834, September 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

LEcuyer:1987:POS

[LG87] Pierre L’Ecuyer and Nataly Giroux. A process-oriented simulation package based on Modula-2. In Proceedings of the 19th Conference on Winter Simulation, WSC ’87, pages 165–174. ACM Press, New York, NY 10036, USA, 1987. ISBN 0-911801-32-4.

Ledgard:1979:PS

[LHN79] Henry F. Ledgard, John F.

Hueras, and Paul A. Nagin. Pascal with Style. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1979. ISBN 0-8104-5124-7. 210 pp. LCCN QA76.73.P2 L4.

Li:1988:IFO

[Li88] Wei Li. Implementing first order logic in Modula-2 using an intuitionistic approach. In Proceedings of the 1988 ACM Sixteenth Annual Conference on Computer Science, CSC ’88, pages 27– 36. ACM Press, New York, NY 10036, USA, 1988. ISBN 0-89791-260-8.

Liddiard:1976:YAL

[Lid76] Lawrence A. Liddiard. Yet another look at code generation for Pascal on CDC 6000 & Cyber machines. Technical Report UCC 76002, University Computer Center, University of Minnesota, Minneapolis, MN 55455, USA, December 1976. i + 14 pp.

URL https://conservancy.
umn.edu/bitstream/handle/

11299/162946/UCC_Technical_
Report_76002_Dec_1976.pdf.

Lins:1990:PEO

[Lin90] C. Lins. Programming without enumerations in Oberon. ACM SIGPLAN Notices, 25(7):19–27, July 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:1986:SPM

[Liu86] Ken-Chih Liu. A string pattern matching extension to Pascal and some comparisons with SNOBOL4. Software—Practice and Experience, 16(6):541–548, June 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Learner:1970:NTC

[LL70] A. Learner and A. L. Lim. A note on transforming contextfree grammars to Wirth–Weber precedence form. The Computer Journal, 13(2):142–144, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/
content/13/2/142.full.pdf+
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_02/130142.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_02/tiff/142.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/143.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_13/Issue_
02/tiff/144.tif.

Luthi:2024:CPN

[LM24] Pauline Lüthi and Florian Meyer. Computer pioneer Niklaus Wirth has died. ETH Web site, July 4, 2024. URL

https://ethz.ch/en/news-and-
events/eth-news/news/2024/
01/computer-pioneer-niklaus-
wirth-has-died.html.

Leino:1998:ESC

[LN98] K. R. M. Leino and G. Nelson. An extended static checker for Modula-3. Lecture Notes in Computer Science, 1383:302– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lo:1985:SPM

[Lo85]  Chingmin Jim Lo.  Simple

patches to Modula-2 text IO. ACM SIGPLAN Notices, 20 (6):20–25, June 1985. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Louden:1990:PCC

[Lou90] Kenneth Louden. P-code and compiler portability: Experience with a Modula-2 optimizing compiler. ACM SIGPLAN Notices, 25(5):53–59, May 1990. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lalonde:1983:STC

[LP83] Wilf R. Lalonde and John R.

Pugh. A simple technique for converting from a PASCAL shop to a C shop. Software—Practice and Experience, 13(9):771– 775, September 1983. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Luckham:1979:VAR

[LS79] David C. Luckham and Norihisa Suzuki. Verification of array, record, and pointer operations in Pascal. ACM Transactions on Programming Languages and Systems, 1(2):226–244, October 1979. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Leeson:1987:TIM

[LS87] J. J. Leeson and M. L. Spear. Type independent modules: The preferred approach to generic ADTs in Modula-2. ACM SIGPLAN Notices, 22(3):65–70, March 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lalis:1994:ACO

[LS94] Spiros Lalis and Beverly A. Sanders. Adding concurrency to the Oberon system. Lecture Notes in Computer Science, 782: 328–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lightfoot:2006:MPL

[LS06] David E. Lightfoot and Clemens Szyperski, editors. Modular Programming Languages: 7th Joint Modular Languages Conference, JMLC 2006 Oxford, UK, September 13–15, 2006 Proceedings, volume 4228 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CO-DEN LNCSD9. ISBN 3-540-40927-0 (print), 3-540-40928-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.
springerlink.com/content/
978-3-540-40928-1.

Luthi:2021:ASM

[Lüt21]  Pauline Lüthi.  “I always

saw myself as an engineer”

— part 1/3. Web video interviews with Niklaus Wirth, December 3, 2021. URL

https://inf.ethz.ch/news-
and-events/spotlights/infk-
news-channel/2021/11/niklaus-
wirth-video-interview.html.

Machura:1979:ISP

[Mac79] Marek Machura. Implementation of a special-purpose language using Pascal implementation methodology. Software— Practice and Experience, 9(11): 931–945, November 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mahjoub:1980:NMC

[Mah80] Ahmed Mahjoub. A new Modula compiler for the LSI-11. ACM SIGPLAN Notices, 15(6):39–45, June 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Malcolm:1971:PRP

[Mal71] Michael A. Malcolm. PL360 (revised) a programming language for the IBM 360. Technical Report STAN-CS-71-215, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1971. v + 94 pp. URL http:/

                                        /infolab.stanford.edu/pub/
                                        cstr/reports/cs/tr/71/215/
                                        CS-TR-71-215.pdf.

Malcolm:1972:PRA

[Mal72] Michael A. Malcolm. PL360 (revised again) a programming language for the IBM 360. Technical Report STAN-CS-71-215, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1972. vi + 103 pp. Original version May 1971.

Marmier:1975:AVP

[Mar75] Édouard Marmier. Automatic Verification of Pascal Programs. Ph.D. thesis, ETH Zürich, Zürich, Switzerland, 1975. v + 138 + 58 pp. ETH Dissertation Number 5629.

Marlin:1979:HBI

[Mar79] Chris D. Marlin. A heap-based implementation of the programming language Pascal. Software —Practice and Experience, 9(2): 101–119, February 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marsden:1984:SPE

[Mar84] B. W. Marsden. A STandard Pascal event simulation package. Software—Practice and Experience, 14(7):659–684, July 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Marais:1994:OS

[Mar94] Johannes L. Marais. Oberon System 3. Dr. Dobb’s Journal of Software Tools, 19(11): 42–50, October 1994. CODEN DDJOEB. ISSN 1044-789X.

Marais:1996:ESS

[Mar96] Johannes L. Marais. Extensible software systems in Oberon. Journal of Computational and Graphical Statistics, 5(3):284– 298, 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1080/10618600.1996.
10474712.

Martynenko:2014:TAW

[Mar14] Boris Martynenko. Towards the 80th anniversary of N. Wirth: Wirth’s syntactic charts in the SYNTAX-technology. In 2014 Third International Conference on Computer Technology in Russia and in the Former Soviet Union, pages 199–206. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2014.

Mastorakis:1998:BSG

[Mas98] Nikos E. Mastorakis. Bookshelf: Stylish guide to Modula-3. IEEE Software, 15(4):89–90, July/ August 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://

                                        dlib.computer.org/so/books/
                                        so1998/pdf/s4087.pdf.

Mattsson:1980:ICP

[Mat80] Sven Erik Mattsson. Implementation of concurrent Pascal on LSI-11. Software—Practice and Experience, 10(3):205–217, March 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See remarks [Hay80].

Mateti:1984:PVC

[Mat84] Prabhaker Mateti. Pascal versus C: a subjective comparison. In Feuer and Gehani [FG84a], page ?? ISBN 0-13-154840-9 (paperback), 0-13-154857-3 (hardcover). LCCN QA76.73.A35 C66 1984.

McCaig:1983:FPP

[McC83] J. M. McCaig. FUMBLR: a Pascal program for initializing memory. Software—Practice and Experience, 13(12):1191–1200, December 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Meek:1987:BRB

[Mee87]  Brian Meek.  Book review:

A Guide to Modula-2, Kaare Christian, Springer-Verlag, 1986.

ISBN 0-387-9642–5. No. of pages: 436 + xix. Price: £33.00 (hardback). Software—Practice and Experience, 17(2):167– 169, February 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Meyer:1999:CLP

[Mey99] Uwe Meyer. Correctness of on-line partial evaluation for a Pascal-like language. Science of Computer Programming, 34(1): 55–73, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Meyer:2024:NWD

[Mey24] Bertrand Meyer. Niklaus Wirth has died. Web site, January 4, 2024. URL https:/

/news.ycombinator.com/item?
id=38858012.

Morales-Fernandez:1989:CMM

[MFMN89] R. Morales-Fernandez and J. J.

Moreno-Navarro. CC-Modula: a Modula-2 tool to teach concurrent programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 21(3):19–25, September 1989. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Matsumoto:1996:AAP

[MHT96] Aki Matsumoto, D. S. Han, and Takao Tsuda. Alias analysis of pointers in Pascal and Fortran 90: Dependence analysis between pointer references. Acta Informatica, 33(2):99–130, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:/

/link.springer-ny.com/link/
service/journals/00236/bibs/
6033002/60330099.htm.

Milanese:1989:AOS

[Mil89]  V. Milanese.  Aspects of

OBERON suitable for graphics applications. ACM SIG-PLAN Notices, 24(8):37–45, August 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Muhlbacher:1997:IO

[MLKK97a] Jörg R. Mühlbacher, Bernhard Leisch, Brian Kirk, and Ulrich Kreuzeder. Introduction to Oberon-2. In Oberon-2 Programming with Windows [MLKK97b], pages 49–174. ISBN 3-540-62522-4, 3-642-45762-2 (e-book). LCCN QA76.76.O63 O24 1997.

URL http://link.springer.
com/chapter/10.1007/978-3-
642-45762-3_4.

Muhlbacher:1997:OPW

[MLKK97b] Jörg R. Mühlbacher, Bernhard Leisch, Brian Kirk, and Ulrich Kreuzeder. Oberon-2 Programming with Windows. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-62522-4, 3-642-45762-2 (ebook). ix + 370 + 66 pp. LCCN QA76.76.O63 O24 1997.

URL http://link.springer.
com/chapter/10.1007/978-3-
642-45762-3.

Muhlbacher:1997:WEP

[MLKK97c] Jörg R. Mühlbacher, Bernhard Leisch, Brian Kirk, and Ulrich Kreuzeder. The working environment of POW! In Oberon-2 Programming with Windows [MLKK97b], pages 15–47. ISBN 3-540-62522-4, 3-642-45762-2 (ebook). LCCN QA76.76.O63 O24 1997. URL http://

link.springer.com/chapter/
10.1007/978-3-642-45762-3_
3.

Moffat:1984:SCA

[Mof84] David V. Moffat. Some concerns about Modula-2. ACM SIGPLAN Notices, 19(12):41– 47, December 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See comments [SS85, Gol85].

Mohilner:1977:UPF

[Moh77] Patricia R. Mohilner. Using Pascal in a Fortran environment. Software—Practice and Experience, 7(3):357–362, May/ June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mohr:1989:UMT

[Moh89] J. Mohr. Using Modula-2 to teach data abstraction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 21(3):60–64, September 1989. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mossenbock:1991:DBOa

[Mös91] Hanspeter Mössenböck. Differences between Oberon and Oberon-2: the programming language Oberon-2. Technical report 160, ETH Zürich, Zürich, Switzerland, May 1991. 32 pp.

URL http://hdl.handle.net/
20.500.11850/68904.

Mossenbock:1993:O

[Mös93a] Hanspeter Mössenböck. Oberon-2. In Object-Oriented Programming, pages 13–27. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-642-97479-1. URL http://

link.springer.com/chapter/
10.1007/978-3-642-97479-3_
2.

Mossenbock:1993:OOP

[Mös93b] Hanspeter Mössenböck. Object-Oriented Programming in Oberon 2. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-56411-X (New York), 3-540-56411-X (Berlin), 3-642-97479-1 (e-book). xiii + 278 pp. LCCN QA76.64 .M67 1993.

Mossenbock:1993:OPO

[Mös93c] Hanspeter Mössenböck. Objektorientierte Programmierung in Oberon-2. (German) [Object Oriented Programming in Oberon-2]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1993. ISBN 3-540-55690-7, 3-540-57789-0. xii + 286 pp. LCCN QA76.6-

76.66.  URL http://www.
zentralblattmath.org/zmath/
en/search/?an=0835.68011.

Mossenbock:1994:EOS

[Mös94] Hanspeter Mössenböck. Extensibility in the Oberon system. Nordic Journal of Computing, 1(1):77–93, Spring 1994. CO-DEN NJCOFR. ISSN 1236-

6064. URL http://www.cs.
helsinki.fi/njc/References/
mossenbock1994:77.html.

Mosli:1995:CCF

[Mös95a] Bernd Mösli. A comparison of C++, FORTRAN 90 and Oberon-2 for scientific programming. In GISI 95, pages 740–748. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-642-79958-2. ISSN 1431-472X.

URL http://link.springer.
com/chapter/10.1007/978-3-
642-79958-7_103.

Mossenbock:1995:O

[Mös95b] Hanspeter Mössenböck. Oberon-2. In Object-Oriented Programming in Oberon-2, pages 13– 27. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-642-79898-5. URL http:/

/link.springer.com/chapter/
10.1007/978-3-642-79898-6_
2.

Mossenbock:1995:OOP

[Mös95c] Hanspeter Mössenböck. Object-Oriented Programming in Oberon-2. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1995. ISBN 3-540-60062-0 (New York), 3-540-60062-0 (Berlin), 3-642-79898-5 (e-book, Berlin). xiii + ii + iii + 278 pp. LCCN QA76.64. URL https:/

/ssw.jku.at/Research/Books/
Oberon2.pdf.

Mossenbock:1997:MPL

[Mos97] Hanspeter Mossenbock, editor. Modular programming languages: Joint Modular Languages Conference, JML ’97, Linz, Austria, March 19–21, 1997: proceedings, volume 1204 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNCSD9. ISBN 3-540-62599-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .J6578 1997. URL http://

link.springer-ny.com/link/
service/series/0558/tocs/
t1204.htm;   http://www.springerlink.
com/content/978-3-540-62599-
5;   http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=1204.

Mossenbock:1998:OPOd

[Mös98]  H. Mössenböck.  Objekto-

rientierte Programmierung in Oberon-2. (German) [Object Oriented Programming in Oberon-2]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1998. ISBN 3-540-64649-3. xvii + 329 pp. LCCN QA76.64 .M67 1998.

Mossenbock:2000:CCA

[Mös00] Hanspeter Mössenböck. Compiler construction: the art of Niklaus Wirth. In Böszörményi [Bös00], pages 55–68. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1 (dpunktVerl.). LCCN QA76.756 .S36 2000. URL http:/

/digitale-objekte.hbz-nrw.
de/storage2/2016/09/08/file_
8/6866144.pdf.

Madhavji:1986:MML

[MPT86] Nazim H. Madhavji, Luc Pinsonneault, and Kamel Toubache. Modula-2/MUPE-2: Language and environment interactions. IEEE Software, 3(6):7–17, November 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Murali:1983:SGC

[MS83] V. Murali and R. K. Shyamasundar. A sentence generator for a compiler for PT, a PASCAL subset. Software—Practice and Experience, 13(9):857–869, September 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McLarty:1984:IMP

[MS84] Hugh McLarty and David W.

Smith. An introduction to Modula-2 for Pascal programmers. Dr. Dobb’s Journal of Software Tools, 9(5):22–24, 26–27, May 1984. CODEN DDJOEB. ISSN 1044-789X.

McCracken:1987:SCC

[MS87] Daniel D. McCracken and William I. Salmon. A Second Course in Computer Science with Modula-2. Wiley, New York, NY, USA, 1987. ISBN 0-471-63111-6. xiii + 464 + 10 pp. LCCN QA76.73.M63 M43 1987.

Malloy:1990:CSP

[MS90] Brian Malloy and Mary Lou Soffa. Conversion of simulation processes to Pascal constructs. Software—Practice and Experience, 20(2):191–207, February 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Moessenboek:1999:ORM

[MS99a] H. Moessenboek and C. Steindl. The Oberon-2 reflection model and its applications. Lecture Notes in Computer Science, 1616:40–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mossenbock:1999:ORM

[MS99b] Hanspeter Mössenböck and Christoph Steindl. The Oberon-2 reflection model and its applications. In Meta-Level Architectures and Reflection, volume 1616 of Lecture Notes in Computer Science, pages 40–53. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNCSD9. ISBN 3-540-48443-4. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
1616/16160040.htm;   http:/
/link.springer-ny.com/link/
service/series/0558/papers/
1616/16160040.pdf;   http:/
/link.springer.com/chapter/
10.1007/3-540-48443-4_4.

Mossenbock:1989:OOO

[MTG89] Hanspeter Mössenböck, Josef Templ, and Robert Griesemer. Object Oberon: an object-oriented extension of Oberon. Technical Report 109, Department Informatik, ETH Zürich, Zürich, Switzerland, June 1989. i + 30 pp. URL

http://hdl.handle.net/20.
500.11850/68697;   https:/
/www.research-collection.
ethz.ch/bitstream/handle/
20.500.11850/68697/eth-3204-
01.pdf.

Magnenat-Thalmann:1981:GPE

[MTT81] Nadia Magnenat-Thalmann and Daniel Thalmann. A graphical Pascal extension based on graphical types. Software— Practice and Experience, 11(1): 53–62, January 1981. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Magnenat-Thalmann:1983:MTD

[MTT83] Nadia Magnenat-Thalmann and Daniel Thalmann. MIRA-3D: a three-dimensional graphical extension of Pascal. Software— Practice and Experience, 13(9): 797–808, September 1983. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Muller:1984:DBM

[Mul84] Hausi A. Muller. Differences between Modula-2 and Pascal. ACM SIGPLAN Notices, 19(10): 32–39, October 1984. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Muller:1986:MPS

[Mul86] Carlo Muller. Modula-Prolog: a software development tool. IEEE Software, 3(6):39–45, November 1986. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mulley:2024:GMC

[Mul24] Gaius Mulley. The GNU Modula-2 Compiler. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 2024. v + 429

                                        pp. URL https://gcc.gnu.
                                        org/onlinedocs/gm2.pdf.

Mossenbock:1991:PLO

[MW91a] H. Mössenböck and N. Wirth. The programming language Oberon-2. Structured Programming, 12(4):179–195, ???? 1991. CODEN STPGEM. ISSN 0935-1183.

Mossenbock:1991:DBOb

[MW91b] Hanspeter Mössenböck and Niklaus Wirth. Differences between Oberon and Oberon-2. Structured Programming, 12 (4):175–178, 1991. CODEN STPGEM. ISSN 0935-1183.

Mossenbock:1993:PLO

[MW93] H. Mössenböck and N. Wirth.

The programming language Oberon-2. Web site and PDF documentation, October 1993. URL https:

//cseweb.ucsd.edu/classes/
fa00/cse131a/oberon2.htm;
https://cseweb.ucsd.edu/~
wgg/CSE131B/oberon2.htm;
https://github.com/shitikanth/
Oberon2-Compiler/blob/master/
Oberon2-Report.pdf.

Nagel:1976:DP

[Nag76] H.-H. Nagel. On DECSystem-10 Pascal. Mitteilung 40, Institute für Informatik, Universität Hamburg, Schlüterstrasse 70, D-2000 Hamburg 13, West Germany, September 9, 1976. URL http:

//pascal.hansotten.com/px-
descendants/dec-pdp-10-pascal-
compiler/.

Nori:1974:PPC

[NAJN74] K. V. Nori, Urs Ammann, Kathleen Jensen, and H. H. Nägeli. The PASCAL (P) compiler: Implementation notes. Report 10, ETH Zürich, Zürich, Switzerland, December 1974. vi +

57 pp. URL http://pascal.
hansotten.com/uploads/wirth/
Pascal-P-Compiler-Implementation-
Notes-eth-3054-01.pdf;   https:
//www.research-collection.
ethz.ch/bitstream/handle/
20.500.11850/68666/eth-3054-
01.pdf.

Nori:1976:PPC

[NAJN76] K. V. Nori, Urs Ammann, Kathleen Jensen, and H. H. Nägeli. The PASCAL (P) compiler: Implementation notes. Report, ETH Zürich, Zürich, Switzerland, ???? 1976. ???? pp. TO DO: Find online, and check authors: cited in [GW78, p. 26]. Is this a mis-citation of [NAJN74]?

Nori:1981:PPC

[NAJN81] K. V. Nori, Urs Ammann, Kathleen Jensen, and H. H. Nägeli. The Pascal P-code compiler: Implementation notes. In Barron [Bar78], pages 125–170. ISBN 0-471-27835-1 (hardcover). LCCN QA76.73.P2 P2.

Nance:1989:TM

[Nan89] B. Nance. TopSpeed Modula-2. Byte Magazine, 14(5):211– 214, May 1989. CODEN BYT-EDJ. ISSN 0360-5280.

Naumann:1994:EO

[Nau94] David A. Naumann. On the essence of Oberon. Lecture Notes in Computer Science, 782:313– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nelson:1979:CPI

[Nel79] Philip A. Nelson. A comparison of PASCAL intermediate languages. In ????, pages 208– 213. ACM Press, New York, NY 10036, USA, 1979. Describes efficient translation of P-Codes for the Cray-1 vector supercomputer Los Alamos National Laboratory, and their use in a compiler for the MODEL language. The closing paragraph states: “Although these versions are not the only versions in existence, duplicate effort will result if new users of P-code do not refer to this or other earlier work. These two versions do show the major areas of modification and future work with P-code would do well to look at this older work before reinventing ‘new’ P-code constructs.”.

Nelson:1991:SPM

[Nel91] Greg Nelson. Systems programming with Modula-3, volume 1991 of Prentice Hall series in innovative technology. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1991. ISBN 0-13-590464-1 (paperback). ix + 267 pp. LCCN QA76.66 .S87 1991.

URL https://archive.org/
details/systems-programming-
with-modula-3.

Nielsen:1988:ESD

[Nie88]  Jan Højlund Nielsen.  Em-

bedded systems development using Modula-2. Computer Physics Communications, 50 (1–2):59–70, July 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0010465588901166.

Nigro:1993:TEO

[Nig93] Libero Nigro. On the type extensions of Oberon-2. ACM SIGPLAN Notices, 28(2):41–44, February 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nikitin:1997:ROI

[Nik97]  Eric W. Nikitin.  Into the

Realm of Oberon: an Introduction to Programming and the Oberon-2 Programming Language. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-98279-5, 1-4612-2300-8 (e-book), 1-4612-7491-5 (print). xii + 197 pp. LCCN QA76.64 .N55 1998.

Nouspikel:2005:TTP

[Nou05] Thierry Nouspikel. The TI-99/ 4A tech pages. Web site, 2005.

                                        URL https://www.unige.ch/
                                        medecine/nouspikel/ti99/titechpages.

htm. See section The P-Code card, Part II: the P-system for a description of the Pascal P-Code language.

Narayana:1979:SAC

[NPJ79] K. T. Narayana, V. R. Prasad, and M. Joseph. Some aspects of concurrent programming in CCNPASCAL. Software—Practice and Experience, 9(9):749– 770, September 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Neal:1978:EPC

[NW78] David Neal and Virgil Wallentine. Experiences with the portability of concurrent PAS-CAL. Software—Practice and Experience, 8(3):341–353, May/ June 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nordstrom:1984:DIP

[NW84] Bengt Nordström and Åke Wikström. The design of an interactive program development system for Pascal. Software— Practice and Experience, 14(2): 177–190, February 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Odersky:1989:MNA

[Ode89] Martin Odersky. MINOS: a new approach to the design of an input/output library for Modula-2. Structured Programming, 10(2):89–105, 1989. CODEN STPGEM. ISSN 0935-1183.

Odersky:2000:PFN

[Ode00]  Martin Odersky.  Program-

ming with functional nets. In Böszörményi [Bös00], pages 172– 202. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1 (dpunktVerl.). LCCN QA76.756 .S36 2000. URL http:/

/digitale-objekte.hbz-nrw.
de/storage2/2016/09/08/file_
8/6866144.pdf.

Ohran:1984:LM

[Ohr84]  Richard Ohran.  Lilith and

Modula-2. Byte Magazine, 9(8): 181–192, August 1984.

Olderog:1984:CPP

[Old84] Ernst-Rüdiger Olderog. Correctness of programs with Pascal-like procedures without global variables. Theoretical Computer Science, 30(1):49–90, 1984. CODEN TCSCDI. ISSN 0304-3975,1879-2294.

Oldehoeft:1981:IMP

[ORT81] R. R. Oldehoeft, W. D. Ralph, and M. H. Tindall. An interactive manager for Pascal software. Software—Practice and Experience, 11(8):867–873, August 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Oliveira:1983:AMI

[OW83] J. N. Oliveira and I. R. Wilson. An analysis of microcomputer implementation of Pascal. Software—Practice and Experience, 13(4):373–384, April 1983. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Padiou:1990:NCE

[Pad90] Gérard Padiou. Nested coroutines for exception handling in Modula-2. Structured Programming, 11(2):79–84, 1990. CO-DEN STPGEM. ISSN 0935-1183.

Partridge:1985:SIT

[Par85] Derek Partridge. Specifications and an implementation of the type-ambiguity problem in Pascal. Software—Practice and Experience, 15(12):1141–1158, December 1985. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Pase:1985:SPM

[Pas85] Douglas M. Pase. System programming in Modula-2. ACM SIGPLAN Notices, 20(11):49– 53, November 1985. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Perrott:1981:EFP

[PD81] R. H. Perrott and P. S. Dhillon.

An experiment with Fortran and Pascal. Software—Practice and Experience, 11(5):491–496, May 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pemberton:2002:PI

[PD02] Steven Pemberton and Martin Daniels. Pascal implementation. Web site, February 2002. URL

https://homepages.cwi.nl/~
steven/pascal/book/0intro.
html.

Pescio:1997:FWN

[Pes97] Carlo Pescio. A few words with Niklaus Wirth. Software Development, 5(6):??, June

1997. URL http://pascal.
hansotten.com/uploads/wirth/
interview%20wirth%20June%
201997%20carlio%20pescio.
txt.

Peterson:1982:SFI

[Pet82] James L. Peterson. Surveyor’s forum: Idiomatic programming. ACM Computing Surveys, 14(4):

625, December 1982. CODEN CMSVAN. ISSN 0010-4892. See [FG82a, FG82b].

Pfister:2001:WNC

[Pfi01]  Cuno Pfister.  What’s new

in Component Pascal? Report, Oberon microsystems, Inc., Zürich, Switzerland, March 2001. 13 pp. URL https:

//github.com/Oleg-N-Cher/
OfrontPlus/.

Anonymous:2004:EOC

[Pfi04] Cuno Pfister. The evolution of Oberon-2 to Component Pascal. Report, Oberon microsystems, Inc., Zürich, Switzerland, February 1, 2004. URL http:

//oberon2005.oberoncore.ru/
paper/p_o2cp.pdf;   http:/
/www.oberon.ch/resources/
component_pascal/language_
report.html;   https://web.
archive.org/web/20220331230135/
http://www.oberon.ch/resources/
component_pascal/language_
report.html.

Pfister:1991:OTN

[PHT91] Cuno Pfister, Beat Heeb, and Josef Templ. Oberon technical notes. Report 156, ETH Zürich, Zürich, Switzerland, March 1991. ???? pp. URL ?

???.

Pirklbauer:2019:ROP

[Pir19] Andreas Pirklbauer. The Revised Oberon-2 programming language. Report, ????, ????, December 1, 2019. 7 pp. URL

https://github.com/Oleg-N-
Cher/OfrontPlus/.

Pomberger:2000:NWP

[PMR00] Gustav Pomberger, Hanspeter Mössenböck, and Peter Rechenberg. Niklaus Wirth: a pioneer of computer science. In Böszörményi [Bös00], pages 5– 20. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1 (dpunktVerl.). LCCN QA76.756 .S36 2000. URL http:/

/digitale-objekte.hbz-nrw.
de/storage2/2016/09/08/file_
8/6866144.pdf.

Pomberger:1985:LMW

[Pom85] Gustav Pomberger. Lilith und Modula-2: Werkzeuge der Softwaretechnik. (German) [Lilith and Modula-2: Tools of Software Engineering]. PC professionell. Hanser, München, West Germany, 1985. ISBN 3-446-14328-9 (paperback). 198 pp. LCCN ????

Pomberger:1993:SEW

[Pom93] Gustav Pomberger. Software engineering — Auf dem Wege vom Handwerk zur industriellen Fertigung von Softwareprodukten? (German) [Software engineering — on the way from craftsmanship to industrial production of software products?]. In Informationswirtschaft, pages 289–307. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-642-87094-5. URL http:/

/link.springer.com/chapter/
10.1007/978-3-642-87094-1_

Porter:1988:SAM

[Por88a] K. Porter. State of the art in Modula-2 [review of four compilers]. Dr. Dobb’s Journal of Software Tools, 13(9):64–81, 100– 106, September 1988. CODEN DDJOEB. ISSN 0888-3076.

Porter:1988:SPI

[Por88b]  Kent Porter.  Structured

programming: Implementing Wirth’s LineDrawing module. Dr. Dobb’s Journal of Software Tools, 13(4):118, 121–122, 124, 127, 128, April 1988. CODEN DDJOEB. ISSN 1044-789X.

Porter:1988:WIM

[Por88c] Kent Porter. What is Modula-2? Dr. Dobb’s Journal of Software Tools, 13(??):??, ?? 1988. CODEN DDJOEB. ISSN 1044-789X.

Posch:2024:RNW

[Pos24]  Maya Posch.  Remembering

Niklaus Wirth: Father of Pascal and inspiration to many. Web site, January 5, 2024. URL

                                        https://hackaday.com/2024/
                                        01/05/remembering-niklaus-
                                        wirth-father-of-pascal-and-
                                        inspiration-to-many/.

Pournelle:1990:MVJ

[Pou90] Jerry Pournelle. Multimedia video: Jerry looks at multimedia video boards, a new Modula-2, and assorted gadgets. Byte Magazine, 15(12):73–??, November 1990. CODEN BYTEDJ. ISSN 0360-5280.

Pountain:1991:OLA

[Pou91] D. Pountain. Oberon: a look at the latest language to come out of Niklaus Wirth’s workshop. Byte Magazine, 16(3):135–136, 138, 140, 142, March 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1993:OGA

[Pou93] Dick Pountain. Oberon: a glimpse at the future: a radically object-oriented design previews future operating systems. Byte Magazine, 18(6):111–??, May 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1995:OFS

[Pou95] Dick Pountain. The Oberon/F system: Niklaus Wirth’s Oberon/F is a lightweight, portable, objectoriented component framework for the Windows and Mac environments. Byte Magazine, 20 (1):227–??, January 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Powell:1984:POC

[Pow84a] Michael L. Powell. A portable optimizing compiler for Modula-2. ACM SIGPLAN Notices, 19 (6):310–319, June 1984. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Powell:1984:POCa

[Pow84b] Michael L. Powell. A portable optimizing compiler for Modula-2. In Proceedings of the 1984 SIGPLAN Symposium on Compiler Construction, SIGPLAN ’84, pages 310–318. ACM Press, New York, NY 10036, USA, 1984. ISBN 0-89791-139-3.

Powell:1984:UMS

[Pow84c] Michael L. Powell. Using Modula-2 for system programming with UNIX. In Software Tools Users Group [Sof84], pages 119–132. LCCN QA76.8.U65 U83 1984.

Pronk:1992:STC

[Pro92] C. Pronk. Stress testing of compilers for Modula-2. Software— Practice and Experience, 22(10): 885–897, October 1992. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Provem:2024:RSD

[Pro24] Liam Provem. RIP: Software design pioneer and Pascal creator Niklaus Wirth. Web site, January 4, 2024. URL https:

//www.theregister.com/2024/
01/04/niklaus_wirth_obituary/

. Includes 1h42m video interview with Niklaus Wirth about his academic career and professional life.

Pountain:1994:ESS

[PS94] Dick Pountain and Clemens Szyperski. Extensible software systems: New programming tools are needed to develop software systems tha can be easily extended with new modules. Byte Magazine, 19(5):57– ??, May 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pronk:1996:IIS

[PS96] C. Pronk and M. Schönhacker.

ISO/IEC 10514-1, the Standard for Modula-2: Process aspects. ACM SIGPLAN Notices, 31(8): 74–83, August 1996. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pronk:1997:SEM

[PSSW97] C. Pronk, R. J. Sutcliffe, M. Schönhacker, and A. Wiedemann. Standardized extensions to Modula-2. ACM SIGPLAN Notices, 32(11):34–48, November 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Philippsen:1992:MC

[PT92] Michael Philippsen and Walter F. Tichy. MODULA-2 and its compilation. Lecture Notes in Computer Science, 591:169– 183, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL https:

//core.ac.uk/download/pdf/

197597056.pdf. In Proceedings of the First International Conference of the Austrian Center for Parallel Computation, Salzburg, Austria, 1991.

Perkins:1984:UPV

[PV84] Daniel R. Perkins and Dennis Volper. UCSD Pascal on the VAX, portability and performance. Software—Practice and Experience, 14(5):473–482, May 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Quandt:2024:NWT

[Qua24] Roland Quandt. Niklaus Wirth ist tot: IT-Welt trauert um Erfinder von Pascal. (German) [Niklaus Wirth is dead: IT world mourns Pascal’s inventor]. Web site, January 3, 2024. URL

https://winfuture.de/news,
140430.html.

Radenski:1993:VO

[Rad93] Atanas Radenski. A voyage to Oberon. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(3): 13–18, September 1993. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Radenski:1994:OSP

[Rad94] Atanas Radenski. Is Oberon as simple as possible? A smaller object-oriented language based on the concept of module type. In Programming Languages and System Architectures, volume 782 of Lecture Notes in Computer Science, pages 298–312. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNCSD9. ISBN 3-540-48356-X. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.
com/chapter/10.1007/3-540-
57840-4_38.

Rand:1986:PPI

[Ran86] D. W. Rand. PASCAL programs for identification of Lie algebras: Part 1. Radical — a program to calculate the radical and nil radical of parameter-free and parameter-dependent Lie algebras. Computer Physics Communications, 41(1):105–125, July 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/001046558690024X.

See erratum [Ran87a].

Rand:1987:ENP

[Ran87a] D. W. Rand. Erratum notice:

“PASCAL programs for identification of Lie algebras. I. RAD-ICAL — a program to calculate the radical and nil radical of parameter-free and parameterdependent Lie algebras” [Comput. Phys. Comm. 41 (1986), no. 1, 105–125]. Computer Physics Communications, 47(2–3):369– 371, November/December 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0010465587901251.

See [Ran86].

Rand:1987:PPIb

[Ran87b] D. W. Rand. PASCAL programs for identification of Lie algebras. III. Levi decomposition and canonical basis. Computer Physics Communications, 46(2):311–322, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Ravn:1982:PVC

[Rav82] A. P. Ravn. Pointer variables in Concurrent Pascal. Software —Practice and Experience, 12 (3):211–222, March 1982. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rudnicki:1985:PPP

[RD85] Piotr Rudnicki and Włodzimierz Drabent. Proving properties of Pascal programs in MIZAR 2. Acta Informatica, 22(3):311–331, August 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rudnicki:1986:PPP

[RD86] Piotr Rudnicki and Włodzimierz Drabent. Proving properties of Pascal programs in MIZAR 2. Acta Informatica, 22(6):699–707, March 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rees:1984:BRP

[Ree84] M. J. Rees. Book review: Pascal Compiler Validation, B. A. Wichmann and Z. J. Ciechanowicz (eds.), Wiley, Chichester, 1983. No. of pages: 176. Price £9.75 Hardback. Software— Practice and Experience, 14(6): 601–602, June 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reeb:1991:TPP

[Ree91] Georges Reeb. A Turbo Pascal program. Annals of Global Analysis and Geometry, 9(1):97, 1991. ISSN 0232-704X (print), 1572-9060 (electronic).

Reed:2000:BYO

[Ree00] Paul Reed. Building your own tools: an Oberon industrial casestudy. In Gutknecht and Weck [GW00], pages 291–298. CODEN LNCSD9. ISBN 3-540-67958-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .J6578 2000. URL http://

link.springer-ny.com/link/
service/series/0558/tocs/
t1897.htm;   http://www.springerlink.
com/content/978-3-540-67958-
5;   http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=1897.

Rehmer:1987:EHU

[Reh87] K. Rehmer. Error handling using Modula-2. ACM SIGPLAN Notices, 22(3):40–48, March 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reimer:1984:IDP

[Rei84] Manuel Reimer. Implementation of the database programming language Modula/R on the personal computer Lilith. Software—Practice and Experience, 14(10):945–956, October 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reiser:1991:OSU

[Rei91] Martin Reiser. The Oberon System: User Guide and Programmer’s Manual. ACM Press, New York, NY 10036, USA, 1991. ISBN 0-201-54422-9. xii + 350 pp. LCCN QA76.76.O63 R45 1991. URL

                                        https://archive.org/stream/
                                        bitsavers_ethceresRe1991_
                                        13502706/Reiser_The_Oberon_
                                        System_1991_djvu.txt.

Ribar:1990:OPG

[Rib90] L. John Ribar. OS/2 programming goes Modula-2. Computer Language Magazine, 7(11): 83–??, November 1990. CODEN COMLEF. ISSN 0749-2839.

Richmond:1974:HP

[Ric74] George H. Richmond. History of Pascal. Pascal Newsletter, (2): 1–2, May 1974. URL http://

                                        pascal.hansotten.com/uploads/
                                        pug/02_Pascal_Newsletter_
                                        May74.pdf.

Rodionov:2001:OSM

[RL01] Alexey S. Rodionov and Dmitry V. Leskov. Oberon-2 as successor of Modula-2 in simulation. In Bjärner et al. [BBZ01], pages 538–545. ISBN 3-540-43075-X (softcover), 3-540-45575-2 (ebook). LCCN QA75.5 .I52 2001.

                                        URL http://link.springer.
                                        com/chapter/10.1007/3-540-
                                        45575-2_52.

Rustan:2003:AD

[RLN03] K. Rustan, M. Leino, and Greg Nelson. Abstraction dependencies. In Programming Methodology, pages 269–289. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 0-387-21798-3. ISSN 0172-603X.

URL http://link.springer.
com/chapter/10.1007/978-0-
387-21798-7_13.

Robson:1982:BRBb

[Rob82] D. J. Robson. Book review:

Pascal implementation: The P4 compiler, S. Pemberton and M. C. Daniels. Ellis Horwood, Chichester. No. of pages: 254. Price: £22.50. Software— Practice and Experience, 12(12): 1167, December 1982. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Robson:1983:TCL

[Rob83] D. J. Robson. Towards a conversational language-sensitive system for Pascal. Software— Practice and Experience, 13(11): 1013–1017, November 1983. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ross:1974:BRB

[Ros74] D. T. Ross. Book review: Systematic programming: An introduction, N. Wirth, Prentice-Hall, Englewood Cliffs, N.J., 1973. No. of pages: 167. Price: £5.25. Software—Practice and Experience, 4(1):112, January/March 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rosenwald:2024:NWV

[Ros24] Michael S. Rosenwald. Niklaus Wirth, visionary software architect, dies at 89. New York Times Web site, February 22, 2024.

                                        URL https://www.nytimes.
                                        com/2024/02/22/technology/
                                        niklaus-wirth-dead.html.

Rovner:1986:EMB

[Rov86]  Paul Rovner.  Extending

Modula-2 to build large, integrated systems. IEEE Software, 3(6):46–57, November 1986. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Russell:1976:IPC

[RS76] David L. Russell and Jeffrey V. Sue. Implementation of a Pascal compiler for the IBM 360. Software—Practice and Experience, 6(3):371–376, July/September 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rueping:1995:M

[RS95] A. Rueping and E. Sekerinski. Modula-3. Lecture Notes in Computer Science, 891:357– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reiser:1992:POS

[RW92] Martin Reiser and Niklaus Wirth. Programming in Oberon: Steps beyond Pascal and Modula. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-201-56543-9 (paperback). xvi + 320 pp. LCCN QA76.76.O63 R46 1992.

Reiser:1994:PON

[RW94] Martin Reiser and Niklaus Wirth. Programmieren in Oberon — das neue Pascal. (German) [Programming in Oberon — the new Pascal]. Addison-Wesley, Reading, MA, USA, 1994. ISBN 3-89319-657-9. xxi + 343 pp.

Rand:1987:PPIa

[RWZ87] D. W. Rand, P. Winternitz, and H. Zassenhaus. PASCAL programs for the identification of Lie algebras. II. SPLIT—a program to decompose parameterfree and parameter-dependent Lie algebras into direct sums. Computer Physics Communications, 46(2):297–309, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Sale:1979:SSA

[Sal79a] A. H. J. Sale. Strings and the sequence abstraction in Pascal. Software—Practice and Experience, 9(8):671–683, August 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sale:1979:ISP

[Sal79b]  Arthur Sale.  Implementing

strings in Pascal-Again. Software —Practice and Experience, 9 (10):839–841, October 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sale:1979:PSR

[Sal79c] Arthur Sale. Pascal stylistics and reserved words. Software— Practice and Experience, 9(10): 821–825, October 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sale:1981:ICS

[Sal81] Arthur Sale. The implementation of case statements in Pascal. Software—Practice and Experience, 11(9):929–942, September 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Salzman:1992:ASM

[Sal92] Eric Salzman. Activity simulation in Modula-2: an exercise in language extension. Computer Languages, 17(1):39–60, ???? 1992. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Savit:1987:UMA

[Sav87] J. Savit. Uninitialized Modula-2 abstract objects, revisited. ACM SIGPLAN Notices, 22(2): 78–84, February 1987. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sawitzki:1996:ESS

[Saw96] Günther Sawitzki. Extensible statistical software: On a voyage to Oberon. Journal of Computational and Graphical Statistics, 5 (3):263–283, 1996. CODEN ????

ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:

//www.tandfonline.com/doi/
abs/10.1080/10618600.1996.
10474711.

Schach:1980:PTP

[Sch80] Stephen R. Schach. A portable trace for the Pascal heap. Software—Practice and Experience, 10(6):421–426, June 1980. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schiper:1986:PCI

[Sch86] A. Schiper. Programmation concurrente: Illustré par des exemples en Portal, Modula-2 et Ada. Collection informatique. Pr. Polytechniques Romandes, Lausanne, Switzerland, 1986. ISBN 2-88074-093-2. 295 pp. LCCN ???? URL http:/

/www.zentralblatt-math.org/
zmath/en/search/?an=0659.
68032.

Schaller:1988:ME

[Sch88] Nan C. Schaller. The Modula-2 experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 20(3):42–44, September 1988. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schneider:1989:CPP

[Sch89a] Victor Schneider. Converting a portable Pascal p-code interpreter to a code generator. Software—Practice and Experience, 19(11):1111–1113, November 1989. CODEN SPEXBL.

ISSN 0038-0644 (print), 1097-024X (electronic).

Schulman:1989:MOJ

[Sch89b] A. Schulman. Modula-2 and OS/2 join forces. Byte Magazine, 14(8):171–174, August 1989. CODEN BYTEDJ. ISSN 0360-5280.

Schweller:1989:SCP

[Sch89c] Kenneth G. Schweller. Scheduling concurrent procedures using Modula-2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 21(2):51–56, June 1989. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schramm:2000:SUA

[Sch00a] Andreas Schramm. The “Structured-Universe approach” for irregular and dynamic spatial structures. In High Performance Computing Systems and Applications, pages 197–214. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 0-7923-7774-5. URL http://

                                        link.springer.com/chapter/
                                        10.1007/0-306-47015-2_24.

Schulthess:2000:LSI

[Sch00b] Peter Schulthess. Lean systems in an intrinsically complex world. In Böszörményi [Bös00], pages 87–94. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1 (dpunktVerl.). LCCN QA76.756 .S36 2000. URL http:

                                        //digitale-objekte.hbz-nrw.

de/storage2/2016/09/08/file_
8/6866144.pdf.

Schramm:2001:IDA

[Sch01] Andreas Schramm. An index domain for adaptive multi-grid methods. In Vector and Parallel Processing — VECPAR 2000, pages 330–343. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-44942-6. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/3-540-44942-6_27.

Syslo:2006:DOA

[SDK06] Maciej M. Syslo, Narsingh Deo, and Janusz S. Kowalik. Discrete Optimization Algorithms with Pascal Programs. Dover, New York, NY, USA, 2006. ISBN 0-486-45353-7 (paperback). xiv + 542 pp. LCCN QA402.5 .S94 2006. Reprint of the 1983 original.

Sedgewick:1993:AM

[Sed93] Robert Sedgewick. Algorithms in Modula-3. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-53351-0. xiv + 656 pp. LCCN QA76.73.M63 S43 1993.

Severance:2012:ATC

[Sev12] Charles Severance. The art of teaching computer science: Niklaus Wirth. Computer, 45(7): 8–10, 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//youtu.be/BJIqHIYSDrk;
http://youtu.be/Wel2b93LjhY.

Sewry:1984:MMC

[Sew84a] D. A. Sewry. Modula-2 and the monitor concept. ACM SIGPLAN Notices, 19(11):33– 41, November 1984. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sewry:1984:MPF

[Sew84b] D. A. Sewry. Modula-2 process facilities. ACM SIGPLAN Notices, 19(11):23–32, November 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sewell:1987:HTM

[Sew87] E. Wayne Sewell. How to MANGLE your software: the WEB system for Modula-2. TUGboat, 8(2):118–122, July 1987. ISSN 0896-3207. URL https:

                                        //tug.org/TUGboat/tb08-2/
                                        tb18sewellmangle.pdf.

Sewry:1988:PSM

[Sew88] D. A. Sewry. Process scheduling in Modula-2. ACM SIGPLAN Notices, 23(12):95–97, December 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shimasaki:1980:APP

[SFIK80] Masaaki Shimasaki, Shigeru Fukaya, Katsuo Ikeda, and Takeshi Kiyono. An analysis of Pascal programs in compiler writing. Software—Practice and Experience, 10(2):149– 157, February 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sumner:1982:MSP

[SG82] Roger T. Sumner and R. E.

Gleaves. Modula-2: A solution to Pascal’s problems. ACM SIGPLAN Notices, 17(9):28–33, September 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schiper:1989:CPI

[SH89] André Schiper and J. (Jack)

Howlett. Concurrent programming: illustrated with examples in Portal, Modula-2 and Ada. North Oxford Academic, London, UK, 1989. ISBN 0-946536-39-2. 234 pp. LCCN QA76.6. Translation to English by Jack Howlett of [Sch86].

Shammas:1986:MCI

[Sha86a] Namir Clement Shammas. Modula-2 compilers for the IBM PC. Dr. Dobb’s Journal of Software Tools, 11(10):48–??, October 1986. CODEN DDJOEB. ISSN 1044-789X.

Shammas:1986:TPP

[Sha86b] Namir Clement Shammas. Turbo Pascal procedural parameters, local modules in Modula-2. Dr. Dobb’s Journal of Software Tools, 11(12):96–??, December 1986. CODEN DDJOEB. ISSN 0888-3076.

Shammas:1987:TBC

[Sha87] Namir Clement Shammas. True Basic challenges Modula-2. Dr. Dobb’s Journal of Software Tools, 12(5):124–126, 128–130, May 1987. CODEN DDJOEB. ISSN 0888-3076.

Shiriaev:1992:PXP

[Shi92] Dmitri Shiriaev. PASCAL-XSC. A portable programming system for scientific computations. Interval Computations = Interval’nye vychisleniia, (4):76–82, 1992. ISSN 0135-4868. Interval ’92 (Moscow, 1992).

Shrivastava:1978:SPR

[Shr78] S. K. Shrivastava. Sequential Pascal with recovery blocks. Software—Practice and Experience, 8(2):177–185, March/April 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shrivastava:1979:CPBb

[Shr79a] S. K. Shrivastava. Concurrent Pascal with backward error recovery: Implementation. Software—Practice and Experience, 9(12):1021–1033, December 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shrivastava:1979:CPBa

[Shr79b] S. K. Shrivastava. Concurrent Pascal with backward error recovery: Language features and examples. Software—Practice and Experience, 9(12):1001– 1020, December 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Savic:2016:MVJ

[SIRB16] Miloš Savić, Mirjana Ivanović, Miloš Radovanović, and Zoran Budimac. Modula-2 versus Java as the first programming language: Evaluation of students’ performance. In Proceedings of the 17th International Conference on Computer Systems and Technologies 2016, CompSys-Tech ’16, pages 415–422. ACM Press, New York, NY 10036, USA, 2016. ISBN 1-4503-4182-9. LCCN ????

Sites:1972:AWR

[Sit72] Richard L. Sites. Algol W reference manual. Technical Report STAN-CS-71-230, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1972. i + 141 pp. URL https://

bitsavers.org/pdf/stanford/
cs_techReports/STAN-CS-71-
230_Algol_W_Reference_Manual_
Feb72.pdf.

Steensgaard-Madsen:1979:PCR

[SM79]  Jørgen  Steensgaard-Madsen.

Pascal—clarifications and recommended extensions. Acta Informatica, 12(1):73–94, June 25, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Steensgaard-Madsen:1981:MPP

[SM81]  Jørgen  Steensgaard-Madsen.

Modular programming with Pascal. Software—Practice and Experience, 11(12):1331–1337, December 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Szarapka:1995:TMP

[SM95] Lehez Szarapka and Dragan Mašulović. On translating Modula-2 programs to C: Local procedures and modules. Filomat, 9(2):295–302, Octo-

                                        ber 1995. URL https://www.
                                        jstor.org/stable/43997738.

Smedema:1983:PLP

[SMB83] C. H. Smedema, P. (Piet) Medema, and M. (Maarten) Boasson. The Programming Languages: Pascal, Modula, CHILL, and Ada. Prentice-Hall International, Upper Saddle River, NJ 07458, USA, 1983. ISBN 0-13-729756-4 (paperback). 154 pp. LCCN QA76.7 .S6 1983.

STUG:1984:UAS

[Sof84] Software Tools Users Group, editor. USENIX Association [and] Software Tools Users Group Summer Conference, Salt Lake City 1984: proceedings, June 12– 15, 1984, Salt Lake City, Utah, USA. USENIX, P.O. Box 7, El Cerrito 94530, CA, USA, 1984. LCCN QA76.8.U65 U83 1984.

Szyperski:1994:EPL

[SOM94] Clemens Szyperski, Stephen Omohundro, and Stephan Murer. Engineering a programming language: The type and class system of Sather. In Programming Languages and System Architectures, pages 208–227. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-48356-X. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.
com/chapter/10.1007/3-540-
57840-4_33.

Sommerer:2006:OSL

[Som06]  Ralph Sommerer.  Oberon

Script: a lightweight compiler and runtime system for the Web. In Modular Programming Languages, pages 73–83. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-540-40928-9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/
10.1007/11860990_6.

Souter:1990:PMA

[Sou90] John Souter. The position of MODULA-2 among programming languages. Microprocessors and Microsystems, 14(3): 145–148, April 1990. CO-DEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0141933190900643.

Schonhacker:1996:IIS

[SP96] M. Schönhacker and C. Pronk.

ISO/IEC 10514-1, the Standard for Modula-2: Changes, clarifications and additions. ACM SIG-PLAN Notices, 31(8):84–95, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spector:1982:AIM

[Spe82] David Spector. Ambiguities and insecurities in Modula-2. ACM SIGPLAN Notices, 17(8):43–51, August 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spector:1983:LPM

[Spe83] David Spector. Lexing and parsing Modula-2. ACM SIGPLAN Notices, 18(10):25–32, October 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spicer:2024:MNW

[Spi24]  Dag Spicer.  In memoriam:

Niklaus Wirth (1934–2024). Computer History Museum Web site, January 5, 2024. URL

                                        https://computerhistory.org/
                                        blog/in-memoriam-niklaus-
                                        wirth-1934-2024/.

Sloane:2015:OK

[SR15] Anthony M. Sloane and Matthew Roberts. Oberon-0 in Kiama. Science of Computer Programming, 114(??):20–32, December 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.sciencedirect.com/science/
                                        article/pii/S0167642315003032.

Segre:1985:SCA

[SS85] Lidia Segre and Michael Stanton. “Some Concerns about Modula-2” considered unwarranted. ACM SIGPLAN Notices, 20(5):31–35, May 1985. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [Mof84].

Sebesta:1986:FIA

[ST86] Robert W. Sebesta and Mark A.

Taylor. Fast identification of Ada and Modula-2 reserved words. Journal of Pascal, Ada and Modula-2, 5(2):36–39, March/ April 1986. CODEN JPAME8, JOPAD5. ISSN 0747-1351, 0735-1232.

Steindl:1998:ISO

[Ste98] Christoph Steindl. Intermodular slicing of object-oriented programs. In Compiler Construction, pages 264–278. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-69724-1. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/
chapter/10.1007/BFb0026437.

Sincovec:1986:DSU

[SW86] R. F. Sincovec and R. S. Wiener.

Data Structures Using Modula-2. John Wiley, New York, NY, USA, 1986. ISBN 0-471-81489-X. xxiii + 500 pp. LCCN QA76.73.M63 S56 1986.

Stubbs:1987:DSA

[SW87] Daniel F. Stubbs and Neil W. Webre. Data Structures with Abstract Data Types and Modula-2. Brooks/Cole, Pacific Grove, CA, USA, 1987. ISBN 0-534-07302-6. xx + 556 pp. LCCN QA76.9.D35 S76 1987. US$43.25. Textbook treatment of hashing.

Sydow:1994:I

[Syd94]  Friedrich V. Sydow.  O.

In InformLex, pages 231–244. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-322-83682-7. URL http:/

/link.springer.com/chapter/
10.1007/978-3-322-83682-3_

15. Glossary with Oberon and Oberon-2 entries.

Szyperski:1990:NCO

[Szy90] Clemens A. Szyperski. Network communication in the Oberon environment. Technical report 126, ETH Zürich, Zürich, Switzerland, February 1990. i + 43 pp. URL

http://hdl.handle.net/20.
500.11850/68708.

Szyperski:1991:WET

[Szy91] Clemens A. Szyperski. Write: an extensible text editor for the Oberon system. Technical report 151, ETH Zürich, Zürich, Switzerland, January 1991. i + 48 pp. URL

http://hdl.handle.net/20.
500.11850/68907.

Szyperski:1992:IIW

[Szy92a] Clemens A. Szyperski. Import is not inheritance why we need both: Modules and classes. In ECOOP ’92 European Conference on Object-Oriented Programming, pages 19–32. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-540-55668-0. URL http:/

/link.springer.com/chapter/
10.1007/BFb0053028.

Szyperski:1992:WIA

[Szy92b] Clemens A. Szyperski. Writeing applications: Designing an extensible text editor as an application framework. In Georg Heeg, Boris Magnusson, and Bertrand Meyer, editors, Technology of object-oriented languages and systems, TOOLS 7: proceedings of the seventh international conference, TOOLS Europe ’92, Dortmund, Germany, pages 247–261. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-917436-2. LCCN QA76.64 .T67 1992. URL https://dl.

acm.org/doi/10.5555/161213.
161244.

Szyperski:2000:MCR

[Szy00] Clemens Szyperski. Modules and components: rivals or partners? In Böszörményi [Bös00], pages 121–132. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1 (dpunktVerl.). LCCN QA76.756 .S36 2000. URL http:

//digitale-objekte.hbz-nrw.

                                        de/storage2/2016/09/08/file_
                                        8/6866144.pdf.

Tagg:1988:LLP

[Tag88] A. G. Tagg. Low-level programming in Modula-2: an experience with local area network access. Software—Practice and Experience, 18(6):495–507, June 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Taylor:1990:MMP

[Tay90]  Dave Taylor.  Metrowerks

Modula-2 PSE. Computer Language Magazine, 7(11):91–??, November 1990. CODEN COM-LEF. ISSN 0749-2839.

Templ:1990:SOU

[Tem90] Josef Templ. SPARC-Oberon: User’s guide and implementation. Technical report 133, ETH Zürich, Zürich, Switzerland, June 1990. i + 30 pp.

                                        URL http://hdl.handle.net/
                                        20.500.11850/68461.

Templ:1994:OPL

[Tem94] Josef Templ. The Oberon programming language. Dr. Dobb’s Journal of Software Tools, 19 (??):60–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Templ:1995:OOC

[Tem95] Josef Templ. Ofront^{T M}: Oberon-2 to C Translator, Version 1.0: User Guide. Software Templ, Linz, Austria, 1995. 101 pp. URL

                                        https://github.com/Oleg-N-
                                        Cher/OfrontPlus/.

Templ:2000:FOR

[Tem00] Josef Templ. FFF97: Oberon in the real world. In Böszörményi [Bös00], pages 227–244. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1 (dpunktVerl.). LCCN QA76.756 .S36 2000. URL http:/

/digitale-objekte.hbz-nrw.
de/storage2/2016/09/08/file_
8/6866144.pdf.

Tennent:1978:ALT

[Ten78] Robert D. Tennent. Another look at type compatibility in Pascal. Software—Practice and Experience, 8(4):429–437, July/ August 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tennent:1985:CAI

[Ten85] Robert D. Tennent. A comparison of the ANSI and ISO Pascal Standards. Software—Practice and Experience, 15(8):821–822, August 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Terry:1986:MKS

[Ter86] P. D. Terry. A Modula-2 kernel for supporting monitors. Software—Practice and Experience, 16(5):457–472, May 1986. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tessler:1985:OPR

[Tes85] Larry Tessler. Object Pascal report. Structured Language World, 9(3):10–14, ???? 1985. CODEN ???? ISSN 0176-9375.

Thomas:1992:POOa

[Tho92a] Richard Thomas. A proposal for object oriented Modula-2. In Addendum to the Proceedings on Object-Oriented Programming Systems, Languages, and Applications (Addendum), OOPSLA ’92, pages 171–173. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-610-7.

Thomas:1992:POOb

[Tho92b] Richard Thomas. A proposal for object oriented Modula-2. SIG-PLAN OOPS Mess., 4(2):171– 173, December 1992. ISSN 1055-6400 (print), 1558-0253 (electronic).

Torbett:1987:MAI

[Tor87a] M. A. Torbett. More ambiguities and insecurities in Modula-2. ACM SIGPLAN Notices, 22(5):11–17, May 1987. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Torbett:1987:NPA

[Tor87b] M. A. Torbett. A note on “Protecting against Uninitialized Abstract Objects in Modula-2”. ACM SIGPLAN Notices, 22(5):8–10, May 1987. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [Wie86].

Tsin:1982:EPP

[Tsi82] Yung H. Tsin. Extending the power of PASCAL’s external procedure mechanism. Software —Practice and Experience, 12 (3):283–292, March 1982. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Timmerer:2011:INEa

[TWB11a] Christian Timmerer, Niklaus Wirth, and Laszlo Böszörmenyi. An interview with Niklaus Emil Wirth, part 1. IEEE TV Web site, June 7, 2011. URL

https://ieeetv.ieee.org/an-
interview-with-niklaus-emil-
wirth-part-1.

Timmerer:2011:INEb

[TWB11b] Christian Timmerer, Niklaus Wirth, and Laszlo Böszörmenyi. An interview with Niklaus Emil Wirth, part 2. IEEE TV Web site, June 7, 2011. URL

https://ieeetv.ieee.org/an-
interview-with-niklaus-emil-
wirth-part-2.

Timmerer:2011:INEc

[TWB11c] Christian Timmerer, Niklaus Wirth, and Laszlo Böszörmenyi. An interview with Niklaus Emil Wirth, part 3. IEEE TV Web site, June 7, 2011. URL

https://ieeetv.ieee.org/an-
interview-with-niklaus-emil-
wirth-part-3.

USENIX:1985:PUA

[USE85] USENIX Association, editor.

Proceedings: USENIX Association Winter Conference, January 23–25, 1985, Dallas, Texas, USA. USENIX, P.O. Box 7, El Cerrito 94530, CA, USA, 1985.

vanDelft:1989:CO

[vD89] A. J. E. van Delft. Comments on Oberon. ACM SIGPLAN Notices, 24(3):23–30, March 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vixie:1992:M

[Vix92] Paul Vixie. Modula-3. ;login: the USENIX Association newsletter, 17(6):30–34, November/December 1992. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://archive.org/
                                        details/login_nov92/page/
                                        n29/mode/2up?view=theater.

Vaughan-Nichols:2024:TNW

[VN24] Steven J. Vaughan-Nichols. Tribute: Niklaus Wirth, 1934–2024: Remembering the creator of Pascal, a champion of lean and elegant code, and a vital figure in the evolution of software development. The Newstack Web site, January 12, 2024. URL https:

                                        //thenewstack.io/tribute-
                                        niklaus-wirth-1934-2024/.

Viera:2015:CCC

[VS15] Marcos Viera and S. Doaitse Swierstra. Compositional compiler construction: Oberon0. Science of Computer Programming, 114(??):45–56, December 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0167642315003019.

Walker:1986:DAM

[WA86] Bill Walker and Stephen Alexander. Data abstraction with Modula-2. Dr. Dobb’s Journal of Software Tools, 11(2):62–64, 94, 95, February 1986. CODEN DDJOEB. ISSN 1044-789X.

Wagner:2000:LMW

[Wag00] Bernhard Wagner. Lilith meets the world of business. In Böszörményi [Bös00], pages 203– 218. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1 (dpunktVerl.). LCCN QA76.756 .S36 2000. URL http:/

/digitale-objekte.hbz-nrw.
de/storage2/2016/09/08/file_
8/6866144.pdf.

Walker:1986:IPP

[Wal86] A. N. Walker. The Ilem Project:

Pascal to Algol in easy stages I: Outward bound. Software— Practice and Experience, 16(12): 1097–1107, December 1986. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wallis:1989:BRB

[Wal89] Peter Wallis. Book review:

Algorithms and data structures, by N. Wirth. Prentice-Hall International, Hemel Hempstead, United Kingdom, 1986, Price £15.95/$27.95 (paperback), ISBN 0-13-021999-1. Science of Computer Programming, 12(2):157–158, July 1989. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0167642389900403.

Ward:2003:VWH

[War03] Brian Ward. VMware Workstation: [das Handbuch; Installation, Konfiguration, Anwendung und Troubleshooting; Gast-Systeme: Windows, Linux, BSD, Novell NetWare, Solaris, FreeDOS und Oberon; virtuelle Netzwerke, Netzwerkkonfiguration und -dienste]. mitp-Verlag, Landsberg, Germany, 2003. ISBN 3-8266-0964-6. 336 pp. LCCN ???? EUR 32.00 (DE). URL http://

www.gbv.de/du/services/agi/
FCC0A57071BE8695C125704A0029797F/

FLMA122525. Translation of American original to German by Gerhard Franken.

Welsh:1979:PPA

[WB79] J. Welsh and D. W. Bustard. Pascal-Plus — another language for modular multiprogramming. Software—Practice and Experience, 9(11):947–957, November 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

White:1985:PPR

[WB85a] N. H. White and K. H. Bennett. PRTDS — A Pascal run-time diagnostics system. Software— Practice and Experience, 15(11): 1041–1056, November 1985. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

White:1985:RTD

[WB85b] N. H. White and K. H. Bennett.

Run-time diagnostics in Pascal. Software—Practice and Experience, 15(4):359–367, April 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Waite:1981:ASI

[WC81] W. M. Waite and Lynn Robert Carter. An analysis/synthesis interface for Pascal compilers. Software—Practice and Experience, 11(8):769–787, August 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wichmann:1983:PCV

[WC83] Brian Anderson Wichmann and Z. J. Ciechanowicz, editors. Pascal compiler validation. Wiley, New York, NY, USA, 1983. ISBN 0-471-90133-4. xiii + 176 pp. LCCN QA76.73.P2 P35 1983.

Weber:1967:MIE

[Web67] Helmut Weber. A microprogrammed implementation of EU-LER on IBM System/360 model 30. Communications of the ACM, 10(9):549–558, September 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weck:1998:IUC

[Wec98] Wolfgang Weck. Inheritance using contracts & object composition. In Object-Oriented Technologys, pages 384–388. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-69687-3. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.
com/chapter/10.1007/3-540-
69687-3_78.

Welsh:1978:ERC

[Wel78]  J. Welsh.  Economic range

checks in Pascal. Software— Practice and Experience, 8(1): 85–97, January/February 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wittie:1984:PMO

[WF84] Larry D. Wittie and Ariel J. Frank. A portable Modula-2 operating system: SAM2S. In Proceedings of the July 9-12, 1984, National Computer Conference and Exposition, AFIPS ’84, pages 283–292. ACM Press, New York, NY 10036, USA, 1984. ISBN 0-88283-043-0.

Wirth:1989:OS

[WG89] Niklaus Wirth and Jürg Gutknecht. The Oberon system. Software— Practice and Experience, 19(9): 857–893, September 1989. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wirth:1992:POD

[WG92] Niklaus Wirth and Jürg Gutknecht. Project Oberon: The Design of an Operating System and Compiler. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-54428-8. xi + 548 pp. LCCN QA76.76.O63 W58 1992.

Wirth:1988:OS

[WG8 ] Niklaus Wirth and Jürg Gutknecht.

The Oberon system. Report 88, ETH Zürich, Zürich, Switzerland, 1988 (??). 32 + 13 figures pp.

Wirth:1966:CDA

[WH66a] Niklaus Wirth and C. A. R.

Hoare. A contribution to the development of ALGOL. Communications of the ACM, 9(6): 413–432, June 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See errata [WH66b]. This is the defining document for the Algol W language; see [BBG^{+}69].

Wirth:1966:ECD

[WH66b] Niklaus Wirth and C. A. R.

Hoare. Errata: “A Contribution to the Development of ALGOL”. Communications of the ACM, 9 (12):878, December 1966. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [WH66a].

Welsh:1986:MIS

[WH86] Jim Welsh and Atholl Hay. A Model Implementation of Standard Pascal. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1986. ISBN 0-13-586454-2. xi + 483 pp. LCCN QA76.73P2 W47.

Wichmann:1984:CPA

[Wic84] B. A. Wichmann. A comparison of Pascal and Ada. In Feuer and Gehani [FG84a], page ?? ISBN 0-13-154840-9 (paperback), 0-13-154857-3 (hardcover). LCCN QA76.73.A35 C66 1984.

Wiener:1984:GSM

[Wie84] Richard S. Wiener. A generic sorting module in Modula-2. ACM SIGPLAN Notices, 19 (3):66–72, March 1984. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wiener:1986:PAU

[Wie86] R. S. Wiener. Protecting against uninitialized abstract objects in Modula-2. ACM SIGPLAN Notices, 21(6):63–69, June 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See note [Tor87b].

Wilson:1980:PSH

[Wil80] I. R. Wilson. Pascal for school and hobby use. Software— Practice and Experience, 10(8): 659–671, August 1980. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Williams:1982:FNS

[Wil82] M. Howard Williams. A flexible notation for syntactic definitions. ACM Transactions on Programming Languages and Systems, 4 (1):113–119, January 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Wirth:1963:GAb

[Wir63a] Niklaus Wirth. A generalization of ALGOL. Communications of the ACM, 6(9):547–554, September 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wirth:1963:GAa

[Wir63b] Niklaus Emil Wirth. A Generalization of Algol. Ph.D. thesis, University of California, Berkeley, Berkeley, CA, 1963. 107 pp.

URL https://www.proquest.
com/pqdtglobal/docview/302115469.

Wirth:1964:PSM

[Wir64a] N. Wirth. A proposal on string manipulation within ALGOL 60. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (17): 13–17, July 1964. CODEN AL-GOBG. ISSN 0084-6198.

Wirth:1964:CAM

[Wir64b] Niklaus Wirth. Certification of Algorithm 146: Multiple integration. Communications of the ACM, 7(5):296, May 1964. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wirth:1965:CPV

[Wir65a] N. Wirth. Comments on a paper by A. van Wijngaarden. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (19):11–12, January 1965. CODEN ALGOBG. ISSN 0084-6198.

Wirth:1965:RIF

[Wir65b] N. Wirth. R65-24 introduction to FORMAC. IEEE Transactions on Electronic Computers, EC-14(2):278, April 1965. CODEN IEECA8. ISSN 0367-7508. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4038432.

Wirth:1965:AFP

[Wir65c] Niklaus Wirth. Algorithm 265: Find precedence functions [L2]. Communications of the ACM, 8 (10):604–605, October 1965. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wirth:1965:PLC

[Wir65d] Niklaus Wirth. A programming language for the 360 computers. Report CS-33, Department of Computer Science, Stanford University, Stanford, CA 94305, USA, December 24, 1965. ii + 25 pp. URL https://

bitsavers.org/pdf/stanford/
cs_techReports/CS33_Wirth_
PL360_Dec65.pdf.

Wirth:1965:AO

[Wir65e] Niklaus E. Wirth. Algorithm 249: Outreal N [15]. Communications of the ACM, 8(2): 104, February 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wirth:1966:ANC

[Wir66a] N. Wirth. Additional notes on — contribution to the development of ALGOL. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (24):13–17, September 1966. CODEN AL-GOBG. ISSN 0084-6198.

Wirth:1966:LEN

[Wir66b] Niklaus Wirth. Letter to the Editor: A note on “Program structures for parallel processing”. Communications of the ACM, 9(5):320–321, May 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wirth:1966:PLC

[Wir66c] Niklaus Wirth. A programming language for the 360 computers. Technical Report CS53, Stanford University, Department of Computer Science, Stanford, CA, USA, December 20, 1966. ii + 81

pp. URL http://i.stanford.
edu/pub/cstr/reports/cs/tr/
66/53/CS-TR-66-53.pdf.

Wirth:1967:IA

[Wir67a] Niklaus Wirth. Information for authors. IEEE Transactions on Electronic Computers, EC-14(3): c2, June 1967. CODEN IEECA8. ISSN 0367-7508.

Wirth:1967:CBC

[Wir67b] Niklaus Wirth. On certain basic concepts of programming languages. Technical Report CS-TR-67-65, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1, 1967. ii + 30 pp. URL http:

//i.stanford.edu/pub/cstr/
reports/cs/tr/67/65/CS-TR-
67-65.pdf.

Wirth:1967:PS

[Wir67c] Niklaus Wirth. The PL360 system. Report CS-TR-67-68, Department of Computer Science, Stanford University, Stanford, CA 94305, USA, June 5, 1967. ix + 63 pp.

URL http://i.stanford.edu/
pub/cstr/reports/cs/tr/67/
68/CS-TR-67-68.pdf;   http:
//infolab.stanford.edu/TR/
CS-TR-67-68.html.

Wirth:1967:PLC

[Wir67d] Niklaus Wirth. A programming language for the 360 computers. Report CS-TR-65, Department of Computer Science, Stanford University, Stanford, CA 94305, USA, 1967.

Wirth:1967:SCT

[Wir67e] Niklaus Wirth. Soviet cybernetics technology: IX, ALGEC — summary and critique. Memorandum RM-5157-PR, Rand Corporation, Santa Monica, CA, USA, 1967. viii + 43 pp.

Wirth:1968:ACC

[Wir68a] Niklaus Wirth. ALGOL colloquium – closing word. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), (29):16–19, May 31, 1968. CODEN ALGOBG. ISSN 0084-6198. URL http:

//archive.computerhistory.
org/resources/text/algol/
ACM_Algol_bulletin/1061166/
p16-wirth.pdf.

Wirth:1968:CPP

[Wir68b] Niklaus Wirth. Corrigendum: “PL360, a programming language for the 360 computers”. Journal of the ACM, 15(3):489, July 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Wir68c]. An appendix was duplicated, but with quite different typography.

Wirth:1968:PPL

[Wir68c] Niklaus Wirth. PL360, a programming language for the 360 computers. Journal of the ACM, 15(1):37–74, January 1968. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [Wir68b].

Wirth:1969:BCC

[Wir69a] Niklaus Wirth. A basic course on compiler principles. BIT (Nordisk tidskrift for informationsbehandling), 9(4):362–386, December 1969. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0006-
3835&volume=9&issue=4&spage=
362.

Wirth:1969:MMC

[Wir69b] Niklaus Wirth. On multiprogramming, machine coding, and computer organization. Communications of the ACM, 12(9): 489–498, September 1969. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See corrigendum [Wir70a].

Wirth:1970:CMM

[Wir70a] N. Wirth. Corrigenda: “On Multiprogramming, Machine Coding, and Computer Organization”. Communications of the ACM, 13(4):266, April 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Wir69b].

Wirth:1970:ADP

[Wir70b] Niklaus Wirth. An axiomatic definition of the programming language PASCAL. Report 6, ETH Zürich, Zürich, Switzerland, December 1970. ???? pp.

Wirth:1970:PLP

[Wir70c] Niklaus Wirth. The programming language PASCAL. Report 1, ETH Zürich, Zürich, Switzerland, November 1970. v

                                        + 58 pp. URL http://pascal.
                                        hansotten.com/uploads/wirth/
                                        The_Programming_Language_
                                        Pascal_1970%20001.pdf;   https:
                                        //www.research-collection.
                                        ethz.ch/bitstream/handle/
                                        20.500.11850/68712/eth-3290-
                                        01.pdf.

Wirth:1971:DPC

[Wir71a] Niklaus Wirth. The design of a PASCAL compiler. Software —Practice and Experience, 1 (4):309–333, October/December 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wirth:1971:PDSa

[Wir71b] Niklaus Wirth. Program development by stepwise refinement. Report 2, ETH Zürich, Zürich, Switzerland, January 1971. iii

                                        + 22 pp. URL http://pascal.
                                        hansotten.com/uploads/wirth/
                                        Program%20development%20by%

20step-wise%20refinement%
20jan%201971%20002.pdf.

Wirth:1971:PDSb

[Wir71c] Niklaus Wirth. Program development by stepwise refinement. Communications of the ACM, 14 (4):221–227, April 1971. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wirth:1971:PLP

[Wir71d] Niklaus Wirth. The programming language Pascal. Acta Informatica, 1(1):35–63, January 1971. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:/

/link.springer.com/article/
10.1007/BF00264291. See com-

ments [Hab73, LD75].

Wirth:1972:PCP

[Wir72a] N. Wirth. Planned changes to the programming language PAS-CAL. Report, ETH Zürich, Zürich, Switzerland, June 1972.

5 pp. URL http://pascal.
hansotten.com/uploads/wirth/
Planned_Changes_To_The_Programming_
Language_PASCAL_Jun72.pdf;
 http://www.bitsavers.org/
pdf/eth/pascal/Planned_Changes_
To_The_Programming_Language_
PASCAL_Jun72.pdf.

Wirth:1972:PLPc

[Wir72b] N. Wirth. The programming language Pascal and its design criteria. In High Level Languages: International Computer State of the Art Report, volume 7, pages 451–473. Infotech, Maidenhead, UK, 1972. ISBN 0-85539-060-3. LCCN QA76.7 .H52.

Wirth:1972:SPG

[Wir72c] N. Wirth. Systematisches Programmieren. (German) [Systematic Programming], volume 17 of Leitfaden der angewandten Mathematik und Mechanik. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1972. ISBN 3-519-02327-X. 160 pp. LCCN QA76.6 .W57.

Wirth:1972:GES

[Wir72d] Niklaus Wirth. Guest editorial: Software and its portability. Software—Practice and Experience, 2(4):311–312, October 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wirth:1972:PCG

[Wir72e] Niklaus Wirth. On “Pascal”, code generation, and the CDC 6000 computer. Report STAN-CS-72-257, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1972. i + 38 pp. URL http:

                                        //infolab.stanford.edu/pub/
                                        cstr/reports/cs/tr/72/257/
                                        CS-TR-72-257.pdf.

Wirth:1972:PLPa

[Wir72f]  Niklaus Wirth.  The pro-

gramming language PASCAL. Report, ETH Zürich, Zürich, Switzerland, August 1972. ???? pp.

Wirth:1972:PLPb

[Wir72g]  Niklaus Wirth.  The pro-

gramming language PASCAL (revised report). Report 5, ETH Zürich, Zürich, Switzerland, November 1972. viii +

48 pp. URL http://pascal.
hansotten.com/uploads/wirth/
The_Programming_Language_
Pascal_1972.pdf.

Wirth:1973:IPS

[Wir73a] Niklaus Wirth. Implementations of Pascal on systems with no control characters. Report, ETH Zürich, Zürich, Switzerland, June 26, 1973. 2 pp.

Wirth:1973:PLP

[Wir73b] Niklaus Wirth. The programming language PASCAL, (revised report). Report 5, ETH Zürich, Zürich, Switzerland, July 1973. iv + 49

pp.  URL http://pascal.
hansotten.com/uploads/wirth/
The_Programming_Language_
Pascal_1973.pdf;   https:/
/www.research-collection.
ethz.ch/handle/20.500.11850/
68910.

Wirth:1973:SPI

[Wir73c]  Niklaus Wirth.  Systematic

Programming: an Introduction. Prentice-Hall series in automatic computation. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1973. ISBN 0-13-880369-2. xiv + 169 pp. LCCN QA76.6 .W5713.

Wirth:1974:CWS

[Wir74a] Niklaus Wirth. On the composition of well-structured programs. ACM Computing Surveys, 6(4): 247–259, December 1974. CO-DEN CMSVAN. ISSN 0010-

4892. URL http://pascal.
hansotten.com/uploads/wirth/
StructuredPrograms.pdf.

Wirth:1974:DPL

[Wir74b] Niklaus Wirth. On the design of programming languages. In Information Processing 74 (Proc. IFIP Congress, Stockholm, 1974), pages 386–393. North-Holland, Amsterdam, The Netherlands, 1974. URL http:/

/pascal.hansotten.com/uploads/
wirth/On%20the%20Design%20of%
20Prog%20Languages%20Wirth.
PDF.

Wirth:1974:P

[Wir74c] Niklaus Wirth. Pascal 6000-3.4. Pascal Newsletter, (2):6–17, May

1974. URL http://pascal.
hansotten.com/uploads/pug/
02_Pascal_Newsletter_May74.
pdf.

Wirth:1974:SSU

[Wir74d] Niklaus Wirth. Systemprogramming aus der Sicht der Universität. (German) [System programming from the university’s perspective]. In Clemens Hackl, editor, Programming Methodology, 4th Informatik Symposium, IBM Germany, Wildbad, September 25–27, 1974, volume 23 of Lecture Notes in Computer Science, pages 192–202.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974.

Wirth:1975:APLc

[Wir75a] N. Wirth. An assessment of the programming language Pascal. IEEE Transactions on Software Engineering, SE-1(2):192–198, June 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312839.

Wirth:1975:GRW

[Wir75b] N. Wirth. A generalization of the read and write procedures. Pascal Newsletter, (3):11–12, February 1975.

Wirth:1975:PSI

[Wir75c] N. Wirth. PASCAL-S: a subset and its implementation. Report 12, ETH Zürich, Zürich, Switzerland, June 1975. 63 pp.

Wirth:1975:PPC

[Wir75d] N. Wirth. The PASCAL(P) compiler: Implementation notes. Report ??, ETH Zürich, Zürich, Switzerland, 1975.

Wirth:1975:AD

[Wir75e] Niklaus Wirth. Algorithmen und Datenstrukturen. (German) [Algorithms and Data Structures]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1975. ISBN 3-519-02330-X. 376 pp. LCCN QA76.9.D35 W57. DM26.80.

Wirth:1975:APLa

[Wir75f] Niklaus Wirth. An assessment of the programming language PAS-CAL. In Proceedings of the International Conference on Reliable Software, pages 23–30. ACM Press, New York, NY 10036, USA, 1975. ISBN 1-4503-7385-2.

Wirth:1975:APLb

[Wir75g] Niklaus Wirth. An assessment of the programming language PAS-CAL. ACM SIGPLAN Notices, 10(6):23–30, June 1975. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wirth:1975:SPE

[Wir75h] Niklaus Wirth. Systematisches Programmieren eine Einführung. (German) [Systematic programming an introduction], volume 17 of Leitfäden der angewandten Mathematik und Mechanik; Teubner-Studienbücher. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, second edition, 1975. ISBN 3-519-12327-4. 160 pp. LCCN QA76.6 .W57 1975.

Wirth:1976:CND

[Wir76a] N. Wirth. Comment on a note on dynamic arrays in PAS-CAL. ACM SIGPLAN Notices, 11(1):37–38, January 1976. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wirth:1976:CE

[Wir76b] N. Wirth. Corrections to the 2nd edition. Pascal Newsletter, (5):6, September 1976.

Wirth:1976:PLW

[Wir76c] N. Wirth. Programming languages: what to demand and how to assess them, and Professor Cleverbyte’s visit to Heaven. Report 17, ETH Zürich, Zürich, Switzerland, 1976. i + 32

pp.  URL http://pascal.
hansotten.com/uploads/wirth/
Programming_Languages_What_
To_Demand_And_How_To_Assess_
Them_Apr76.pdf.

Wirth:1976:ADS

[Wir76d] Niklaus Wirth. Algorithms + Data Structures = Programs. Prentice-Hall Series in Automatic Computation. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1976. ISBN 0-13-022418-9. xvii + 366 pp. LCCN QA76.6 .W561.

Wirth:1976:RBP

[Wir76e]  Niklaus Wirth.  Revidierter

Bericht über die Programmiersprache PASCAL. (German) [Revised report on the PAS-CAL programming language]. Akademie-Verlag, Berlin, Germany, 1976. ISBN 3-11-257911-9 (print), 3-11-257912-7. x + 48 pp. LCCN QA76.73.P2 W57.

Wirth:1977:GEG

[Wir77a] N. Wirth. Guest editorial: Guest editorial extract from Professor Cleverbyte’s visit to heaven.

Software—Practice and Experience, 7(2):155–158, March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wirth:1977:TDRa

[Wir77b] N. Wirth. Towards a discipline of real-time programming. ACM SIGSOFT Software Engineering Notes, 2(2):142, March 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wirth:1977:TDRb

[Wir77c] N. Wirth. Towards a discipline of real-time programming. Operating Systems Review, 11(2): 142, April 1977. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586x (electronic).

Wirth:1977:TDRc

[Wir77d] N. Wirth. Towards a discipline of real-time programming. In Proceedings of an ACM Conference on Language Design for Reliable Software, page 142. ACM Press, New York, NY 10036, USA, 1977. ISBN 1-4503-7380-1.

Wirth:1977:TDRd

[Wir77e] N. Wirth. Towards a discipline of real-time programming. ACM SIGPLAN Notices, 12(3):142, March 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wirth:1977:CE

[Wir77f]  Niklaus Wirth.  Compiler-

bau: Eine Einführung. (German) [Compiler Construction: an Introduction], volume 36 of Leitfäden der angewandten Mathematik und Mechanik LAMM, Band; Teubner Studienbücher: Informatik. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1977. ISBN 3-519-02338-5. 94 pp. LCCN QA76.6 .W563.

Wirth:1977:DIM

[Wir77g] Niklaus Wirth. Design and implementation of Modula. Software—Practice and Experience, 7(1):67–84, January/ March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wirth:1977:MLM

[Wir77h] Niklaus Wirth. Modula: a language for modular multiprogramming. Software—Practice and Experience, 7(1):3–35, January/March 1977. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wirth:1977:TDRe

[Wir77i] Niklaus Wirth. Toward a discipline of real-time programming. Communications of the ACM, 20 (8):577–583, August 1977. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wirth:1977:UM

[Wir77j] Niklaus Wirth. The use of Modula. Software—Practice and Experience, 7(1):37–65, January/ March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wirth:1977:WCW

[Wir77k] Niklaus Wirth. What can we do about the unnecessary diversity of notation for syntactic definitions? Communications of the ACM, 20(11): 822–823, November 1977. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://pascal.
hansotten.com/uploads/wirth/
What%20can%20we%20do%20Notation%
20EBNF.pdf.

Wirth:1978:M

[Wir78a] N. Wirth. MODULA-2. Report 27, ETH Zürich, Zürich, Switzerland, December 1978.

Wirth:1978:OPI

[Wir78b] N. Wirth. Obeisance to Pascal inventor. Electronics: the International Magazine of Electronics Technology, 51(26):6, December 21, 1978. ISSN 0883-4989. URL https://

www.worldradiohistory.com/
Archive-Electronics/70s/78/
Electronics-1978-12-21.pdf.

Wirth:1978:TCI

[Wir78c] Niklaus Wirth. Technical correspondence: Interlude on signals and semaphores revisited. author’s response. Communications of the ACM, 21(7):592, July 1978. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Wirth:1979:CPP

[Wir79a]  N. Wirth.  A collection of

Pascal programs. Report 33, ETH Zürich, Zürich, Switzerland, July 1979. iii + 75

pp.  URL http://pascal.
hansotten.com/uploads/wirth/
A%20Collection%20of%20Pascal%
20Programs%20Wirth%201979%
20033.pdf.

Wirth:1979:MSS

[Wir79b] Niklaus Wirth. The module: a system structuring facility in high-level programming languages. In Jeffrey M. Tobias, editor, Language Design and Programming Methodology, Proceedings of a Symposium Held in Sydney, Australia, 10– 11 September, 1979, volume 79 of Lecture Notes in Computer Science, pages 1–24. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979.

Wirth:1979:PCB

[Wir79c] Niklaus Wirth. A personal computer based on a highlevel language. In Jeffrey M. Tobias, editor, Language Design and Programming Methodology, Proceedings of a Symposium Held in Sydney, Australia, 10–11 September, 1979, volume 79 of Lecture Notes in Computer Science, pages 191–194. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979.

Wirth:1979:PCD

[Wir79d] Niklaus Wirth. A personal computer designed for use with a high-level language. In Werner Remmele and Heinz Schecher, editors, Microcomputing, Tagung III/1979 des German Chapter of the ACM am 24. und 25.10.1979 in München, volume 3 of Berichte des German Chapter of the ACM, pages 115– 134. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1979.

Wirth:1980:CRG

[Wir80a] N. Wirth. [Cross reference generator for Pascal programs]: Program Xref. Pascal Newsletter, (17):41–46, March 1980. Program dated 10 February 1976.

Wirth:1980:M

[Wir80b] N. Wirth. MODULA-2. Report 36, ETH Zürich, Zürich, Switzerland, 1980.

Wirth:1980:PP

[Wir80c] N. Wirth. Program PASCAL-S. Pascal Newsletter, (19):30–40, September 1980.

Wirth:1981:LPCa

[Wir81a] N. Wirth. Lilith: a personal computer for the software engineer. In Proceedings of the 5th International Conference on Software Engineering, ICSE ’81, pages 2–15. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1981. ISBN 0-89791-146-6.

Wirth:1981:LPCb

[Wir81b] Niklaus Wirth. LILITH: a personal computer for the software engineer. In Michael J. Flynn, Neville Harris, and Daniel P. McCarthy, editors, Microcomputer System Design, An Advanced Course, Trinity College Dublin, Ireland, June 1981, volume 126 of Lecture Notes in Computer Science, pages 349– 397. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981.

Wirth:1981:PSI

[Wir81c] Niklaus Wirth. Pascal-S: a subset and its implementation. In Barron [Bar78], pages 199–259. ISBN 0-471-27835-1 (hardcover). LCCN QA76.73.P2 P2.

Wirth:1982:PM

[Wir82] Niklaus Wirth. Programming in Modula-2. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-11674-5 (New York), 3-540-11674-5 (Berlin). 176 pp. LCCN QA76.73.M63 W5713 1982.

Wirth:1983:AD

[Wir83a] N. Wirth. Algorithmen und Datenstrukturen. (German) [Algorithms and Data Structures]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, third edition, 1983. ISBN 3-519-02250-8, 3-322-94704-1. 320 pp. LCCN ????

Wirth:1983:PDS

[Wir83b] Niklaus Wirth. Program development by stepwise refinement (reprint). Communications of the ACM, 26(1):70–74, January 1983. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Reprint of [?].

Wirth:1983:PM

[Wir83c] Niklaus Wirth. Programming in Modula-2. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1983. ISBN 0-387-12206-0 (New York), 3-540-12206-0 (Berlin), 3-540-15078-1, 3-642-96757-4 (e-book), 3-642-96759-0 (print), 3-642-96880-5 (print). 176 pp. LCCN QA76.73.M63 W5713 1983.

Wirth:1984:HGM

[Wir84a] N. Wirth. History and goals of Modula-2. Byte Magazine, 9 (??):145–152, August 1984. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Modula-2.

Wirth:1984:APL

[Wir84b] Niklaus Wirth. An assessment of the programming language Pascal. In Feuer and Gehani [FG84a], page ?? ISBN 0-13-154840-9 (paperback), 0-13-154857-3 (hardcover). LCCN QA76.73.A35 C66 1984.

Wirth:1984:DSAa

[Wir84c] Niklaus Wirth. Data structures and algorithms. Report 60, ETH Zürich, Zürich, Switzerland, 1984. 18 pp. URL http://

pascal.hansotten.com/uploads/
wirth/Data%20Structures%20Algorithms%
20Informatics%20060.pdf.

Wirth:1984:DSAb

[Wir84d] Niklaus Wirth. Data structures and algorithms. Scientific American, 251(3):60–69 (Intl. ed. 48–57), September 1984. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/
scientificamerican/journal/
v251/n3/pdf/scientificamerican0984-
60.pdf.

Wirth:1984:PLW

[Wir84e] Niklaus Wirth. Programming languages: what to demand and how to assess them. In Feuer and Gehani [FG84a], page ?? ISBN 0-13-154840-9 (paperback), 0-13-154857-3 (hardcover). LCCN QA76.73.A35 C66 1984.

Wirth:1985:PLD

[Wir85a] Niklaus Wirth. From programming language design to computer construction. Communications of the ACM, 28(2): 160–164, February 1985. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.
acm.org/pubs/toc/Abstracts/
0001-0782/2789.html.

Wirth:1985:PMG

[Wir85b] Niklaus Wirth. Programmieren in Modula-2. (German) [Programming in Modula-2]. Texts and Monographs in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third corrected edition, 1985. ISBN 0-387-13301-1 (New York), 3-540-13301-1 (Berlin). xiv + 220 pp. LCCN ????

Wirth:1985:PM

[Wir85c] Niklaus Wirth. Programming in Modula-2. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third corrected edition, 1985. ISBN 0-387-15078-1, 3-540-15078-1 (West Germany). 176 pp. LCCN QA76.73.M63 W5713 1983.

Wirth:1986:ADM

[Wir86a] Niklaus Wirth. Algorithmen und Datenstrukturen mit Modula-2. (German) [Algorithms and Data Structures with Modula-2]. Leitfäden und Monographien der Informatik. [Guides and Monographs in Information Science]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, fourth edition, 1986. ISBN 3-519-02260-5. 299 pp.

Wirth:1986:ADS

[Wir86b] Niklaus Wirth. Algorithms and Data Structures. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1986. ISBN 0-13-021999-1, 0-13-022005-1. 288 pp.

LCCN QA76.9.D35 W58 1986. US$32.95.

Wirth:1986:MAC

[Wir86c] Niklaus Wirth. Microprocessor architectures: a comparison based on code generation by compiler. Communications of the ACM, 29(10):978–990, October 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/
0001-0782/6619.html.

Wirth:1987:ERT

[Wir87a] Niklaus Wirth. Extension of record types. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 19(2):2–9, June 1987. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wirth:1987:PLD

[Wir87b] Niklaus Wirth. From programming language design to computer construction. In Ashenhurst [Ash87], page ?? ISBN 0-201-07794-9, 0-89791-183-0. LCCN QA76.24 .A33 1987. ACM Turing Award lecture, reprinted from [Wir85a].

Wirth:1987:HAPa

[Wir87c] Niklaus Wirth. Hardware architectures for programming languages and programming languages for hardware architectures. ACM SIGARCH Computer Architecture News, 15(5): 2–8, October 1987. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Wirth:1987:HAPb

[Wir87d] Niklaus Wirth. Hardware architectures for programming languages and programming languages for hardware architectures. Operating Systems Review, 21(4):2–8, October 1987. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wirth:1987:HAPc

[Wir87e] Niklaus Wirth. Hardware architectures for programming languages and programming languages for hardware architectures. ACM SIGPLAN Notices, 22(10):2–8, October 1987. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wirth:1987:HAPd

[Wir87f] Niklaus Wirth. Hardware architectures for programming languages and programming languages for hardware architectures. In Proceedings of the Second International Conference on Architectural Support for Programming Languages and Operating Systems, ASPLOS II, pages 2–8. ACM Press, New York, NY 10036, USA, 1987. ISBN 0-8186-0805-6.

Wirth:1988:OSW

[Wir88a] N. Wirth. Oberon: a system for workstations. Microprocessing and Microprogramming, 24 (1–5):3–8, August 1988. CO-DEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (electronic).

Wirth:1988:MO

[Wir88b] Niklaus Wirth. From Modula to Oberon. Software—Practice and Experience, 18(7):661–670, July 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wirth:1988:PM

[Wir88c] Niklaus Wirth. Programming in Modula-2. Texts and Monographs in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 1988. ISBN 0-387-50150-9, 3-540-50150-9. 182 pp. LCCN QA76.73.M63 W5713

1988. URL http://freepages.
modula2.org/report4/modula-
2.html.

Wirth:1988:PLO

[Wir88d] Niklaus Wirth. The programming language Oberon. Software—Practice and Experience, 18(7):671–690, July 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wirth:1988:TE

[Wir88e] Niklaus Wirth. Type extensions.

ACM Transactions on Programming Languages and Systems, 10 (2):204–214, April 1988. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL http://pascal.
hansotten.com/uploads/wirth/
Type%20Extensions%20nw1988.
pdf;   http://www.acm.org/
pubs/toc/Abstracts/0164-0925/

46167.html. See remarks

[Coh91, Wir91c].

Wirth:1989:DSS

[Wir89] Niklaus Wirth. Designing a system from scratch. Structured Programming, 10(1):10–18, 1989. CODEN STPGEM. ISSN 0935-1183.

Wirth:1990:MOO

[Wir90a] N. Wirth. MODULA-2 and object-oriented programming. Microprocessors and Microsystems, 14(3):149–152, April 1990. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

                                        tronic). URL https://www.
                                        sciencedirect.com/science/
                                        article/abs/pii/0141933190900654.

Wirth:1990:CNL

[Wir90b] Niklaus Wirth. Ceres-Net: a low-cost computer network. Software—Practice and Experience, 20(1):13–24, January 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wirth:1990:POG

[Wir90c]  Niklaus Wirth.  Die Pro-

grammiersprache Oberon. (German) [The Oberon programming language]. Technical report 140, ETH Zürich, Zürich, Switzerland, October 1990.

                                        URL http://hdl.handle.net/
                                        20.500.11850/68703.

Wirth:1990:DLC

[Wir90d] Niklaus Wirth. Drawing lines, circles, and ellipses in a raster. In Feijen et al.

[FvGGM90], chapter 51, pages 427–434. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326 1990. URL http:/

/www.zentralblatt-math.org/
zmath/en/search/?an=0718.

68004. This paper rederives the Bresenham line and circle drawing algorithms, and then extends the latter to handle ellipses.

Wirth:1990:MOP

[Wir90e] Niklaus Wirth. From Modula to Oberon: The programming language Oberon. Technical report 143, ETH Zürich, Zürich, Switzerland, November 1990. i + 28 pp.

URL http://hdl.handle.net/
20.500.11850/68901.

Wirth:1990:PLO

[Wir90f] Niklaus Wirth. The programming language Oberon. Note, ETH Zürich, Zürich, Switzerland, October 1, 1990. 16

pp.  URL https://people.
inf.ethz.ch/wirth/Oberon/
Oberon.Report.pdf.

Wirth:1991:RTE

[Wir91a] N. Wirth. Reply to Type-Extension Tests Can Be Performed in Constant Time. ACM Transactions on Programming Languages and Systems, 13(4): 630, October 1991. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [Coh91].

Wirth:1991:PCS

[Wir91b] Niklaus Wirth. Perspectives on computer science education. Education and Computing, 7(1–2): 105–109, January 1991. ISSN 0167-9287 (print), 1878-304X (electronic).

Wirth:1991:TCR

[Wir91c] Niklaus Wirth. Technical correspondence: Reply to “Typeextension tests can be performed in constant time”. ACM Transactions on Programming Languages and Systems, 13 (4):630, October 1991. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic).  URL http://www.
                                        acm.org/pubs/toc/Abstracts/
                                        0164-0925/214521.html. See

[Wir88e, Coh91].

Wirth:1993:RADa

[Wir93a] N. Wirth. Recollections about the development of Pascal. ACM SIGPLAN Notices, 28(3): 333–342, March 1993. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

                                        http://www.acm.org:80/pubs/
                                        citations/proceedings/plan/
                                        154766/p333-wirth/.

Wirth:1993:RADb

[Wir93b] N. Wirth. Recollections about the development of Pascal. In The Second ACM SIGPLAN Conference on History of Programming Languages, HOPL-II, pages 333–342. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-570-

4.  URL http://pascal.
hansotten.com/uploads/wirth/
Recollection%20On%20Dev%20of%
20Pascal.pdf.

Wirth:1994:GSE

[Wir94a] Niklaus Wirth. Gedanken zur Software-Explosion. (German) [Thoughts on the Software Explosion]. Informatik-Spektrum, 17(1):5–10, 1994. ISSN 0170-6012 (print), 1432-122X (electronic).

Wirth:1994:LLD

[Wir94b] Niklaus Wirth. Lola-2: a logic description language. Note, ETH Zürich, Zürich, Switzerland, 1994. 4 pp. URL https:

//people.inf.ethz.ch/wirth/
Lola/Lola2.pdf. Revised 1

September 2015.

Wirth:1995:DCD

[Wir95a] N. Wirth. Digital Circuit Design for Computer Science Students: an Introductory Textbook. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-58577-X (paperback), 3-642-57780-6 (e-book). xiii + 204 + 147 pp. LCCN TK7888.4 .W57 199.

Wirth:1995:ETT

[Wir95b] Niklaus Wirth. Effective technology transfer. The Oberon Tribune, 1(1):1–2, July

1995. URL http://pascal.
hansotten.com/uploads/wirth/

                                        Effective%20Technology%20transfer%
                                        20nw1995.pdf.

Wirth:1995:LSN

[Wir95c] Niklaus Wirth. Lola System notes. Technical report 236, ETH Zürich, Zürich, Switzerland, June 3, 1995. iii + 40 pp.

                                        URL http://hdl.handle.net/
                                        20.500.11850/69287.

Wirth:1995:PLS

[Wir95d] Niklaus Wirth. A plea for lean software. Computer, 28 (2):64–68, February 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://pascal.
                                        hansotten.com/uploads/wirth/
                                        A%20Plea%20for%20Lean%20Software.
                                        pdf;   http://pascal.hansotten.
                                        com/uploads/wirth/wirth%20a%
                                        20plea%20for%20lean%20software.
                                        pdf.

Wirth:1996:LLF

[Wir96a] N. Wirth. The language Lola, FPGAs, and PLDs in teaching digital circuit design. In Bjørner et al. [BBP96], pages 2–20. CO-DEN LNCSD9. ISBN 3-540-62064-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I5443 1996.

Wirth:1996:RAD

[Wir96b] N. Wirth. Recollections about the development of Pascal. In T. J. Bergin and R. G. Gibson, editors, History of Programming Languages—II [Proceedings of the Second Conference on the History of Programming Languages, Cambridge, Massachusetts, April 1993], pages 97–120. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-201-

89502-1. URL http://pascal.
hansotten.com/niklaus-wirth/
recollections-about-the-development-
of-pascal/;   http://pascal.
hansotten.com/uploads/wirth/
recollections.pdf.

Wirth:1996:CC

[Wir96c] Niklaus Wirth. Compiler Construction. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-40353-6 (paperback). xi + 176 pp. LCCN QA76.76.C65 W5713 1996.

Wirth:1996:GTC

[Wir96d] Niklaus Wirth. Grundlagen und Techniken des Compilerbaus. (German) [Fundamentals and Techniques of Compiler Construction]. Addison-Wesley, Reading, MA, USA, 1996. ISBN 3-89319-931-4. ix + 195 pp.

Wirth:1996:TVT

[Wir96e] Niklaus Wirth. Tasks versus threads: an alternative multiprocessing paradigm. Software

— Concepts and Tools, 17(1): 6–12, 1996. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Wirth:1998:CSMa

[Wir98a] Niklaus Wirth. A computer system for model helicopter flight control: Technical memo nr. 1: the hardware core. Report 284, ETH Zürich, Zürich, Switzerland, January 1998. ii + 9

pp. URL http://hdl.handle.
net/20.500.11850/68999;
https://people.inf.ethz.ch/
wirth/Oberon/284.pdf.

Wirth:1998:CSMb

[Wir98b] Niklaus Wirth. A computer system for model helicopter flight control technical memo nr. 2: The programming language Oberon SA. Report 285, ETH Zürich, Zürich, Switzerland, January 1998. i + 6 pp. URL https:

//people.inf.ethz.ch/wirth/
Oberon/285.pdf. Second edition

March 1999.

Wirth:1998:HCTa

[Wir98c] Niklaus Wirth. Hardware compilation: the translation of programs into circuits. Technical report 286, ETH Zürich, Zürich, Switzerland, January 1998. ii +

9 pp. URL http://hdl.handle.
net/20.500.11850/69000;
https://people.inf.ethz.ch/
wirth/Oberon/286.pdf.

Wirth:1998:HCTb

[Wir98d] Niklaus Wirth. Hardware compilation: Translating programs into circuits. Computer, 31 (6):25–31, June 1998. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://computer.
org/computer/r6025abs.htm;
 http://dlib.computer.org/
co/books/co1998/pdf/r6025.
pdf.

Wirth:1999:ADP

[Wir99a] Niklaus Wirth. Algorithmen und Datenstrukturen — Pascal-Version. (German) [Algorithms and Data Structures — Pascal Version]. Leitfäden der Informatik. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, fifth edition, 1999. ISBN 3-519-22250-7. 320 pp. LCCN QA76.9.A43.

Wirth:1999:CSM

[Wir99b] Niklaus Wirth. A computer system for model helicopter flight control. Technical memo 6, ETH Zürich, Zürich, Switzerland, March 1999. iii +

19 pp. URL https://www.
research-collection.ethz.
ch/bitstream/handle/20.500.
11850/69202/eth-4355-01.pdf.

Wirth:1999:EP

[Wir99c] Niklaus Wirth. An essay on programming. Technical Report 315, ETH Zürich, Zürich, Switzerland, 1999. i + 5

pp.  URL http://pascal.
hansotten.com/uploads/wirth/
Essay%20on%20programming%
201999%20eth315.pdf;   https:
//www.research-collection.
ethz.ch/bitstream/handle/
20.500.11850/68815/eth-4356-
01.pdf.

Wirth:2000:DPP

[Wir00a] Niklaus Wirth. The development of procedural programming languages: Personal contributions and perspectives. In Gutknecht and Weck [GW00], pages 1–10. CODEN LNCSD9. ISBN 3-540-67958-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .J6578 2000. URL http://

link.springer-ny.com/link/
service/series/0558/tocs/
t1897.htm;   http://www.springerlink.
com/content/978-3-540-67958-
5;   http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=1897.

Wirth:2000:PLD

[Wir00b] Niklaus Wirth. From programming language design to computer construction. In Böszörményi [Bös00], pages 21–

30. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1 (dpunktVerl.). LCCN QA76.756 .S36 2000. URL http:/

/digitale-objekte.hbz-nrw.
de/storage2/2016/09/08/file_
8/6866144.pdf.

Wirth:2000:SPW

[Wir00c] Niklaus Wirth. Summary of projects by N. Wirth, 1962–1999. Web site, 2000. URL https:

//people.inf.ethz.ch/wirth/
projects.html.

Wirth:2001:ESR

[Wir01a] Niklaus Wirth. Embedded systems and real-time programming. Lecture Notes in Computer Science, 2211:486–492, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/

service/series/0558/bibs/
2211/22110486.htm;   http:/
/link.springer-ny.com/link/
service/series/0558/papers/
2211/22110486.pdf.

Wirth:2001:PDS

[Wir01b] Niklaus Wirth. Program development by stepwise refinement. In Broy and Denert [BD01], pages 545–569. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN QA76.758.

URL http://link.springer.
com/chapter/10.1007/978-3-
642-48354-7_23/.

Wirth:2001:PLP

[Wir01c] Niklaus Wirth. The programming language Pascal. In Broy and Denert [BD01], pages 515–543. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN QA76.758. URL http:/

/link.springer.com/chapter/
10.1007/978-3-642-48354-7_
22/.

Wirth:2002:CSEa

[Wir02a] Niklaus Wirth. Computing science education: The road not taken. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3): 1–3, September 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wirth:2002:CSEb

[Wir02b] Niklaus Wirth. Computing science education: The road not taken. In Proceedings of the 7th Annual Conference on Innovation and Technology in Computer Science Education, ITiCSE ’02, pages 1–3. ACM Press, New York, NY 10036, USA, 2002. ISBN 1-58113-499-1.

Wirth:2002:PS

[Wir02c] Niklaus Wirth. Pascal and its successors. In Broy and Denert [BD02], pages 108– 119. ISBN 3-540-43081-4, 3-642-59412-3 (e-book), 3-642-59413-1, 3-642-63970-4 (print). LCCN QA76.758. URL https:

                                        //www.swissdelphicenter.ch/
                                        en/niklauswirth.php.

Wirth:2002:PDS

[Wir02d] Niklaus Wirth. Program development by stepwise refinement (reprint). In Broy and Denert [BD02], pages 149– 169. ISBN 3-540-43081-4, 3-642-59412-3 (e-book), 3-642-59413-1, 3-642-63970-4 (print). LCCN

                                        QA76.758. URL http://link.
                                        springer.com/10.1007/978-3-
                                        642-59412-0.

Wirth:2002:PLP

[Wir02e] Niklaus Wirth. The programming language Pascal (reprint). In Broy and Denert [BD02], pages 121–148. ISBN 3-540-43081-4, 3-642-59412-3 (e-book), 3-642-59413-1, 3-642-63970-4 (print). LCCN QA76.758. URL

                                        http://link.springer.com/
                                        10.1007/978-3-642-59412-0.

Wirth:2003:EPL

[Wir03a] Niklaus Wirth. The essence of programming languages. In László Böszörményi and Peter Schojer, editors, Modular Programming Languages, Joint Modular Languages Conference, JMLC 2003, Klagenfurt, Austria, August 25–27, 2003, Proceedings, volume 2789 of Lecture Notes in Computer Science, pages 1–11. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003.

Wirth:2003:HSC

[Wir03b] Niklaus Wirth. Hardware/ software co-design then and now. Information Processing Letters, 88(1–2):83–87, October 17, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). Special issue to honour Professor W. M. Turski’s contribution to computing science on the occasion of his 65th birthday (Warsaw, 2003).

Wirth:2004:ADS

[Wir04a] Niklaus Wirth. Algorithms and Data Structures: Oberon version. ETH Zürich, Zürich, Switzerland, 2004. 212 pp.

URL  https://github.com/
opalagin/ad2012-oberon-examples;
 https://people.inf.ethz.
ch/wirth/AD.pdf.

Wirth:2004:POD

[Wir04b] Niklaus Wirth. Programming in Oberon: a derivative of Programming in Modula-2. Report, ETH Zürich, Zürich, Switzerland, 2004. 63 pp. URL https:

//people.inf.ethz.ch/wirth/
ProgInOberon2004.pdf.

Wirth:2005:CC

[Wir05a] Niklaus Wirth. Compiler construction. Report, ETH Zürich, Zürich, Switzerland, November 2005. 131 pp. URL

                                        https://www.inf.ed.ac.uk/
                                        teaching/courses/ct/other/
                                        CBEAll.pdf.

Wirth:2005:GIT

[Wir05b] Niklaus Wirth. Good ideas, through the looking glass. Report, ETH Zürich, Zürich, Switzerland, February 2, 2005.

                                        28 pp. URL http://pascal.
                                        hansotten.com/uploads/wirth/
                                        GoodIdeas_origFig.pdf.

Wirth:2005:PPL

[Wir05c] Niklaus Wirth. PICL: A programming language for the microcontroller PIC. Note, ETH Zürich, Zürich, Switzerland, February 2005. 5 pp. URL

                                        https://people.inf.ethz.ch/
                                        wirth/PICL/PICL.pdf. Revised

October 2007.

Wirth:2005:TDC

[Wir05d] Niklaus Wirth. A triptychon of digital circuits. Note, ETH Zürich, Zürich, Switzerland, August 2005. 6 pp. URL http://

                                        pascal.hansotten.com/uploads/
                                        wirth/CounterShifter.pdf;
                                        https://people.inf.ethz.ch/
                                        wirth/Miscellaneous/CounterShifter.
                                        pdf.

Wirth:2006:GIT

[Wir06] Niklaus Wirth. Good ideas, through the looking glass. Computer, 39(1):28–39, January 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/comp/
mags/co/2006/01/r1toc.xml;
 http://csdl2.computer.org/
dl/mags/co/2006/01/r1028.
pdf;   http://pascal.hansotten.
com/uploads/wirth/Good%20Ideas%
20Wirth.pdf.

Wirth:2007:PLD

[Wir07a] Niklaus Wirth. From Programming Language Design to Computer Construction, pages 159–

164. ACM Press, New York, NY 10036, USA, 2007. ISBN 1-4503-

1049-4. URL http://pascal.
hansotten.com/uploads/wirth/
TuringAward.pdf. ACM Turing

Award lecture, reprinted from [Wir85a].

Wirth:2007:LPI

[Wir07b] Niklaus Wirth. The language PICL and its implementation. Note, ETH Zürich, Zürich, Switzerland, September 20, 2007. 10 pp. URL https:

//people.inf.ethz.ch/wirth/
PICL/PICLcompiler.pdf.

Wirth:2007:MO

[Wir07c] Niklaus Wirth. Modula-2 and Oberon. In Proceedings of the Third ACM SIGPLAN Conference on History of Programming Languages, HOPL III, pages 3.1–3.10. ACM Press, New York, NY 10036, USA, 2007. ISBN 1-59593-766-

8.  URL https://people.

                                        inf.ethz.ch/wirth/Articles/
                                        Modula-Oberon-June.doc.

Wirth:2007:OCA

[Wir07d] Niklaus Wirth. An Oberon compiler for the ARM processor. Report, ETH Zürich, Zürich, Switzerland, 2007. 48

                                        pp.  URL https://people.
                                        inf.ethz.ch/wirth/Oberon/
                                        Oberon.ARM.Compiler.pdf. Re-

vised 16 April 2008.

Wirth:2007:POC

[Wir07e] Niklaus Wirth. Porting the Oberon compiler from Oberon to Oberon-07. Note, ETH Zürich, Zürich, Switzerland, August 15, 2007. 3 pp. URL https:

                                        //people.inf.ethz.ch/wirth/
                                        Oberon/PortingOberon.pdf.

Wirth:2007:SND

[Wir07f] Niklaus Wirth. SET: A neglected data type, and its compilation for the ARM. Note, ETH Zürich, Zürich, Switzerland, August 8, 2007. 3 pp. URL https:

                                        //people.inf.ethz.ch/wirth/
                                        Oberon/SETs.pdf.

Wirth:2008:CNE

[Wir08a] Niklaus Wirth. Als Computer noch erklärbar waren. (German) [When computers were still explainable]. Essay, ETH Zürich, Zürich, Switzerland, October 3,

                                        2008. URL http://pascal.
                                        hansotten.com/uploads/wirth/
                                        Informatik68.pdf;   https:
                                        //people.inf.ethz.ch/wirth/
                                        Miscellaneous/Informatik68.
                                        pdf.

Wirth:2008:BHSa

[Wir08b] Niklaus Wirth. A brief history of software engineering. Report, ETH Zürich, Zürich, Switzerland, February

25, 2008. URL http://pascal.
hansotten.com/uploads/wirth/
IEEE-Annals.pdf.

Wirth:2008:BHSb

[Wir08c] Niklaus Wirth. A brief history of software engineering. IEEE Annals of the History of Computing, 30(3):32–39, July/September 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL https:/

/people.inf.ethz.ch/wirth/
Miscellaneous/IEEE-Annals.
pdf.

Wirth:2008:DBR

[Wir08d] Niklaus Wirth. Differences between Revised Oberon and Oberon. Note, ETH Zürich, Zürich, Switzerland, March 22, 2008. 2 pp. URL https:

//people.inf.ethz.ch/wirth/
Oberon/Oberon07.pdf. Revised

15 July 2011.

Wirth:2008:FSM

[Wir08e] Niklaus Wirth. Finite state machines, programmable logic devices, and the crux of growing complexity. Essay, ETH Zürich, Zürich, Switzerland, September 20, 2008. 18

pp.  URL http://pascal.
hansotten.com/uploads/wirth/
PLD.pdf;   https://people.
inf.ethz.ch/wirth/Miscellaneous/
PLD.pdf.

Wirth:2008:GTC

[Wir08f]  Niklaus Wirth.  Grundlagen

und Techniken des Compilerbaus. (German) [Fundamentals and Techniques of Compiler Construction]. R. Oldenbourg, München, Germany, second edition, 2008. ISBN 3-486-58581-9. xi + 191 pp. LCCN ???? URL

                                        http://www.oldenbourg-link.
                                        de/isbn/978348658581.

Wirth:2008:ITO

[Wir08g] Niklaus Wirth. Interrupts and traps in Oberon-ARM. Note, ETH Zürich, Zürich, Switzerland, February 22, 2008. 3

                                        pp.  URL https://people.
                                        inf.ethz.ch/wirth/Oberon/
                                        Interrupts.pdf.

Wirth:2008:MSE

[Wir08h] Niklaus Wirth. A microcontroller system for experimentation. Report, ETH Zürich, Zürich, Switzerland, April 16, 2008. 16 pp. URL https:

                                        //people.inf.ethz.ch/wirth/
                                        PICL/PIC.pdf.

Wirth:2008:ND

[Wir08i] Niklaus Wirth. A note on division. Note, ETH Zürich, Zürich, Switzerland, August 20, 2008.

                                        4 pp. URL http://pascal.
                                        hansotten.com/uploads/wirth/
                                        Division.pdf;   https://people.
                                        inf.ethz.ch/wirth/Miscellaneous/
                                        Division.pdf.

Wirth:2008:PS

[Wir08j] Niklaus Wirth. On programming styles. Memorandum, ETH Zürich, Zürich, Switzerland, March 24, 2008. 1 pp. URL

https://people.inf.ethz.ch/
wirth/Miscellaneous/Styles.
pdf.

Wirth:2008:WDW

[Wir08k] Niklaus Wirth. Was ist ein Denkplatz, und warum erstreben wir ihn?. (German) [What is a place of thought and why do we strive for it?]. Essay, ETH Zürich, Zürich, Switzerland, October 30, 2008. 3

pp.  URL http://pascal.
hansotten.com/uploads/wirth/
Denkplatz.pdf;   https://
people.inf.ethz.ch/wirth/
Miscellaneous/Denkplatz.pdf.

Wirth:2008:YI

[Wir08l] Niklaus Wirth. The year of informatics. Note, ETH Zürich, Zürich, Switzerland, March 2008. 9 pp. URL http://

pascal.hansotten.com/uploads/
wirth/Informatika2008.pdf;
 https://people.inf.ethz.
ch/wirth/Miscellaneous/Informatika2008.
pdf.

Wirth:2010:CSH

[Wir10a] Niklaus Wirth. Computer science: a historical perspective and a current assessment. In Sebastian Nanz, editor, The Future of Software Engineering, page 151. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010.

Wirth:2010:ECS

[Wir10b] Niklaus Wirth. Experiments in computer system design. Technical report, ETH Zürich, Zürich, Switzerland, August 2010. 60 pp. URL https:

                                        //people.inf.ethz.ch/wirth/
                                        FPGA-relatedWork/ComputerSystemDesign.
                                        pdf.

Wirth:2010:RA

[Wir10c] Niklaus Wirth. The RISC architecture. Technical report, ETH Zürich, Zürich, Switzerland, December 5, 2010. 3 pp. URL

                                        https://people.inf.ethz.ch/
                                        wirth/FPGA-relatedWork/RISC-
                                        Arch.pdf.

Wirth:2013:SHH

[Wir13] Niklaus Wirth. Selected honours and honorary doctorates. Note, ETH Zürich, Zürich, Switzerland, December 26, 2013. 1 pp. URL https:

                                        //people.inf.ethz.ch/wirth/
                                        SelectedHonours.pdf.

Wirth:2014:ILT

[Wir14]  Niklaus Wirth.  An imple-

mentation of Lola-2, or, translating from Lola to Verilog N.Wirth, 30.11.201. Technical report, ETH Zürich, Zürich, Switzerland, November 30, 2014. 7 pp. URL https:

                                        //people.inf.ethz.ch/wirth/
                                        Lola/LolaCompiler.pdf.

Wirth:2015:CCP

[Wir15a] Niklaus Wirth. Computers and computing: a personal perspective. Note, ETH Zürich, Zürich, Switzerland, December

2015. URL http://pascal.
hansotten.com/uploads/wirth/
Informatika2008.pdf;   https:
//people.inf.ethz.ch/wirth/
Miscellaneous/ComputersAndComputing.
pdf.

Wirth:2015:DRA

[Wir15b] Niklaus Wirth. The design of a RISC architecture and its implementation with an FPGA. Report, ETH Zürich, Zürich, Switzerland, September 1, 2015. 24 pp. URL https:

//people.inf.ethz.ch/wirth/
FPGA-relatedWork/RISC.pdf.

Wirth:2015:PT

[Wir15c] Niklaus Wirth. Programming:

a tutorial. Note, ETH Zürich, Zürich, Switzerland, October 5, 2015. 64 pp. URL https:

//people.inf.ethz.ch/wirth/
Oberon/PIO.pdf.

Wirth:2015:TC

[Wir15d] Niklaus Wirth. Three counters. Note, ETH Zürich, Zürich, Switzerland, October 2015. 3 pp. URL https:

//people.inf.ethz.ch/wirth/
FPGA-relatedWork/ThreeCounters.
pdf.

Wirth:2016:PLO

[Wir16]  Niklaus Wirth.  The pro-

gramming language Oberon. Technical report, ETH Zürich, Zürich, Switzerland, May 3, 2016. 17 pp. URL https:

//people.inf.ethz.ch/wirth/
Oberon/Oberon07.Report.pdf.

Wirth:2017:RAT

[Wir17a] Niklaus Wirth. 9. a RISCarchitecture as target. Technical report, ETH Zürich, Zürich, Switzerland, September 25, 2017. 64 pp. URL

                                        https://people.inf.ethz.ch/
                                        wirth/CompilerConstruction/
                                        CompilerConstruction2.pdf.

Wirth:2017:CC

[Wir17b] Niklaus Wirth. Compiler construction. Report, ETH Zürich, Zürich, Switzerland, May 2017. 44 pp. URL

                                        https://people.inf.ethz.ch/
                                        wirth/CompilerConstruction/
                                        CompilerConstruction1.pdf.

This is a slightly revised version of the book published by Addison-Wesley in 1996 ISBN 0-201-40353-6.

Wirth:2018:ORO

[Wir18] Niklaus Wirth. Oberon-07 (revised Oberon) (Oberon at a glance). Note, ETH Zürich, Zürich, Switzerland, November 20, 2018. 1 pp. URL https:

                                        //people.inf.ethz.ch/wirth/
                                        Oberon/OberonAtAGlance.pdf.

Wirth:2021:VYP

[Wir21]  Niklaus Wirth.  Viewpoint:

50 years of Pascal. Communications of the ACM, 64(3): 39–41, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL https://cacm.acm.org/
                                        magazines/2021/3/250705-50-
                                        years-of-pascal.

Wirth:1975:MLM

[Wir5 a] Niklaus Wirth. MODULA: a language for multiprogramming. Report 18, ETH Zürich, Zürich, Switzerland, 1975 (??).

Wirth:1975:UMD

[Wir5 b] Niklaus Wirth. The use of MOD-ULA and design and implementation of MODULA. Report 19, ETH Zürich, Zürich, Switzerland, 1975 (??).

Wortman:1992:CCMa

[WJ92a] David B. Wortman and Michael D.

Junkin. A concurrent compiler for Modula-2+. In Proceedings of the ACM SIGPLAN 1992 Conference on Programming Language Design and Implementation, PLDI ’92, pages 68–81. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-475-9.

Wortman:1992:CCMb

[WJ92b] David B. Wortman and Michael D.

Junkin. A concurrent compiler for Modula-2+. ACM SIGPLAN Notices, 27(7):68–81, July 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wirth:1993:EBF

[WL93] Niklaus Wirth and Stefan Ludwig. An extension-board with an FPGA for experimental circuit design. Technical report 198, ETH Zürich, Zürich, Switzerland, 1993. 44 pp.

URL http://hdl.handle.net/
20.500.11850/68988.

Woodman:1986:FSM

[Woo86] Mark Woodman. Formatted syntaxes and Modula-2. Software —Practice and Experience, 16 (7):605–626, July 1986. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Woodman:1993:TMS

[Woo93] Mark Woodman. A taste of the Modula-2 Standard. ACM SIGPLAN Notices, 28(9):15–24, September 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Woodman:1994:PLS

[Woo94] Mark Woodman. Programming language standards scene, ten years on paper 10: Modula-2. Computer Standards and Interfaces, 16(5-6):487–494, September 1994. CODEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Welsh:1972:PCI

[WQ72] J. Welsh and C. Quinn. A PAS-CAL compiler for ICL 1900 series computers. Software—Practice and Experience, 2(1):73– 77, January/March 1972. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wiener:1984:SEM

[WS84] Richard S. Wiener and Richard Sincovec. Software engineering with Modula-2 and Ada. Wiley, New York, NY, USA, 1984. ISBN 0-471-89014-6. xix + 451 pp. LCCN ????

Wiener:1985:TAI

[WS85] Richard S. Wiener and Richard F.

Sincovec. Two approaches to implementing generic data structures in Modula-2. ACM SIGPLAN Notices, 20(6):56–64, June 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Welsh:1977:AIP

[WSH77] J. Welsh, W. J. Sneeringer, and C. A. R. Hoare. Ambiguities and insecurities in Pascal. Software —Practice and Experience, 7(6): 685–696, November/December 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See also [Ker81, Ker84].

Wirth:1965:EGA

[WW65] Niklaus Wirth and Helmut Weber. Euler: a generalization of Algol, and its formal definition. Technical Report CS20, Stanford University, Department of Computer Science, Stanford, CA, USA, April 27, 1965. xv + 115

pp. URL http://i.stanford.
edu/pub/cstr/reports/cs/tr/
65/20/CS-TR-65-20.pdf.

Wirth:1966:EEG

[WW66a] Niklaus Wirth and Helmut Weber. Errata: “EULER: a generalization of ALGOL and it formal definition: Part 1”. Communications of the ACM, 9(12): 878, December 1966. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [WW66b].

Wirth:1966:EGAa

[WW66b] Niklaus Wirth and Helmut Weber. EULER: a generalization of ALGOL and its formal definition: Part 1. Communications of the ACM, 9(1):13–25, January 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See errata [WW66a].

Wirth:1966:EGAb

[WW66c] Niklaus Wirth and Helmut Weber. EULER: a generalization of ALGOL, and its formal definition: Part II. Communications of the ACM, 9(2):89–99, February 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wirth:1968:PS

[WWS68] Niklaus Wirth, Joseph W. Wells, Jr., and Edwin H. Satterthwaite, Jr. The PL360 system. Technical Report CS 91, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1, 1968. viii + 89

                                        pp. URL http://i.stanford.
                                        edu/pub/cstr/reports/cs/tr/
                                        68/91/CS-TR-68-91.pdf.

Wyant:1994:IM

[Wya94]  Geoff Wyant.  Introducing

Modula-3. Linux Journal, 8:??, December 1994. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Yip:1984:PGS

[Yip84] C. K. Yip. The Pascal graphics system. Software—Practice and Experience, 14(2):101– 118, February 1984. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Young:1981:ISL

[You81] S. J. Young. Improving the structure of large Pascal programs. Software—Practice and Experience, 11(9):913–927, September 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zamulin:1996:ASO

[Zam96] A. V. Zamulin. Algebraic semantics of the Oberon target machine. In Bjørner et al. [BBP96], pages 41–54. CODEN LNCSD9. ISBN 3-540-62064-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I5443 1996. URL http://

link.springer.com/chapter/
10.1007/3-540-62064-8_5.

Zeller:1997:SIO

[Zel97] E. Zeller. Seamless integration of online services in the Oberon document system. Lecture Notes in Computer Science, 1204:366– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zimmer:1985:MMW

[Zim85] J. A. Zimmer. A modest Modula wish list. ACM SIGPLAN Notices, 20(11):69–77, November 1985. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).